### v1

# FLex&Chill: Improving Local Federated Learning Training with Logit Chilling

Abstract
Federated learning are inherently hampered by data heterogeneity: non-iid distributed training data over local clients. We propose a novel model training approach for federated learning, FLex&\&&Chill, which exploits the Logit Chilling method. Through extensive evaluations, we demonstrate that, in the presence of non-iid data characteristics inherent in federated learning systems, this approach can expedite model convergence and improve inference accuracy. Quantitatively, from our experiments, we observe up to 6$\times$ improvement in the global federated learning model convergence time, and up to 3.37% improvement in inference accuracy.
1 Introduction
Federated learning represents a paradigm enabling the training of effective models within distributed environments without explicitly exposing local data McMahan et al. (2017). This approach provides opportunities for communication and computation efficiency during training by sharing model parameters/gradients learned independently from local leaning processes while suppressing raw data transmissions. Furthermore, since training operations take place in a de-centralized manner, federated learning opens the possibility to parallelize the model training process via the assistance of local client devices. It presents an attractive solution for handling and exploiting the vast amounts of data generated by many mobile and Internet of Things (IoT) device entities consisting the overall system Park et al. (2023a).
In contrast to conventional centralized training methods, federated learning faces challenges associated with delayed model convergence and limited performance. A fundamental challenge of federated learning arises from the disparity of training data used at each client device. Federated stochastic optimization encounters difficulties in determining optimal parameters, leading to increased communication overhead between local devices and the central server, as well as the need for additional training on (typically resource limited) local devices Zhao et al. (2018).
To address these challenges, a number of prior research have focused on optimizing the local training operations on federated learning clients or on refining the server-side aggregation process. For instance, FedProx Li et al. (2020) controls the number of iterations for each local device, aiming to train models resilient to challenges posed by non-independent and non-iid data environments. SCAFFOLD Karimireddy et al. (2020) achieves expedited convergence and improved model accuracy McMahan et al. (2017) by introducing a correction term during the model aggregation phase to balance the influence of each client. These operations alleviate the problems posed by the non-iid environment, a common dataset configuration for federated learning clients.
This paper proposes a novel model training approach for federated learning, FLex&\&&Chill (Federated Learning EXploiting logit Chilling), that targets to boost the convergence and model performance via temperature scaling, namely logit chilling, during the local model training operations. Note that a typical neural network generates a probability ($q_{i}$) for each class from its neural network logits ($z_{i}$) by applying a softmax operation (Eq. 1) with the temperature $T$ conventionally set to 1. However, it is not uncommon in other application scenarios to exploit even higher temperature values (i.e., $T\in(1,\infty)$), for example, in knowledge distillation Hinton et al. (2015); Touvron et al. (2021) or neural network calibration Guo et al. (2017).
$\xi(z,i,T)=q_{i}=\frac{exp(z_{i}/T)}{\sum_{j}exp(z_{j}/T)}$ (1)
When employing a temperature $T$ higher than 1, the probability distribution across classes naturally undergoes smoothing due to the narrowed interval between the logits Guo et al. (2017), making such an approach suitable for applications such as knowledge distillation. Notably, despite altering the probabilities, the order of labels’ probability remains unaffected with respect to changes in $T$. While these works successfully explore the impact of varying temperatures for neural network operations, they have mainly focused on applying high temperatures after the model training process, mostly for the inference operations. Raising $T$ to a higher value reduces the confidence of the training operations. This is especially useful to tame DNNs which are typically overconfident.
On the other hand, in federated learning, where disparate training data is common, increasing the model confidence of each individually trained client model (using a lower $T$) prior to their parameter aggregation can be a more effective approach. Our investigation involves measuring both the gradient flow and the data position shift within the representation space to provide guidelines in this unexplored territory.
Specifically, this work targets to show that utilizing fractional temperature values (i.e., $T\in(0,1)$) during the training process enables effective gradient propagation towards the input layer when applied to federated learning data distribution scenarios. Consequently, we observed that lower temperatures facilitate more efficient data updates within the representation space when dealing with non-iid datasets, a common characteristic of federated learning systems. From our evaluations with three datasets and three baseline federated learning models, we show that these effects accelerate the federated learning model convergence time by up to $\mathbf{6.00\times}$ and improves the inference accuracy by up to $\mathbf{3.37\%}$, despite potential concerns on model instability when training with with low temperatures. We summarize the contributions of this work in three-fold as the following:
  • To the best of our knowledge, this is the first work that explores the impact of integrating temperature scaling during the training process in a federated learning scenario.
  • We offer empirical evidence that demonstrates the performance enhancements of exploiting logit chilling in federated learning scenarios, by examining the gradient flow and data position shifts within the representation space.
  • We introduce a novel federated learning model training approach, FLex&\&&Chill, showcasing the application of low temperature usage during the training process.
2 Background and Related Work
2.0.1 Temperature scaling
In machine learning, the output probability distribution generated by a model serves diverse purposes. Primarily, classification models offer probabilities for each class, showcasing the model’s confidence in its predictions. This aids in assessing how certain the model is in its predictions (i.e., classification outputs) and allows consideration of the likelihood of each class prior to making a final classification decision. Moreover, this distribution offered by the model helps in understanding the uncertainty of a model on its decisions Pearce et al. (2021); Park et al. (2023b).
In neural networks, computing probabilities from logits often involves employing softmax operations, as shown in Eq. 1. Here, the temperature parameter ($T>0$) controls the probability distribution’s concentration towards one class, and is used to manage how the categorical probability is dispersed over the possible classes Jang et al. (2016). Notably, this categorical probability aligns with the concept of entropy in information theory, offering a perspective where the temperature parameter governs the level of disorder or uncertainty within the distribution. Specifically, when $T\in(1,\infty)$ is applied, the probability distribution tends to be smoother due to increased $T$. This arises from the diminished differences between logit values, enhancing the distribution’s entropy. In contrast, with $T\in(0,1)$ within the softmax function, the probability distribution becomes sharper, since lower temperatures amplify the differences between logits, effectively polarizing the distribution and concentrating probabilities closer to the maximum logit value. Consequently, the model becomes more confident in its predictions, as it allocates higher probabilities to fewer classes, thereby reducing uncertainty in the distribution.
Therefore, temperature scaling operates as a means to adjust the probability distribution within a neural network-based system. Previous work in Knowledge Distillation Hinton et al. (2015); Touvron et al. (2021) exploit the temperature value by applying a high temperature $T\in(1,\infty)$ to achieve a smoothed probability distribution. This manipulation aims to magnify small values within the distribution, often representing hidden information encapsulated within the model. Additionally, in generative models Shih et al. (2023); Wang et al. (2020), temperature scaling regulates the randomness inherent in model outputs. This technique allows for the control of variability levels in generated samples, enabling adjustments to generated output diversity. Temperature scaling is also used in in model calibration Guo et al. (2017). Applying temperature adjustments to the model’s output probabilities, typically through a scaling factor $T$, aids in aligning these probabilities more accurately with the true likelihoods or uncertainties present in the data. Likewise, previous work have focused on applying high temperature values for various purposes, mostly for the purpose of conservatively better-understanding the full dataset without bias.
2.0.2 Federated learning
In recent years, federated learning has emerged as a promising paradigm in machine learning, particularly in scenarios where data privacy, security, and the decentralization of model training are core concerns. Unlike traditional centralized training approaches, federated learning enables model training across distributed devices or servers while keeping the data localized; thus, addressing privacy challenges associated with centralized data aggregation. This decentralized learning paradigm involves iterative model training where local updates occur on individual devices, and only aggregated model updates are shared with the central server.
Despite its benefits, federated learning poses several challenges. A body of work have tackled the issue of communication overhead stemming from the exchange of frequent model parameter updates. A number of previously proposed work have proposed schemes to effectively abstract the model parameters in various forms to relieve the communication and networking overhead Park et al. (2023a). Furthermore, federated learning systems are prone to facing issues rising from data heterogeneity caused from the diversity or differences in the characteristics, distributions, or formats of data across federated learning clients. Previous work have attempted to address this challenge by addressing relevant issues such as variations in feature/input variations Yu et al. (2021) and spatio-temporal variations Liu et al. (2020).
Figure 1: Distribution of gradient norm observed at the input layer for correctly and incorrectly inferenced samples with varying temperature values. We observe noticeable improvements in gradient flow to the when training with lower temperatures.
3 Design of FLex&Chill
We initiate our discussions by exploring the impact of logit scaling, also known as temperature scaling, providing empirical evidence on the rationale behind our proposed system. Furthermore, in this section, we introduce our system, referred to as FLex&\&&Chill, a novel local training approach designed for efficiently training a federated learning framework.
3.1 Impact of Training with Lower Temperatures
As aforementioned, conventionally, temperature scaling has primarily been focused on calibrating the probability distribution of a model’s output after the model training phase, with limited exploration into its application during the training process itself. In this section, we present empirical results on three initial experiments specifically crafted to showcase the implications of employing temperature scaling throughout the training phase operations in different dimensions. Through these investigations, our aim is to establish a foundation for comprehending the potential implications and advantages associated with the utilization of temperature scaling techniques in neural network training. We conduct our experiments with the CIFAR10 dataset Krizhevsky et al. (2009) and use a 2-layered CNN architecture McMahan et al. (2017) as the baseline neural network architecture.
3.1.1 Gradient flow patterns
Our first experiment focuses on assessing the norm of computed gradients within the input layer for various temperature values ($T$), and we present sample distribution plot of how the gradients flow for 50K training samples in Figure 1. Note that here we separate the plots for the correctly inferenced samples (in the foward pass) and the incorrect ones to observe their differences. The propagation of gradients to the input layer serves as an indicative metric for the model’s learning efficacy. Successful gradient propagation to the input layer implies that the model effectively utilizes and comprehends information from diverse features within the input data during the training phase Evci et al. (2022).
As shown in Figure 1, the mode of distribution shifts to larger values with decreasing temperatures for both correct and incorrect predictions. One reason behind this mode shift stems from the gradient function characteristics of softmax. Eq. 2 presents the gradient of cross-entropy for the softmax activation with varying temperature $T$ and hard target, one-hot encoded labels. For brevity, we present the final derivation result and leave details in the Appendix. Since the gradient is multiplied by $\frac{1}{T}>1$, for low temperatures $T<1$, we can expect a gradient boost.
$\frac{\partial\xi}{\partial{z_{i}}}=\frac{1}{T}(p_{i}-y_{i})=\frac{1}{T}(\frac{e^{z_{i}/T}}{\sum_{j}e^{z_{j}/T}}-y_{i})$ (2)
Furthermore, from the perspective of the loss term $(p_{i}-y_{i})$, when the model yields an incorrect decision for a given sample, the loss term significantly increases as $p_{i}\xrightarrow{}0$ (more details later in this section). The double-folded effect accounts for substantial mode shift with thick and long tails in the gradient distribution for lower training temperatures.
3.1.2 Data position shift in the representation space
To better understand the impact of temperature scaling on the training process, we conducted a study into the shift of a data’s position in the representation space (i.e., the distance to the decision boundary). Drawing inspiration from the concept of ‘travel to decision boundary’ as proposed by Kim et al. Kim et al. (2022), we quantified the distance of a data point’s inference result to the decision boundary by determining the magnitude of epsilon ($\epsilon$) value multiplied by the adversarial perturbation Goodfellow et al. (2014) needed to alter the model’s decision.
For this experiment, we configure an initial neural network and perform inference for 1,000 samples to identify the results’ positions on the representation space. Among the 1,000, we select the ones that were misclassified, and recompute their positions after performing one training operation using that specific data point and compute the difference in positions on the representation space with respect to the decision boundary. By doing so, we can observe how drastically one training operation affects the next inference output with respect to different temperatures used in the training phase.
Figure 2: Distributions depicting differences between distances to the decision boundary before and after model updates for varying training temperatures. Notice that lower temperatures show noticeable shift in estimations’ positions on the representation space, suggesting their aggressiveness in modifying the model even with a small number of training samples.
Figure 2 plots our results in the form of distributions depicting differences between distances to the decision boundary before and after model updates. When higher temperatures ($T={2,4}$) are applied, we can notice that the changes in inference results’ positions are not significant (i.e., plots skewed towards 0). Whereas, with a small $T$, we can notice more pronounced shifts in data positions within the representation space. This suggests that using a small $T$ at the softmax operations can actively steer the model towards (potentially) substantial optimization to better align with the training data.
3.1.3 Impact of temperature on individual samples
Figure 3: Probability distribution for single training sample (class index 4) with different training temperatures. A low $T$ hallucinates the model’s probability distribution, causing the neural network to be trained more aggressively even with a small number of samples.
We now examine the impact of different degrees of temperatures on a single sample prediction. Specifically, we take a look at the probability distribution of prediction for a single sample with varying temperature during training process. In Figure 3 we present the the resulting probability distribution for a single training sample (with ground truth index of 4). As the plots show, a higher temperature, $T$=4 in this case, leads to a widely spread probability distribution, and the neural network tends to less confident in prediction. On the other hand, with lower temperatures the information embedded in the data sample becomes more concentrated to the predicted label. If the prediction is incorrect, this increased prediction on the wrong label results in reduced probability on the correct label and increases the loss. This leads to more aggressive changes despite training from a single sample.
In federated learning, clients will train themselves with a small amount of local data (much smaller than typical centralized training). For the models trained at the clients to hold value at the server during the aggregation process, it is important that the local data characteristics are effectively embedded in its model parameters despite the small local dataset size. The aggressive behavior when applying low training temperatures that we see in Figure 3 and throughout our initial experimental results serve as evidence that the core information from each training element can be effectively extracted and learned at the neural network when low temperatures are applied in the federated learning training process.
3.2 FLex&\&&Chill and Logit Chilling
Based on such empirical observations, we hypothesise that training neural networks with lower temperatures can benefit the model convergence and performance for systems exploiting with non-iid data, which is common for federated learning scenarios. In FLex&\&&Chill we term this as logit chilling, as a we ‘chill down’ the temperature at the softmax operations, thereby altering the logit computation derived from the neural network during the training process. The focus of a FLex&\&&Chill-based federated learning system is to exploit logit chilling to train a client-side model with low temperature values. As we observed through the studies above and as our evaluations will show, this has an effect of effectively exploiting crucial information from the training samples to actively alter the model towards the new knowledge.
The lower temperatures used by Logit Chilling directly influences the local training process on each federated learning client. Notably, the seamless integration of Logit Chilling and FLex&\&&Chill with existing federated learning frameworks is a distinct advantage, allowing for a straightforward implementation within already deployed frameworks. Furthermore, we note that FLex&\&&Chill can be used orthogonality with other efforts that accelerate federated learning training operations such as FedProx and SCAFFOLD.
We formulate the operations of FLex&\&&Chill using Algorithm 1. Here, typical federated learning operations take place at both the server and clients except for the fact that the server determines a target $T$ at each round and local model training operations are performed with $T$.
Algorithm 1 FLex&\&&Chill
Data ($D_{1},D_{2},...,D_{N}$) where $D_{i}$ is the user $i$’s local data.
Model ($w^{1}_{t},w^{2}_{t},...,w^{N}_{t}$) where $w_{t}^{i}$ is the user $i$’s model in $t$-th federated round.
Parameter: Temperature $T\in(0,1)$
1:Server executes:
2:Initialize $w_{0}$
3:for each round $t=1,2,...,R$ do
4:    $S_{t}\xleftarrow{}$ (random set of $m$ clients)
5:    for $k\in S_{t}$ do
6:         $w_{t}^{k}\xleftarrow{}$ ClientUpdate($w_{t},T$)
7:    end for
8:    for $k\in S_{t}$ do
9:         $w_{t+1}\xleftarrow{}$ Aggregation($w_{t}^{k}$)
10:    end for
11:end for
12:ClientUpdate($w_{t}$, $T$):
13:$w_{t}^{k}\xleftarrow{}w_{t}$
14:for data, label $\in D_{k}$ do
15:    $\mathbf{z}=w_{t}^{k}(data)/T$
16:    $l(\mathbf{z})=loss(\mathbf{z};label)$
17:    $w_{t}^{k}\xleftarrow{}w_{t}^{k}-\eta\nabla l(\mathbf{z})$
18:end for
19:Return $w_{t}^{k}$ to server
4 Experiment
We evaluate the performance of FLex&\&&Chill using three datasets: (i) FEMNIST from the LEAF database Caldas et al. (2018b), including the Extended MNIST handwriting data Cohen et al. (2017) , (ii) the CIFAR10 dataset, and (iii) the CIFAR100 dataset Krizhevsky et al. (2009). As metrics, we use the average accuracy and training loss observed at the clients, selected to observe both the performance gain and the effectiveness of FLex&\&&Chill in its training operations. Furthermore, we also present the number of federated learning rounds needed to achieve a pre-defined target accuracy.
(a) FEMNIST-DNN
(b) CIFAR10-CNN
(c) CIFAR100-ResNet18
Figure 4: Average federated learning client-side model accuracy for different temperatures used in training. Overall, lower temperatures show higher average model accuracy for federated learning systems.
4.1 Experiment Setup
Dataset. We detail the datasets used in our work as follows.
  • FEMNIST dataset: Sourced from the LEAF dataset, the FEMNIST dataset comprises of 10 handwritten digits (0-9) and 52 characters (26 lowercase and 26 upper case) images contributed by 712 users, totaling 157,132 samples. We selected a subset of 36 users, each with diverse amounts of local data to account for the non-iid environment. We followed the implementations provided by the original work Caldas et al. (2018a).
  • CIFAR10 and CIFAR100 dataset: We use the baseline CIFAR10 and CIFAR100 datasets Krizhevsky et al. (2009) and manually partition the data over 50 federated learning users, with each user holding 1,000 samples.
Model. For each dataset, we utilized distinct model architectures, taking into account the specific challenges and characteristics inherent to each dataset. For FEMNIST, an MLP model featuring four linear layers was employed. We term this combination as “FEMNIST-DNN” in the rest of our evaluations. In the case of the CIFAR10 and CIFAR100 datasets, experiments were conducted using a two-layer CNN model McMahan et al. (2017) and a ResNet18 He et al. (2016), respectively. These two cases are denoted as “CIFAR10-CNN” and “CIFAR100-ResNet18.”
Experiment Details. Each case underwent local training for 300 federated training rounds, utilizing FedAvg McMahan et al. (2017) as the baseline federated learning framework. FedAvg is a widely used framework for federated learning. In each round, the server selects 10 participating clients, employing Stochastic Gradient Descent (SGD) as the optimizer with a learning rate of 0.001 Ruder (2016). Each client trained its model for 10 local epochs, employing a batch size of 16, and utilized cross-entropy loss as the loss function.
4.2 Average Client-side Accuracy
(a) FEMNIST-DNN
(b) CIFAR10-CNN
(c) CIFAR100-ResNet18
Figure 5: Federated learning client-side model training loss observed for different training temperature values. We can notice that overall a lower training loss is observed when lower temperature values are used for model training.
We report the average client-side model accuracy results obtained for models trained with different temperatures in Figure 4. For all three datasets (and the three respective models used for each test case) we see an increasing accuracy trend observed for lower temperatures. Note that for the FEMNIST-DNN case with $T$=0.05, we noticed that the accuracy converges after 120 rounds; thus, perform an early exit for this case. One interesting aspect we can notice is that the accuracy for the $T$=0.05 case, which is the lowest temperature we test for, does not show the best performance for the CIFAR100-ResNet18 experiment. While the $T$=0.05 case shows the best performance of FEMNIST-DNN and CIFAR10-CNN, for CIFAR100-ResNet18, we see a accuracy worse than the $T$=0.5 case. This suggests that training with the lowest possible temperature value is not always ideal, rather this result highlights the importance of identifying a proper training temperature as we will further discuss in Section 5. A closer zoom into the plots suggest that the best performing configuration can change over the training round and can vary for different datasets, but overall, we can notice that exploiting a temperature value lower than “1” in the training process can help improve the classification accuracy.
The average accuracy for all federated learning clients after 300 training rounds presented in Table 1 confirm our observations, indicating that with $T$=0.05 the CIFAR10 case shows a 3.37% improvement in accuracy compared to the $T$=1 case. Even for other cases, the use of a low temperature value show promising results in terms of accuracy. We emphasize once more that this does not mean that the accuracy will always perform better at the lowest temperature. Rather, our results suggest the need to carefully explore the fractional temperature space in the federated learning training process.
4.3 Training Loss
Next we compare the training loss observed at the clients. In Figure 5 we present the average the loss observed at each federated learning training round. Again, we note that we perform early exit for the FEMNIST-DNN case at 120 rounds when $T$=0.05 due to quick model convergence. We can notice there that the use of lower temperatures in the training phase generally helps quickly lower the training loss during the local training phase for federated learning clients. At the same time, when observing the CIFAR100-ResNet18 plots, we can see that the loss for $T$=0.05 does not converge well. In fact, this serves as a reason behind the lower accuracy reported for CIFAR100-ResNet18.
From the quantitative values in Table 1 where we plot the average loss observed at all federated learning clients after 300 training rounds (120 for the FEMNIST-DNN @ $T$=0.05), we can make similar observations. Especially, for FEMNIST-DNN, we see a six-fold improvement in loss values compared to $T$=1. For the CIFAR100-ResNet18 case, however, we notice that the lowest loss convergence is in fact the $T$=0.25 case rather than 0.05. This observation confirms the need to properly consider and explore fractional temperature values for client-side model training.
4.4 Federated Learning Convergence Analysis
FEMNIST-DNN CIFAR10-CNN CIFAR100-ResNet18 Temperature ($T$) Accuracy (%) Loss Num. of Round Accuracy (%) Loss Num. of Round Accuracy (%) Loss Num. of Round $T=4.0$ 73.60 (-5.06) 0.893 299 (0.44×) 37.58 (-3.76) 1.724 299 (0.84×) 28.05 (-0.20) 2.984 288 (0.69×) $T=2.0$ 77.20 (-1.46) 0.749 205 (0.64×) 40.35 (-0.99) 1.638 270 (0.93×) 27.89 (-0.36) 2.918 288 (0.69×) $T=1.0$ 78.66 (0.00) 0.684 132 (1.00×) 41.34 (0.00) 1.594 251 (1.00×) 28.25 (0.00) 2.918 200 (1×) $T=0.5$ 79.78 (+1.13) 0.654 88 (1.50×) 41.52 (+0.18) 1.551 161 (1.56×) 29.77 (+1.52) 2.875 178 (1.12×) $T=0.25$ 79.50 (+0.85) 0.650 57 (2.32×) 43.41 (+2.07) 1.532 132 (1.90×) 30.63 (+2.38) 2.997 152 (1.32×) $T=0.05$ 80.86 (+2.20) 0.643 22 (6.00×) 44.71 (+3.37) 1.528 106 (2.37×) 27.05 (-2.72) 4.228 239 (0.74×)
Table 1: Average classification accuracy, model training loss, and number of federated learning training rounds needed to achieve the maximum accuracy of the $T$=4 case for each dataset-model combination. Quantitative results suggest the effectiveness of exploiting low temperatures for model training in federated learning both in terms of model accuracy and training efficiency.
Centralized - iid Federated - iid Federated - non-iid $T$ Acc. (%) Loss Acc. (%) Loss Acc. (%) Loss 4.0 58.05 1.152 48.22 1.473 37.58 1.724 2.0 63.27 1.041 51.90 1.403 40.35 1.638 1.0 65.18 1.044 54.66 1.325 41.34 1.594 0.5 64.19 1.030 53.19 1.320 41.52 1.551 0.25 63.88 1.092 54.64 1.265 43.41 1.532 0.05 60.98 1.227 47.60 1.701 44.71 1.528
Table 2: Average inference accuracy and model training loss for CIFAR10-CNN with different configurations. The impact of FLex&\&&Chill is prominent for federated learning scenarios with non-iid datasets, whereas its impact is less when elsewise.
Finally, the last columns for each test case in Table 1 we present the number of federated learning training rounds needed to achieve a preset target accuracy for different temperature values using for model training. This number provides an understanding of how efficient the local training processes are when different temperature values are applied. S
[…]
From the results in Table 1 we can notice that with $T$=0.05 the FEMNIST-DNN case achieves the target accuracy after only 22 training rounds. We see similar trends of low temperatures positively affecting the training efficiency throughout our results. This is from the generally aggressiveness of training behaviors when applying lower temperature training. Given a single training sample, while high temperature-based training conservatively absorbs the available information, lower temperatures are more progressive in making model changes. Given the small datasets used for local training and the need to effectively deliver its knowledge to the server for aggregation, we see FLex&\&&Chill as an effective way to train client-side models in the federated learning process.
4.5 Impact of Dataset Dispersion and Federated Learning on FLex&\&&Chill
We now examine the impact of FLex&\&&Chill on different model training configurations. Specifically, using the CIFAR10-CNN test case, we examine two additional configurations: (i) centralized learning with iid data and (ii) federated learning with iid data. Note that the federated learning network is identical to our previous experiments except for the dataset distribution and was tested for 300 rounds. By comparing these results with the non-iid federated learning results, we target to understand the impact of FLex&\&&Chill on federated learning and dataset disparity in greater detail.
As Table 2 shows, applying FLex&\&&Chill, thus a lower $T$ does not induce a positive impact on the performance of centralized learning scenarios. In fact, for centralized learning, controlling $T$ for the training process does not show a noticeable impact. We see the same phenomena for the federated learning system with iid samples as well, suggesting that the performance enhancements we see in Table 1 is not an effect of the federated learning process itself. We do point out that holding iid datasets in federated learning will show overall higher performance compared to non-iid configurations (regardless from $T$), but is not a practical case to assume when deploying federated learning systems. Rather, as we discussed, the impact of exploiting low training temperatures is beneficial when dealing with non-iid datasets.
Figure 6: Accuracy drop observed before and after server-side model aggregation. The aggregation process of client models trained with lower temperatures better retain the individual knowledge.
4.6 Model stability of low temperature training during federated learning aggregation
As discussed in Section 3, low temperatures can expedite local training even when dealing with a small amount of training data. Nevertheless, the effect of low temperature training on the model generalization and aggregation stability of federated learning has not yet been thoroughly studied. To inspect the stability of federated learning aggregation for models trained with low temperatures, we measured the the performance change of each client-side model before and after the server-side aggregation round.
Figure 6 plots the changes in accuracy observed for both participating and non-participating clients at each federated learning round. Specifically, once the model is aggregated, each client computes its accuracy with both the pre-update and the post-update model and computes their difference. Note that pre-update models for clients participating in the round would have updated their local models (very recently) with their own data. Thus, the pre-model performance will be high. However, since each model is trained aggressively over heterogeneous data, the aggregation process could puddle up this knowledge and the aggregated model may lose information. Ideally, with their local information properly aggregated, the performance drop will be minimal. Non-participating clients on the other hand, would hope to have an improved updated model the federated learning process.
As Figure 6 shows the models trained with low temperatures retain the knowledge better than the medium and hot models for participating clients (solid plots groups at bottom). At the early rounds, the models with high temperature show less performance drop, but the fast learning speed of low temperature model amortizes the performance drop, which results in faster learning of the aggregated model. Furthermore, we can also notice that the aggregation process does not interfere with the performance of non-participating clients (grouped on top), implying that federated learning with low temperature training preserves the generalization power.
5 Discussion
$\bullet$ Identifying the optimal temperature. Our study shows the effectiveness of exploiting lower temperatures as a way to enhance federated learning performance. However, an important relevant question remains unanswered: “How do we determine the appropriate or optimal temperature for different applications?” With varying dataset and models that different federated learning application use, the ideal temperature value will also differ. The scope of this work is on identifying and quantifying the potential benefits of lower temperature-based local training in federated learning and we leave the task of identifying per-purpose optimal temperatures to individual system designers and future research.
$\bullet$ Dynamic per-sample temperature scaling. While our paper focuses on understanding the advantages of maintaining a constant low temperature for the training process, this may not be the ideal approach when targeting maximum performance. In reality, varying temperatures can be applied in the training process with respect to their significance Shin et al. (2022); Zeng et al. (2021). That would mean that the training process can exploit lower temperatures for the samples that is “harder” to learn, while higher temperatures are used for those easier to understand. Alternatively, a more course-grain approach of dynamically adjusting the temperature on a per-federated learning round-basis can be another meaningful approach to adapt towards changing local data characteristics.
$\bullet$ Supporting heterogeneous federated learning. While many federated learning frameworks assume homogeneous devices, recent work have pointed out that this is not always the case Park and Ko (2023). In a federated learning network consisting of devices with heterogeneous computing/network capabilities, the issue of “stragglers”, which are devices that take longer to complete local training, becomes important to address. Leveraging insights gleaned from our empirical observations, applying lower temperatures for potential stranglers can be an interesting way of uniforming the local training duration. This would require a prior analysis on the conditions expected at each device and a global consensus on the temperature configuration. We find this as an interesting topic that requires further investigation.
6 Conclusion
This research introduces a novel training approach for federated learning systems, FLex&\&&Chill, aimed at enhancing convergence and improving model performance by exploiting a temperature scaling method termed logit chilling in federated learning local model training operations. Specifically, our work revolves around the implications of employing fractional temperature values ($10<T<$ 1) during the model training process. Our findings, derived from an extensive set of evaluations using three datasets and three widely applied neural networks for federated learning, demonstrate that exploiting the FLex&\&&Chill approach in the client-side model training process holds the potential to expedite model convergence by inducing more pronounced changes despite a limited number of local samples. Consequently, we show that by doing so, we can also achieve an overall higher inference accuracy. While the impact of lower temperature vary with respect to the neural network architecture and dataset, our evaluations reveal up to 6.00$\times$ improvement in federated learning model convergence speed and 3.37% improvement in inference accuracy. We see this work as a first step into devising a more fine-tuned (and optimal) training mechanism for federated learning systems.

### v2

@@
-FLex&Chill: Improving Local Federated Learning Training with Logit Chilling
+Improving Local Training in Federated Learning via Temperature Scaling
 Abstract
-Federated learning are inherently hampered by data heterogeneity: non-iid distributed training data over local clients.
+Federated learning is inherently hampered by data heterogeneity: non-i.i.d.
+training data over local clients.
 We propose a novel model training approach for federated learning, FLex&\&&Chill, which exploits the Logit Chilling method.
-Through extensive evaluations, we demonstrate that, in the presence of non-iid data characteristics inherent in federated learning systems, this approach can expedite model convergence and improve inference accuracy.
+Through extensive evaluations, we demonstrate that, in the presence of non-i.i.d.
+data characteristics inherent in federated learning systems, this approach can expedite model convergence and improve inference accuracy.
 Quantitatively, from our experiments, we observe up to 6$\times$ improvement in the global federated learning model convergence time, and up to 3.37% improvement in inference accuracy.
@@
 (2017).
-This approach provides opportunities for communication and computation efficiency during training by sharing model parameters/gradients learned independently from local leaning processes while suppressing raw data transmissions.
-Furthermore, since training operations take place in a de-centralized manner, federated learning opens the possibility to parallelize the model training process via the assistance of local client devices.
-It presents an attractive solution for handling and exploiting the vast amounts of data generated by many mobile and Internet of Things (IoT) device entities consisting the overall system Park et al.
+This approach provides opportunities for communication and computation efficiency during training by sharing model parameters/gradients learned independently from local learning processes while suppressing raw data transmissions.
+Furthermore, since training operations take place in a decentralized manner, federated learning opens the possibility of parallelizing the model training process via the assistance of local client devices.
+It presents an attractive solution for handling and exploiting the vast amounts of data generated by many mobile and Internet of Things (IoT) device entities consisting of the overall system Park et al.
 (2023a).
@@
 A fundamental challenge of federated learning arises from the disparity of training data used at each client device.
-Federated stochastic optimization encounters difficulties in determining optimal parameters, leading to increased communication overhead between local devices and the central server, as well as the need for additional training on (typically resource limited) local devices Zhao et al.
+Federated stochastic optimization encounters difficulties in determining optimal parameters, leading to increased communication overhead between local devices and the central server, as well as the need for additional training on (typically resource-limited) local devices Zhao et al.
 (2018).
-To address these challenges, a number of prior research have focused on optimizing the local training operations on federated learning clients or on refining the server-side aggregation process.
+To address these challenges, a number of prior research have focused on optimizing local training operations on federated learning clients or on refining the server-side aggregation process.
 For instance, FedProx Li et al.
-(2020) controls the number of iterations for each local device, aiming to train models resilient to challenges posed by non-independent and non-iid data environments.
+(2020) controls the number of local device iterations, aiming to train models resilient to challenges posed by non-independent and non-i.i.d.
+data environments.
 SCAFFOLD Karimireddy et al.
 (2020) achieves expedited convergence and improved model accuracy McMahan et al.
-(2017) by introducing a correction term during the model aggregation phase to balance the influence of each client.
-These operations alleviate the problems posed by the non-iid environment, a common dataset configuration for federated learning clients.
+(2017) by introducing a correction term in the model aggregation phase to balance the influence of each client.
+These operations alleviate challenges posed by the non-i.i.d.
+environment, a common dataset configuration for federated learning clients.
 This paper proposes a novel model training approach for federated learning, FLex&\&&Chill (Federated Learning EXploiting logit Chilling), that targets to boost the convergence and model performance via temperature scaling, namely logit chilling, during the local model training operations.
-Note that a typical neural network generates a probability ($q_{i}$) for each class from its neural network logits ($z_{i}$) by applying a softmax operation (Eq.
+Note that a typical neural network generates a probability ($p_{i}$) for each class from its neural network logits ($z_{i}$) by applying a softmax operation (Eq.
 1) with the temperature $T$ conventionally set to 1.
@@
 (2017).
-$\xi(z,i,T)=q_{i}=\frac{exp(z_{i}/T)}{\sum_{j}exp(z_{j}/T)}$ (1)
-When employing a temperature $T$ higher than 1, the probability distribution across classes naturally undergoes smoothing due to the narrowed interval between the logits Guo et al.
+$\xi(z,i,T)=p_{i}=\frac{exp(z_{i}/T)}{\sum_{j}exp(z_{j}/T)}$ (1)
+When employing a temperature $T~{}>$ 1, the probability distribution across classes undergoes smoothing due to the narrowed interval between logits Guo et al.
 (2017), making such an approach suitable for applications such as knowledge distillation.
 Notably, despite altering the probabilities, the order of labels’ probability remains unaffected with respect to changes in $T$.
-While these works successfully explore the impact of varying temperatures for neural network operations, they have mainly focused on applying high temperatures after the model training process, mostly for the inference operations.
-Raising $T$ to a higher value reduces the confidence of the training operations.
-This is especially useful to tame DNNs which are typically overconfident.
-On the other hand, in federated learning, where disparate training data is common, increasing the model confidence of each individually trained client model (using a lower $T$) prior to their parameter aggregation can be a more effective approach.
-Our investigation involves measuring both the gradient flow and the data position shift within the representation space to provide guidelines in this unexplored territory.
+While these works successfully explore the impact of varying temperatures for neural network operations, they mainly focus on applying high temperatures after the model training process, mostly for inference operations.
+Since increasing $T$ reduces the training confidence, such schemes are effective in taming overconfident DNNs.
+On the other hand, in federated learning, where disparate training data is common, increasing model confidence of each individually trained client model (using small $T$) prior to parameter aggregation can be a more effective approach.
+Our investigation involves measuring both the gradient flow and data position shift within the representation space to provide guidelines in this unexplored territory.
 Specifically, this work targets to show that utilizing fractional temperature values (i.e., $T\in(0,1)$) during the training process enables effective gradient propagation towards the input layer when applied to federated learning data distribution scenarios.
-Consequently, we observed that lower temperatures facilitate more efficient data updates within the representation space when dealing with non-iid datasets, a common characteristic of federated learning systems.
-From our evaluations with three datasets and three baseline federated learning models, we show that these effects accelerate the federated learning model convergence time by up to $\mathbf{6.00\times}$ and improves the inference accuracy by up to $\mathbf{3.37\%}$, despite potential concerns on model instability when training with with low temperatures.
+Consequently, we observed that lower temperatures facilitate more efficient data updates within the representation space when dealing with non-i.i.d.
+datasets, a common characteristic of federated learning systems.
+From our evaluations with three datasets and three baseline federated learning models, we show that these effects accelerate the federated learning model convergence time by up to $\mathbf{6.00\times}$ and improve the inference accuracy by up to $\mathbf{3.37\%}$, despite potential concerns on model instability when training with with low temperatures.
 We summarize the contributions of this work in three-fold as the following:
@@
 2 Background and Related Work
-2.0.1 Temperature scaling
-In machine learning, the output probability distribution generated by a model serves diverse purposes.
+2.1 Temperature scaling
+In machine learning, a model’s output probability distribution serves diverse purposes.
 Primarily, classification models offer probabilities for each class, showcasing the model’s confidence in its predictions.
-This aids in assessing how certain the model is in its predictions (i.e., classification outputs) and allows consideration of the likelihood of each class prior to making a final classification decision.
-Moreover, this distribution offered by the model helps in understanding the uncertainty of a model on its decisions Pearce et al.
-(2021); Park et al.
-(2023b).
-In neural networks, computing probabilities from logits often involves employing softmax operations, as shown in Eq.
+This certainty measure allows consideration of the likelihood of each class prior to making a final classification decision.
+Moreover, this distribution offered by the model helps in understanding the uncertainty of a model on its decisions Park et al.
+(2023b); Pearce et al.
+(2021).
+For neural networks, computing probabilities from logits often involves employing softmax operations, as in Eq.
 1.
-Here, the temperature parameter ($T>0$) controls the probability distribution’s concentration towards one class, and is used to manage how the categorical probability is dispersed over the possible classes Jang et al.
+Here, the temperature parameter ($T>0$) controls the probability distribution’s concentration towards one class, and manages how the categorical probability is dispersed over possible classes Jang et al.
 (2016).
-Notably, this categorical probability aligns with the concept of entropy in information theory, offering a perspective where the temperature parameter governs the level of disorder or uncertainty within the distribution.
-Specifically, when $T\in(1,\infty)$ is applied, the probability distribution tends to be smoother due to increased $T$.
+Notably, this categorical probability aligns with the concept of entropy in information theory, offering a perspective where $T$ governs the level of disorder or uncertainty within the distribution.
+Specifically, with $T\in(1,\infty)$, the probability distribution tends to be smoother as $T$ increases.
 This arises from the diminished differences between logit values, enhancing the distribution’s entropy.
-In contrast, with $T\in(0,1)$ within the softmax function, the probability distribution becomes sharper, since lower temperatures amplify the differences between logits, effectively polarizing the distribution and concentrating probabilities closer to the maximum logit value.
-Consequently, the model becomes more confident in its predictions, as it allocates higher probabilities to fewer classes, thereby reducing uncertainty in the distribution.
+In contrast, $T\in(0,1)$ sharpens the probability distribution, since a lower $T$ amplifies inter-logit differences, effectively polarizing the distribution and concentrating probabilities closer to the maximum logit value.
+Consequently, model predictions are made more confidently, by allocating higher probabilities to fewer classes, thereby reducing uncertainty in the distribution.
 Therefore, temperature scaling operates as a means to adjust the probability distribution within a neural network-based system.
@@
 This technique allows for the control of variability levels in generated samples, enabling adjustments to generated output diversity.
-Temperature scaling is also used in in model calibration Guo et al.
+Temperature scaling is also used in model calibration Guo et al.
 (2017).
 Applying temperature adjustments to the model’s output probabilities, typically through a scaling factor $T$, aids in aligning these probabilities more accurately with the true likelihoods or uncertainties present in the data.
-Likewise, previous work have focused on applying high temperature values for various purposes, mostly for the purpose of conservatively better-understanding the full dataset without bias.
-2.0.2 Federated learning
+Likewise, previous work has focused on applying high-temperature values for various purposes, mostly for the purpose of conservatively better understanding the full dataset without bias.
+2.2 Federated learning
 In recent years, federated learning has emerged as a promising paradigm in machine learning, particularly in scenarios where data privacy, security, and the decentralization of model training are core concerns.
@@
 Despite its benefits, federated learning poses several challenges.
-A body of work have tackled the issue of communication overhead stemming from the exchange of frequent model parameter updates.
-A number of previously proposed work have proposed schemes to effectively abstract the model parameters in various forms to relieve the communication and networking overhead Park et al.
+A body of work has tackled the issue of communication overhead stemming from the exchange of frequent model parameter updates.
+A number of previously proposed works have proposed schemes to effectively abstract the model parameters in various forms to relieve the communication and networking overhead Park et al.
 (2023a).
-Furthermore, federated learning systems are prone to facing issues rising from data heterogeneity caused from the diversity or differences in the characteristics, distributions, or formats of data across federated learning clients.
-Previous work have attempted to address this challenge by addressing relevant issues such as variations in feature/input variations Yu et al.
+Furthermore, federated learning systems are prone to facing issues arising from data heterogeneity caused by the diversity or differences in the characteristics, distributions, or formats of data across federated learning clients.
+Previous works have attempted to address this challenge by addressing relevant issues such as variations in feature/input variations Yu et al.
 (2021) and spatio-temporal variations Liu et al.
 (2020).
-Figure 1: Distribution of gradient norm observed at the input layer for correctly and incorrectly inferenced samples with varying temperature values.
-We observe noticeable improvements in gradient flow to the when training with lower temperatures.
 3 Design of FLex&Chill
@@
 3.1 Impact of Training with Lower Temperatures
-As aforementioned, conventionally, temperature scaling has primarily been focused on calibrating the probability distribution of a model’s output after the model training phase, with limited exploration into its application during the training process itself.
+As aforementioned, conventionally, temperature scaling has primarily been focused on calibrating the probability distribution of a model’s output after the model training phase, with a limited exploration into its application during the training process itself.
 In this section, we present empirical results on three initial experiments specifically crafted to showcase the implications of employing temperature scaling throughout the training phase operations in different dimensions.
@@
 (2017) as the baseline neural network architecture.
+For experimental details, see APPENDIX B
 3.1.1 Gradient flow patterns
-Our first experiment focuses on assessing the norm of computed gradients within the input layer for various temperature values ($T$), and we present sample distribution plot of how the gradients flow for 50K training samples in Figure 1.
-Note that here we separate the plots for the correctly inferenced samples (in the foward pass) and the incorrect ones to observe their differences.
+Figure 1: Distribution of gradient norm at input layer for correctly/incorrectly inferred samples with varying training temperatures.
+Figure 2: Distributions depicting differences between distances to the decision boundary before and after model updates for varying training temperatures.
+Notice that lower temperatures show a noticeable shift in estimations’ positions on the representation space, suggesting their aggressiveness in modifying the model even with a small number of training samples.
+Our first experiment focuses on assessing the norm of computed gradients within the input layer for various temperature values ($T$), and we present a sample distribution plot of how the gradients flow for 50K training samples in Figure 2.
+Note that here we separate the plots for the correctly inferred samples (in the forward pass) and the incorrect ones to observe their differences.
 The propagation of gradients to the input layer serves as an indicative metric for the model’s learning efficacy.
@@
 Eq.
-2 presents the gradient of cross-entropy for the softmax activation with varying temperature $T$ and hard target, one-hot encoded labels.
-For brevity, we present the final derivation result and leave details in the Appendix.
+2 presents the gradient of cross-entropy for the softmax activation ($\mathcal{L}$) with varying temperature $T$ and hard target, one-hot encoded labels.
+For brevity, we present the final derivation result and leave details in the Appendix (see Appendix A).
 Since the gradient is multiplied by $\frac{1}{T}>1$, for low temperatures $T<1$, we can expect a gradient boost.
-$\frac{\partial\xi}{\partial{z_{i}}}=\frac{1}{T}(p_{i}-y_{i})=\frac{1}{T}(\frac{e^{z_{i}/T}}{\sum_{j}e^{z_{j}/T}}-y_{i})$ (2)
-Furthermore, from the perspective of the loss term $(p_{i}-y_{i})$, when the model yields an incorrect decision for a given sample, the loss term significantly increases as $p_{i}\xrightarrow{}0$ (more details later in this section).
-The double-folded effect accounts for substantial mode shift with thick and long tails in the gradient distribution for lower training temperatures.
+$\frac{\partial\mathcal{L}}{\partial{z_{i}}}=\frac{1}{T}(p_{i}-y_{i})=\frac{1}{T}(\frac{e^{z_{i}/T}}{\sum_{j}e^{z_{j}/T}}-y_{i})$ (2)
+Furthermore, from the perspective of the loss term $(p_{i}-y_{i})$, when the model yields an incorrect decision, the loss term significantly increases as $p_{i}\xrightarrow{}0$.
+The double-folded effect accounts for a substantial mode shift with thick and long tails in the gradient distribution for lower temperatures.
 3.1.2 Data position shift in the representation space
@@
 For this experiment, we configure an initial neural network and perform inference for 1,000 samples to identify the results’ positions on the representation space.
-Among the 1,000, we select the ones that were misclassified, and recompute their positions after performing one training operation using that specific data point and compute the difference in positions on the representation space with respect to the decision boundary.
-By doing so, we can observe how drastically one training operation affects the next inference output with respect to different temperatures used in the training phase.
-Figure 2: Distributions depicting differences between distances to the decision boundary before and after model updates for varying training temperatures.
-Notice that lower temperatures show noticeable shift in estimations’ positions on the representation space, suggesting their aggressiveness in modifying the model even with a small number of training samples.
+Among the 1,000, we select the ones that were misclassified, recompute their positions after performing one training operation using that specific data point, and compute the difference in positions on the representation space concerning the decision boundary.
+By doing so, we can observe how drastically one training operation affects the next inference output to different temperatures used in the training phase.
 Figure 2 plots our results in the form of distributions depicting differences between distances to the decision boundary before and after model updates.
@@
 This suggests that using a small $T$ at the softmax operations can actively steer the model towards (potentially) substantial optimization to better align with the training data.
-3.1.3 Impact of temperature on individual samples
-Figure 3: Probability distribution for single training sample (class index 4) with different training temperatures.
-A low $T$ hallucinates the model’s probability distribution, causing the neural network to be trained more aggressively even with a small number of samples.
-We now examine the impact of different degrees of temperatures on a single sample prediction.
-Specifically, we take a look at the probability distribution of prediction for a single sample with varying temperature during training process.
-In Figure 3 we present the the resulting probability distribution for a single training sample (with ground truth index of 4).
-As the plots show, a higher temperature, $T$=4 in this case, leads to a widely spread probability distribution, and the neural network tends to less confident in prediction.
-On the other hand, with lower temperatures the information embedded in the data sample becomes more concentrated to the predicted label.
-If the prediction is incorrect, this increased prediction on the wrong label results in reduced probability on the correct label and increases the loss.
-This leads to more aggressive changes despite training from a single sample.
-In federated learning, clients will train themselves with a small amount of local data (much smaller than typical centralized training).
-For the models trained at the clients to hold value at the server during the aggregation process, it is important that the local data characteristics are effectively embedded in its model parameters despite the small local dataset size.
-The aggressive behavior when applying low training temperatures that we see in Figure 3 and throughout our initial experimental results serve as evidence that the core information from each training element can be effectively extracted and learned at the neural network when low temperatures are applied in the federated learning training process.
+3.1.3 Visualization in the representation space
+To examine the effect of applying a low temperature during the model training phase, we conducted a simple example that classifies 2-dimensional data via logistic regression.
+We split the data into three different subsets and assign them as three devices with non-i.i.d data.
+The Softmax layer was applied at different temperatures ($T$=1.0, 0.5) during the local training phase and average model weights for aggregation after 10 local updates.
+Figure 3 illustrates the classification boundaries and samples distributed in the representation space for the three clients and the aggregated model.
+The results suggest that applying $T$ = 0.5 effectively reflects the local dataset characteristics compared to $T$ = 1.0.
+Specifically, as an example, for clients 0, 1 and the aggregate model, using $T$ = 1.0 fails to make a clear distinction between the red and light green samples; whereas, when applying $T$ = 0.5, this distinction becomes pronounced.
+Figure 3: Example of data points in the 2D representation space with their respective classification boundaries for different federated learning clients with varying training temperatures.
 3.2 FLex&\&&Chill and Logit Chilling
-Based on such empirical observations, we hypothesise that training neural networks with lower temperatures can benefit the model convergence and performance for systems exploiting with non-iid data, which is common for federated learning scenarios.
-In FLex&\&&Chill we term this as logit chilling, as a we ‘chill down’ the temperature at the softmax operations, thereby altering the logit computation derived from the neural network during the training process.
-The focus of a FLex&\&&Chill-based federated learning system is to exploit logit chilling to train a client-side model with low temperature values.
+Based on such empirical observations, we hypothesize that training neural networks with lower temperatures can benefit the model convergence and performance for systems exploiting non-i.i.d.
+data, which is common for federated learning scenarios.
+In FLex&\&&Chill we term this as logit chilling, as we ‘chill down’ the temperature at the softmax operations, thereby altering the logit computation derived from the neural network during the training process.
+The focus of a FLex&\&&Chill-based federated learning system is to exploit logit chilling to train a client-side model with low-temperature values.
 As we observed through the studies above and as our evaluations will show, this has an effect of effectively exploiting crucial information from the training samples to actively alter the model towards the new knowledge.
-The lower temperatures used by Logit Chilling directly influences the local training process on each federated learning client.
+The lower temperatures used by Logit Chilling directly influence the local training process of each federated learning client.
 Notably, the seamless integration of Logit Chilling and FLex&\&&Chill with existing federated learning frameworks is a distinct advantage, allowing for a straightforward implementation within already deployed frameworks.
 Furthermore, we note that FLex&\&&Chill can be used orthogonality with other efforts that accelerate federated learning training operations such as FedProx and SCAFFOLD.
+For a detailed explanation and experimental result about the orthogonality of FLex&\&&Chill, please refer to Appendix  C.
+4 Experiment
+We evaluate the performance of FLex&\&&Chill using three datasets: (i) FEMNIST from the LEAF database Caldas et al.
+(2018a), including the Extended MNIST handwriting data Cohen et al.
+(2017) , (ii) the CIFAR10 dataset, and (iii) the CIFAR100 dataset Krizhevsky et al.
+(2009).
+As metrics, we use the average accuracy and training loss observed at the clients, selected to observe both the performance gain and the effectiveness of FLex&\&&Chill in its training operations.
+Furthermore, we also present the number of federated learning rounds needed to achieve a pre-defined target accuracy.
+(a) FEMNIST-DNN
+(b) CIFAR10-CNN
+(c) CIFAR100-ResNet18
+Figure 4: Average federated learning client-side model accuracy for different temperatures used in training.
+Overall, lower temperatures show higher average model accuracy for federated learning systems.
+4.1 Experiment Setup
+Dataset. We detail the datasets used in our work as follows.
+  • FEMNIST dataset: The FEMNIST dataset comprises of 10 handwritten digits (0-9) and 52 characters (26 lowercase and 26 uppercase) images contributed by 712 users, totaling 157,132 samples.
+We selected a subset of 36 users, each with diverse amounts of local data to account for the non-i.i.d.
+environment.
+We followed the implementations provided by the original work Caldas et al.
+(2018b).
+  • CIFAR10 and CIFAR100 dataset: We use baseline CIFAR10 and CIFAR100 datasets Krizhevsky et al. (2009) and manually partition data over 50 federated learning users, with each holding 1,000 samples.
+Model.
+For each dataset, we utilized distinct model architectures, taking into account the specific challenges and characteristics inherent to each dataset.
+For FEMNIST, an MLP model featuring four linear layers was employed.
+We term this combination as “FEMNIST-DNN” in the rest of our evaluations.
+In the case of the CIFAR10 and CIFAR100 datasets, experiments were conducted using a two-layer CNN model McMahan et al.
+(2017) and a ResNet18 He et al.
+(2016), respectively.
+These two cases are denoted as “CIFAR10-CNN” and “CIFAR100-ResNet18.”
+Experiment Details.
+Each case underwent local training for 300 federated training rounds, utilizing FedAvg McMahan et al.
+(2017) as the baseline federated learning framework.
+FedAvg is a widely used framework for federated learning.
+In each round, the server selects 10 participating clients, employing Stochastic Gradient Descent (SGD) as the optimizer with a learning rate of 0.001 Ruder (2016).
+Each client trained its model for 10 local epochs, employing a batch size of 16, and utilized cross-entropy loss as the loss function.
+Unless specified otherwise, the experiment was conducted using the same hyperparameters as mentioned in this section.
+4.2 Average Client-side Accuracy
+We report the average client-side model accuracy obtained for models trained with different temperatures in Figure 4.
+For all three datasets (and the three respective models used for each test case) we see an increasing accuracy trend observed for lower temperatures.
+Note that for the FEMNIST-DNN case with $T$=0.05, we noticed that the accuracy converges after 120 rounds; thus, perform an early exit for this case.
+One interesting aspect we can notice is that the accuracy for the $T$=0.05 case, which is the lowest temperature we tested for, does not show the best performance for the CIFAR100-ResNet18 experiment.
+While the $T$=0.05 case shows the best performance of FEMNIST-DNN and CIFAR10-CNN, for CIFAR100-ResNet18, we see an accuracy worse than the $T$=0.5 case.
+This suggests that training with the lowest possible temperature value is not always ideal, rather this result highlights the importance of identifying a proper training temperature as we will further discuss in Section 5.
+A closer zoom into the plots suggests that the best-performing configuration can change over the training round and can vary for different datasets, but overall, we can notice that exploiting a temperature value lower than “1” in the training process can help improve the classification accuracy.
+The average accuracy for all federated learning clients after 300 training rounds presented in Table 1 confirm our observations, indicating that with $T$=0.05, the CIFAR10 case shows a 3.37% improvement in accuracy compared to the $T$=1 case.
+Even for other cases, the use of a low-temperature value shows promising results in terms of accuracy.
+We emphasize once more that this does not mean that the accuracy will always perform better at the lowest temperature.
+Rather, our results suggest the need to carefully explore the fractional temperature space in the federated learning training process.
+We note that the performance for model performance and convergence speed tends to be similar regardless of federated learning hyperparameters (see Appendix D).
+4.3 Federated Learning Convergence Analysis
+FEMNIST-DNN CIFAR10-CNN CIFAR100-ResNet18 Temperature ($T$) Accuracy (%) Loss Num.
+of Round Accuracy (%) Loss Num.
+of Round Accuracy (%) Loss Num.
+of Round $T=4.0$ 73.60 (-5.06) 0.893 299 (0.44×) 37.58 (-3.76) 1.724 299 (0.84×) 28.05 (-0.20) 2.984 288 (0.69×) $T=2.0$ 77.20 (-1.46) 0.749 205 (0.64×) 40.35 (-0.99) 1.638 270 (0.93×) 27.89 (-0.36) 2.918 288 (0.69×) $T=1.0$ 78.66 (0.00) 0.684 132 (1.00×) 41.34 (0.00) 1.594 251 (1.00×) 28.25 (0.00) 2.918 200 (1×) $T=0.5$ 79.78 (+1.13) 0.654 88 (1.50×) 41.52 (+0.18) 1.551 161 (1.56×) 29.77 (+1.52) 2.875 178 (1.12×) $T=0.25$ 79.50 (+0.85) 0.650 57 (2.32×) 43.41 (+2.07) 1.532 132 (1.90×) 30.63 (+2.38) 2.997 152 (1.32×) $T=0.05$ 80.86 (+2.20) 0.643 22 (6.00×) 44.71 (+3.37) 1.528 106 (2.37×) 27.05 (-2.72) 4.228 239 (0.74×)
+Table 1: Average classification accuracy, model training loss, and number of federated learning training rounds needed to achieve the maximum accuracy of the $T$=4 cases for each dataset-model combination.
+Finally, in the last columns for each test case in Table 1 we present the number of federated learning training rounds needed to achieve a preset target accuracy for different temperature values used for model training.
+This number provides an understanding of how efficient the local training processes are when different temperature values are applied.
+Specifically, in this experiment, we configure the target accuracy to be the maximum accuracy achieved within 300 training rounds with $T$=4.
+For example, for the FEMNIST-DNN and CIFAR10-CNN cases, the maximum accuracy for $T$=4 was achieved at the 299th round, while CIFAR100-ResNet18 achieved such accuracy after 288 rounds.
+From the results in Table 1 we can notice that with $T$=0.05, the FEMNIST-DNN case achieves the target accuracy after only 22 training rounds.
+We see similar trends of low temperatures positively affecting the training efficiency throughout our results.
+This is from the general aggressiveness of training behaviors when applying lower temperature training.
+Given a single training sample, while high temperature-based training conservatively absorbs the available information, lower temperatures are more progressive in making model changes.
+Given the small datasets used for local training and the need to effectively deliver its knowledge to the server for aggregation, we see FLex&\&&Chill as an effective way to train client-side models in the federated learning process.
+4.4 Feature Space Similarity Analysis
+Figure 5: Test accuracy of the global model and feature space similarity in each layer between reference model trained with i.i.d.
+dataset and the global model trained with non-i.i.d.
+dataset with different temperatures applied during the local training.
+(a) $T$=0.05
+(b) $T$=1.0
+Figure 6: Client-wise feature space similarity matrix measured in penultimate layers by using CKA and inference accuracy in different temperatures (0.05, 1.0) and federated rounds (1, 5, 50, 100).
+(INIT: Initial dispatched model in the given round, AGG: Aggregated global model, REF: Reference model, ACC: Accuracy)
+To better understand the mechanism and rationale behind the performance of FLex&\&&Chill, we analyze the similarity of the feature space between global (aggregated) models trained with different temperatures against a reference model trained in an i.i.d.
+data configuration.
+Specifically, we employ the Centered Kernel Alignment (CKA) Kornblith et al.
+(2019) for computing the similarity of feature spaces, a widely used approach in evaluating model divergence in federated learning Collins et al.
+(2022); Luo et al.
+(2021).
+Note that a limitation of CKA measures is that they only report network similarity without considering training quality.
+Thus, we use a well-trained reference model from an i.i.d.
+environment to observe CKA variations of the penultimate and classification layers between the global models and the reference model.
+We note that the reference model is trained with $T$=1, while comparison models (e.g., global models from federated learning) are trained in non-i.i.d.
+environments with different $T$.
+In the experiments, we used a 2-layered CNN and the CIFAR-10 dataset as in Section 4.1, with the dataset distributed among 10 clients using a Dirichlet distribution Hsu et al.
+(2019).
+All clients participated in the federated training process.
+For detailed discussions on the experimental setup and CKA, please refer to Appendix E.
+As the accuracy and CKA plots for different layers in Figure 5 show, introducing the non-i.i.d.
+environment itself drops the model accuracy (73.86% $\xrightarrow{}$ 58.19%) at $T$=1, with the CKA of the penultimate/classification layers between the two models remaining at 0.7644/0.8214 after 300 rounds.
+Nevertheless, with $T$=0.05, despite the dataset disparity, the global model achieved a comparable accuracy of 72.76% with relatively higher CKA values (0.8898/0.9026) in both the penultimate and classification layers.
+We point out that improved gradient flows via applying low temperatures help models to learn similar local features (i.e., layers 1, 2), affecting the following layers (i.e., layers 3, 4) in the early stage of the learning process.
+As a result, applying low temperatures during local training effectively improves the overall accuracy performance in a non-i.i.d.
+data distribution setting and helps models learn similar feature spaces compared to the ideal i.i.d.
+data configuration.
+We now target to better understand how local model updates progress with different $T$.
+For brevity of the paper we only report the CKA at the penultimate layer, and plot all other cases in Appendix E.
+Specifically, Figure 6 plots the model accuracy (ACC) and measured CKA at the penultimate layer of a client’s local model with (i) the initially distributed model (INIT), (ii) the aggregated model (AGG), (iii) the ideal reference model trained with i.i.d data distribution (REF), and (iv) other client models while $T$=0.05 and 1.0.
+In Round 1, clients’ local models trained with $T$=1 show high feature space similarity against each other while failing to learn the features of the ideal reference model.
+With $T$=0.05, the CKA between local models are relatively low while showing high similarity with the reference model.
+After five rounds (and beyond), local models trained with both $T$=1.0 and $T$=0.05 showed high similarity amongst themselves and the initial/aggregated models, but the $T$=0.05 case continuously shows higher similarity with REF and reports higher accuracy.
+Hence, we argue that adapting low temperatures allows local models to learn effective and diverse features to quickly converge the federated learning process.
+4.5 Impact of Dataset Dispersion and Federated Learning on FLex&\&&Chill
+We now examine the impact of FLex&\&&Chill on different model training configurations.
+Specifically, using the CIFAR10-CNN test case, we examine two additional configurations: (i) centralized learning with i.i.d.
+data and (ii) federated learning with i.i.d.
+data.
+Note that the federated learning network is identical to our previous experiments except for the dataset distribution and was tested for 300 rounds.
+By comparing these results with the non-i.i.d.
+federated learning results, we target to understand the impact of FLex&\&&Chill on federated learning and dataset disparity in greater detail.
+As Table 2 shows, applying FLex&\&&Chill, thus a lower $T$ does not induce a positive impact on the performance of centralized learning scenarios.
+In fact, for centralized learning, controlling $T$ for the training process does not show a noticeable impact.
+We see the same phenomena for the federated learning system with i.i.d.
+samples as well, suggesting that the performance enhancements we see in Table 1 are not an effect of the federated learning process itself.
+Centralized - i.i.d.
+Federated - i.i.d.
+Federated - non-i.i.d.
+$T$ Acc.
+(%) Loss Acc.
+(%) Loss Acc.
+(%) Loss 4.0 58.05 1.152 48.22 1.473 37.58 1.724 2.0 63.27 1.041 51.90 1.403 40.35 1.638 1.0 65.18 1.044 54.66 1.325 41.34 1.594 0.5 64.19 1.030 53.19 1.320 41.52 1.551 0.25 63.88 1.092 54.64 1.265 43.41 1.532 0.05 60.98 1.227 47.60 1.701 44.71 1.528
+Table 2: Average inference accuracy and model training loss for CIFAR10-CNN with different configurations.
+The impact of FLex&\&&Chill is prominent for federated learning scenarios with non-i.i.d.
+datasets, whereas its impact is less than elsewise.
+We do point out that holding i.i.d.
+datasets in federated learning will show overall higher performance compared to non-i.i.d.
+configurations (regardless from $T$), but is not a practical case to assume when deploying federated learning systems.
+Rather, as we discussed, the impact of exploiting low training temperatures is beneficial when dealing with non-i.i.d.
+datasets.
+4.6 Model stability of low-temperature training during federated learning aggregation
+Figure 7: Accuracy drop observed before and after server-side model aggregation. The aggregation process of client models trained with lower temperatures better retains the individual knowledge.
+Figure 7 plots the changes in accuracy observed for both participating and non-participating clients at each federated learning round.
+Specifically, once the model is aggregated, each client computes its accuracy with both the pre-update and the post-update model and computes their difference.
+Note that pre-update models for clients participating in the round would have updated their local models (very recently) with their own data.
+As discussed in Section 3, low temperatures expedite local training even with a small amount of training data.
+Nevertheless, the effect of low-temperature training on the model generalization and aggregation stability of federated learning has not yet been thoroughly studied.
+To inspect the stability of federated learning aggregation for models trained with low temperatures, we measured the performance change of each client-side model before and after the server-side aggregation round.
+Thus, the pre-model performance will be high.
+However, since each model is trained aggressively over heterogeneous data, the aggregation process could puddle up this knowledge and the aggregated model may lose information.
+Ideally, with their local information properly aggregated, the performance drop will be minimal.
+Non-participating clients on the other hand, would hope to have an improved updated model of the federated learning process.
+As Figure 7 shows the models trained with low temperatures retain the knowledge better than the medium and hot models for participating clients (solid plots groups at the bottom).
+In the early rounds, the models with high temperatures show less performance drop, but the fast learning speed of the low-temperature model amortizes the performance drop, which results in faster learning of the aggregated model.
+Furthermore, we can also notice that the aggregation process does not interfere with the performance of non-participating clients (grouped on top), implying that federated learning with low-temperature training preserves the generalization power.
+5 Discussion
+$\bullet$ Identifying the optimal temperature.
+Our study exhibits the effectiveness of exploiting lower temperatures as a way to enhance federated learning performance.
+However, since the softmax layer training temperature can be considered a hyperparameter, optimal temperatures can vary depending on the applications and model architectures.
+Thus, like other hyperparameters used in model training (e.g., batch size, local epoch count), identifying an optimal temperature value is challenging to analytically solve, being beyond the scope of this paper.
+Nevertheless, based on our experiences, we can offer heuristics for selecting temperatures regarding the model capacity and task complexity.
+Large capacity models with relatively easy tasks (e.g., low-resolution data, small number of labels to classify) often exhibit unstable convergence with low temperatures.
+We hypothesize that this phenomenon arises from lower temperatures compelling models to rapidly adapt to local data, which can potentially bias the model towards local data and disrupt the federated learning process.
+Thus, we carefully conclude our heuristics in two-fold.
+(i) Balancing the model capacity and task complexity is crucial to best leverage the temperature scaling during training; (ii) Large capacity models tend to work well with relatively high temperatures since they tend to suffer from instability with a too-low temperature.
+$\bullet$ Dynamic per-sample temperature scaling.
+While our paper focuses on understanding the advantages of maintaining a constant low temperature for the training process, this may not be the ideal approach when targeting maximum performance.
+In reality, varying temperatures can be applied in the training process with respect to their significance Shin et al.
+(2022); Zeng et al.
+(2021).
+That would mean that the training process can exploit lower temperatures for the samples that are “harder” to learn, while higher temperatures are used for those easier to understand.
+Alternatively, a more course-grain approach of dynamically adjusting the temperature on a per-federated learning round-basis can be another meaningful approach to adapt towards changing local data characteristics.
+$\bullet$ Supporting heterogeneous federated learning.
+While many federated learning frameworks assume homogeneous devices, recent work points out that this is not always the case Park and Ko (2023).
+In a federated learning network consisting of devices with heterogeneous computing/network capabilities, the issue of “stragglers”, which take longer to complete local training, becomes important to address.
+Leveraging insights gleaned from our empirical observations, applying lower temperatures for potential stranglers can be an interesting way of uniforming the local training duration.
+This would require a prior analysis on the conditions expected at each device and a global consensus on the temperature configuration.
+We find this as an interesting topic that requires further investigation.
+6 Conclusion
+This research introduces a novel training approach for federated learning systems, FLex&\&&Chill, aimed at enhancing convergence and improving model performance by exploiting a temperature scaling method termed logit chilling in federated learning local model training operations.
+Specifically, our work revolves around the implications of employing fractional temperature values ($T\in(0,1)$) during the model training process.
+Our findings, derived from an extensive set of evaluations using three datasets and three widely applied neural networks for federated learning, demonstrate that exploiting the FLex&\&&Chill approach in the client-side model training process holds the potential to expedite model convergence by inducing more pronounced changes despite a limited number of local samples.
+Consequently, we show that by doing so, we can also achieve an overall higher inference accuracy.
+While the impact of lower temperature varies with respect to the neural network architecture and dataset, our evaluations reveal up to 6.00$\times$ improvement in federated learning model convergence speed and 3.37% improvement in inference accuracy.
+We see this work as a first step into devising a more fine-tuned (and optimal) training mechanism for federated learning systems.
+Appendix A Gradient of softmax with varying temperature
+Proof.
+Let cross-entropy loss for softmax as $\mathcal{L}=-\sum_{j=1}^{C}y_{j}\log{(p_{j})}$, where $C$ is the number of classes and $y_{j}$ and $p_{j}$ corresponds to the hard label (i.e., onehot-encoded label) and probability for given class.
+Here, the probability $p_{j}$ is computed by exploiting the softmax with varying temperatures which can be formalized as follows.
+$p_{j}=\frac{e^{z_{j}/T}}{\sum_{i=1}^{C}{e^{z_{i}/T}}}$
+Here, $T$ is the temperature, and $z_{j}$ is the logit used as the input of the softmax function.
+$\begin{split}\frac{\partial\mathcal{L}}{\partial z_{i}}&=-\sum_{j=1}^{C}\frac{\partial y_{j}\log{(p_{j})}}{\partial z_{i}}\\ &=\sum_{j=1}^{C}y_{j}\frac{\partial\log{(p_{j})}}{\partial z_{i}}\\ &=-\sum_{j=1}^{C}y_{j}\frac{\partial\log{(p_{j})}}{\partial z_{i}}\\ &=-\sum_{j=1}^{C}y_{j}\frac{1}{p_{j}}\frac{\partial p_{j}}{\partial z_{i}}\\ &=-\frac{y_{i}}{p_{i}}\frac{\partial p_{i}}{\partial z_{i}}-\sum_{j\neq i}^{C}\frac{y_{j}}{p_{j}}\frac{\partial p_{j}}{\partial z_{i}}\\ &=-\frac{y_{i}}{p_{i}}\frac{1}{T}p_{i}(1-p_{i})-\sum_{j\neq i}^{C}\frac{y_{j}}{p_{j}}(-p_{j}p_{i})\\ &=-\frac{y_{i}}{T}+\frac{1}{T}y_{i}p_{i}+\frac{1}{T}\sum_{j\neq i}^{C}y_{j}p_{i}\\ &=-\frac{y_{i}}{T}+\frac{1}{T}\sum_{j=1}^{C}y_{j}p_{i}\\ &=-\frac{y_{i}}{T}+\frac{1}{T}p_{i}\sum_{j=1}^{C}y_{j}\\ &=\frac{1}{T}(p_{i}-y_{i})\\ \therefore\frac{\partial\mathcal{L}}{\partial z_{i}}&=\frac{1}{T}(p_{i}-y_{i})\\ &=\frac{1}{T}(\frac{e^{z_{i}/T}}{\sum_{j}e^{z_{j}/T}}-y_{i})\end{split}$
+Appendix B Experimental Details
+B.1 Model Architecture
+Layer Details Repetition FC 1 Linear(784, 512) ReLU() ×$\times 1$ FC 2 Linear(512, 256) ReLU() ×$\times 1$ FC 3 Linear(256, 128) ReLU() ×$\times 1$ FC 4 Linear(256, 62) ×$\times 1$
+Table 3: Model architectures details of DNN used to train the FEMNIST dataset
+Layer Details Repetition Conv 1 Conv2d(3, 10, k=(5, 5), s=(1, 1)) ReLU() MaxPool2d(k=(2, 2)) ×$\times 1$ Conv 2 Conv2d(10, 20, k=(5, 5),s=(1, 1)) ReLU() MaxPool2d(k=(2, 2)) ×$\times 1$ FC 1 Linear(500, 256) ReLU() ×$\times 1$ FC 2 Linear(256, 10) ×$\times 1$
+Table 4: Model architectures details of 2-layered CNN used to train the CIFAR10 dataset
+For the ResNet18 model used to train the CIFAR100 dataset, we exploited PyTorch implementation Paszke et al.
+[2019] by only replacing the final classification head from Linear(512, 1000) to Linear(512, 100).
+B.2 Hyperparameters
+Hyperparameters FEMNIST-DNN CIFAR10-CNN CIFAR100-ResNet18 CKA Evaluation total clients 36 50 10 communication rounds 300 100 optimizer SGD learning rate 0.001 0.005 local epoch 10 10 participants per round 10 10 batch size 16 32 shard size | $\alpha$ N/A 200 $\alpha$=0.5 loss function Cross Entropy
+Table 5: Hyperparameters used for Section 4. $\alpha$ denotes the importance of the Dirichlet distribution.
+B.3 Data distribution
+(a) FEMNIST
+(b) CIFAR10
+(c) CIFAR100
+Figure 8: Visualization of the distribution of training data used in FEMNIST, CIFAR10-CNN and CIFAR100-ResNet experiments. Best viewed in color.
+B.4 Hardware
+For the experiments, we exploited server with four GeForce RTX 2080 Ti GPUs, Intel(R) Xeon(R) Silver 4210 CPU @ 2.20GHz, and 64GB RAM.
+Appendix C Orthogonality of FLex&\&&Chill
+C.1 FLex&\&&Chill As An Algorithm
 We formulate the operations of FLex&\&&Chill using Algorithm 1.
 Here, typical federated learning operations take place at both the server and clients except for the fact that the server determines a target $T$ at each round and local model training operations are performed with $T$.
+Since FLex&\&&Chill only manipulates the local training process which inherently provides orthogonality to our proposed method, it is extremely easy to integrate with existing federated learning frameworks.
 Algorithm 1 FLex&\&&Chill
@@
 16:    $l(\mathbf{z})=loss(\mathbf{z};label)$
-17:    $w_{t}^{k}\xleftarrow{}w_{t}^{k}-\eta\nabla l(\mathbf{z})$
+17:    $w_{t}^{k}\xleftarrow{}LocalUpdate(w_{t}^{k},l(\mathbf{z}))$
 18:end for
 19:Return $w_{t}^{k}$ to server
-4 Experiment
-We evaluate the performance of FLex&\&&Chill using three datasets: (i) FEMNIST from the LEAF database Caldas et al.
-(2018b), including the Extended MNIST handwriting data Cohen et al.
-(2017) , (ii) the CIFAR10 dataset, and (iii) the CIFAR100 dataset Krizhevsky et al.
-(2009).
-As metrics, we use the average accuracy and training loss observed at the clients, selected to observe both the performance gain and the effectiveness of FLex&\&&Chill in its training operations.
-Furthermore, we also present the number of federated learning rounds needed to achieve a pre-defined target accuracy.
-(a) FEMNIST-DNN
-(b) CIFAR10-CNN
-(c) CIFAR100-ResNet18
-Figure 4: Average federated learning client-side model accuracy for different temperatures used in training.
-Overall, lower temperatures show higher average model accuracy for federated learning systems.
-4.1 Experiment Setup
-Dataset. We detail the datasets used in our work as follows.
-  • FEMNIST dataset: Sourced from the LEAF dataset, the FEMNIST dataset comprises of 10 handwritten digits (0-9) and 52 characters (26 lowercase and 26 upper case) images contributed by 712 users, totaling 157,132 samples.
-We selected a subset of 36 users, each with diverse amounts of local data to account for the non-iid environment.
-We followed the implementations provided by the original work Caldas et al.
-(2018a).
-  • CIFAR10 and CIFAR100 dataset: We use the baseline CIFAR10 and CIFAR100 datasets Krizhevsky et al.
-(2009) and manually partition the data over 50 federated learning users, with each user holding 1,000 samples.
-Model.
-For each dataset, we utilized distinct model architectures, taking into account the specific challenges and characteristics inherent to each dataset.
-For FEMNIST, an MLP model featuring four linear layers was employed.
-We term this combination as “FEMNIST-DNN” in the rest of our evaluations.
-In the case of the CIFAR10 and CIFAR100 datasets, experiments were conducted using a two-layer CNN model McMahan et al.
-(2017) and a ResNet18 He et al.
-(2016), respectively.
-These two cases are denoted as “CIFAR10-CNN” and “CIFAR100-ResNet18.”
-Experiment Details.
-Each case underwent local training for 300 federated training rounds, utilizing FedAvg McMahan et al.
-(2017) as the baseline federated learning framework.
-FedAvg is a widely used framework for federated learning.
-In each round, the server selects 10 participating clients, employing Stochastic Gradient Descent (SGD) as the optimizer with a learning rate of 0.001 Ruder (2016).
-Each client trained its model for 10 local epochs, employing a batch size of 16, and utilized cross-entropy loss as the loss function.
-4.2 Average Client-side Accuracy
-(a) FEMNIST-DNN
-(b) CIFAR10-CNN
-(c) CIFAR100-ResNet18
-Figure 5: Federated learning client-side model training loss observed for different training temperature values.
-We can notice that overall a lower training loss is observed when lower temperature values are used for model training.
-We report the average client-side model accuracy results obtained for models trained with different temperatures in Figure 4.
-For all three datasets (and the three respective models used for each test case) we see an increasing accuracy trend observed for lower temperatures.
-Note that for the FEMNIST-DNN case with $T$=0.05, we noticed that the accuracy converges after 120 rounds; thus, perform an early exit for this case.
-One interesting aspect we can notice is that the accuracy for the $T$=0.05 case, which is the lowest temperature we test for, does not show the best performance for the CIFAR100-ResNet18 experiment.
-While the $T$=0.05 case shows the best performance of FEMNIST-DNN and CIFAR10-CNN, for CIFAR100-ResNet18, we see a accuracy worse than the $T$=0.5 case.
-This suggests that training with the lowest possible temperature value is not always ideal, rather this result highlights the importance of identifying a proper training temperature as we will further discuss in Section 5.
-A closer zoom into the plots suggest that the best performing configuration can change over the training round and can vary for different datasets, but overall, we can notice that exploiting a temperature value lower than “1” in the training process can help improve the classification accuracy.
-The average accuracy for all federated learning clients after 300 training rounds presented in Table 1 confirm our observations, indicating that with $T$=0.05 the CIFAR10 case shows a 3.37% improvement in accuracy compared to the $T$=1 case.
-Even for other cases, the use of a low temperature value show promising results in terms of accuracy.
-We emphasize once more that this does not mean that the accuracy will always perform better at the lowest temperature.
-Rather, our results suggest the need to carefully explore the fractional temperature space in the federated learning training process.
-4.3 Training Loss
-Next we compare the training loss observed at the clients.
-In Figure 5 we present the average the loss observed at each federated learning training round.
-Again, we note that we perform early exit for the FEMNIST-DNN case at 120 rounds when $T$=0.05 due to quick model convergence.
-We can notice there that the use of lower temperatures in the training phase generally helps quickly lower the training loss during the local training phase for federated learning clients.
-At the same time, when observing the CIFAR100-ResNet18 plots, we can see that the loss for $T$=0.05 does not converge well.
-In fact, this serves as a reason behind the lower accuracy reported for CIFAR100-ResNet18.
-From the quantitative values in Table 1 where we plot the average loss observed at all federated learning clients after 300 training rounds (120 for the FEMNIST-DNN @ $T$=0.05), we can make similar observations.
-Especially, for FEMNIST-DNN, we see a six-fold improvement in loss values compared to $T$=1.
-For the CIFAR100-ResNet18 case, however, we notice that the lowest loss convergence is in fact the $T$=0.25 case rather than 0.05.
-This observation confirms the need to properly consider and explore fractional temperature values for client-side model training.
-4.4 Federated Learning Convergence Analysis
-FEMNIST-DNN CIFAR10-CNN CIFAR100-ResNet18 Temperature ($T$) Accuracy (%) Loss Num.
-of Round Accuracy (%) Loss Num.
-of Round Accuracy (%) Loss Num.
-of Round $T=4.0$ 73.60 (-5.06) 0.893 299 (0.44×) 37.58 (-3.76) 1.724 299 (0.84×) 28.05 (-0.20) 2.984 288 (0.69×) $T=2.0$ 77.20 (-1.46) 0.749 205 (0.64×) 40.35 (-0.99) 1.638 270 (0.93×) 27.89 (-0.36) 2.918 288 (0.69×) $T=1.0$ 78.66 (0.00) 0.684 132 (1.00×) 41.34 (0.00) 1.594 251 (1.00×) 28.25 (0.00) 2.918 200 (1×) $T=0.5$ 79.78 (+1.13) 0.654 88 (1.50×) 41.52 (+0.18) 1.551 161 (1.56×) 29.77 (+1.52) 2.875 178 (1.12×) $T=0.25$ 79.50 (+0.85) 0.650 57 (2.32×) 43.41 (+2.07) 1.532 132 (1.90×) 30.63 (+2.38) 2.997 152 (1.32×) $T=0.05$ 80.86 (+2.20) 0.643 22 (6.00×) 44.71 (+3.37) 1.528 106 (2.37×) 27.05 (-2.72) 4.228 239 (0.74×)
-Table 1: Average classification accuracy, model training loss, and number of federated learning training rounds needed to achieve the maximum accuracy of the $T$=4 case for each dataset-model combination.
-Quantitative results suggest the effectiveness of exploiting low temperatures for model training in federated learning both in terms of model accuracy and training efficiency.
-Centralized - iid Federated - iid Federated - non-iid $T$ Acc.
-(%) Loss Acc.
-(%) Loss Acc.
-(%) Loss 4.0 58.05 1.152 48.22 1.473 37.58 1.724 2.0 63.27 1.041 51.90 1.403 40.35 1.638 1.0 65.18 1.044 54.66 1.325 41.34 1.594 0.5 64.19 1.030 53.19 1.320 41.52 1.551 0.25 63.88 1.092 54.64 1.265 43.41 1.532 0.05 60.98 1.227 47.60 1.701 44.71 1.528
-Table 2: Average inference accuracy and model training loss for CIFAR10-CNN with different configurations.
-The impact of FLex&\&&Chill is prominent for federated learning scenarios with non-iid datasets, whereas its impact is less when elsewise.
-Finally, the last columns for each test case in Table 1 we present the number of federated learning training rounds needed to achieve a preset target accuracy for different temperature values using for model training.
-This number provides an understanding of how efficient the local training processes are when different temperature values are applied.
-S
-[…]
-From the results in Table 1 we can notice that with $T$=0.05 the FEMNIST-DNN case achieves the target accuracy after only 22 training rounds.
-We see similar trends of low temperatures positively affecting the training efficiency throughout our results.
-This is from the generally aggressiveness of training behaviors when applying lower temperature training.
-Given a single training sample, while high temperature-based training conservatively absorbs the available information, lower temperatures are more progressive in making model changes.
-Given the small datasets used for local training and the need to effectively deliver its knowledge to the server for aggregation, we see FLex&\&&Chill as an effective way to train client-side models in the federated learning process.
-4.5 Impact of Dataset Dispersion and Federated Learning on FLex&\&&Chill
-We now examine the impact of FLex&\&&Chill on different model training configurations.
-Specifically, using the CIFAR10-CNN test case, we examine two additional configurations: (i) centralized learning with iid data and (ii) federated learning with iid data.
-Note that the federated learning network is identical to our previous experiments except for the dataset distribution and was tested for 300 rounds.
-By comparing these results with the non-iid federated learning results, we target to understand the impact of FLex&\&&Chill on federated learning and dataset disparity in greater detail.
-As Table 2 shows, applying FLex&\&&Chill, thus a lower $T$ does not induce a positive impact on the performance of centralized learning scenarios.
-In fact, for centralized learning, controlling $T$ for the training process does not show a noticeable impact.
-We see the same phenomena for the federated learning system with iid samples as well, suggesting that the performance enhancements we see in Table 1 is not an effect of the federated learning process itself.
-We do point out that holding iid datasets in federated learning will show overall higher performance compared to non-iid configurations (regardless from $T$), but is not a practical case to assume when deploying federated learning systems.
-Rather, as we discussed, the impact of exploiting low training temperatures is beneficial when dealing with non-iid datasets.
-Figure 6: Accuracy drop observed before and after server-side model aggregation. The aggregation process of client models trained with lower temperatures better retain the individual knowledge.
-4.6 Model stability of low temperature training during federated learning aggregation
-As discussed in Section 3, low temperatures can expedite local training even when dealing with a small amount of training data.
-Nevertheless, the effect of low temperature training on the model generalization and aggregation stability of federated learning has not yet been thoroughly studied.
-To inspect the stability of federated learning aggregation for models trained with low temperatures, we measured the the performance change of each client-side model before and after the server-side aggregation round.
-Figure 6 plots the changes in accuracy observed for both participating and non-participating clients at each federated learning round.
-Specifically, once the model is aggregated, each client computes its accuracy with both the pre-update and the post-update model and computes their difference.
-Note that pre-update models for clients participating in the round would have updated their local models (very recently) with their own data.
-Thus, the pre-model performance will be high.
-However, since each model is trained aggressively over heterogeneous data, the aggregation process could puddle up this knowledge and the aggregated model may lose information.
-Ideally, with their local information properly aggregated, the performance drop will be minimal.
-Non-participating clients on the other hand, would hope to have an improved updated model the federated learning process.
-As Figure 6 shows the models trained with low temperatures retain the knowledge better than the medium and hot models for participating clients (solid plots groups at bottom).
-At the early rounds, the models with high temperature show less performance drop, but the fast learning speed of low temperature model amortizes the performance drop, which results in faster learning of the aggregated model.
-Furthermore, we can also notice that the aggregation process does not interfere with the performance of non-participating clients (grouped on top), implying that federated learning with low temperature training preserves the generalization power.
-5 Discussion
-$\bullet$ Identifying the optimal temperature.
-Our study shows the effectiveness of exploiting lower temperatures as a way to enhance federated learning performance.
-However, an important relevant question remains unanswered: “How do we determine the appropriate or optimal temperature for different applications?” With varying dataset and models that different federated learning application use, the ideal temperature value will also differ.
-The scope of this work is on identifying and quantifying the potential benefits of lower temperature-based local training in federated learning and we leave the task of identifying per-purpose optimal temperatures to individual system designers and future research.
-$\bullet$ Dynamic per-sample temperature scaling.
-While our paper focuses on understanding the advantages of maintaining a constant low temperature for the training process, this may not be the ideal approach when targeting maximum performance.
-In reality, varying temperatures can be applied in the training process with respect to their significance Shin et al.
-(2022); Zeng et al.
-(2021).
-That would mean that the training process can exploit lower temperatures for the samples that is “harder” to learn, while higher temperatures are used for those easier to understand.
-Alternatively, a more course-grain approach of dynamically adjusting the temperature on a per-federated learning round-basis can be another meaningful approach to adapt towards changing local data characteristics.
-$\bullet$ Supporting heterogeneous federated learning.
-While many federated learning frameworks assume homogeneous devices, recent work have pointed out that this is not always the case Park and Ko (2023).
-In a federated learning network consisting of devices with heterogeneous computing/network capabilities, the issue of “stragglers”, which are devices that take longer to complete local training, becomes important to address.
-Leveraging insights gleaned from our empirical observations, applying lower temperatures for potential stranglers can be an interesting way of uniforming the local training duration.
-This would require a prior analysis on the conditions expected at each device and a global consensus on the temperature configuration.
-We find this as an interesting topic that requires further investigation.
-6 Conclusion
-This research introduces a novel training approach for federated learning systems, FLex&\&&Chill, aimed at enhancing convergence and improving model performance by exploiting a temperature scaling method termed logit chilling in federated learning local model training operations.
-Specifically, our work revolves around the implications of employing fractional temperature values ($10<T<$ 1) during the model training process.
-Our findings, derived from an extensive set of evaluations using three datasets and three widely applied neural networks for federated learning, demonstrate that exploiting the FLex&\&&Chill approach in the client-side model training process holds the potential to expedite model convergence by inducing more pronounced changes despite a limited number of local samples.
-Consequently, we show that by doing so, we can also achieve an overall higher inference accuracy.
-While the impact of lower temperature vary with respect to the neural network architecture and dataset, our evaluations reveal up to 6.00$\times$ improvement in federated learning model convergence speed and 3.37% improvement in inference accuracy.
-We see this work as a first step into devising a more fine-tuned (and optimal) training mechanism for federated learning systems.
+C.2 FLex&\&&Chill with Advanced FL Schemes
+As mentioned in Section 3.2, FLex&\&&Chill is an orthogonal approach to enhance the federated learning process.
+To demonstrate this, we trained the 2-layer CNN with FedProx Li et al.
+[2020] and SCAFFOLD Karimireddy et al.
+[2020] representing more recently proposed FL schemes compared to FedAvg used in this work.
+Hyperparameters FedProx SCAFFOLD total clients 50 communication rounds 300 optimizer SGD learning rate 0.001 local epoch 5 participants per round 5 batch size 32 $\alpha$ 0.5 loss function Cross Entropy $\mu$ 0.01 N/A
+Table 6: Default hyperparameters used in this section. $\alpha$ denotes the importance of the Dirichlet distribution.
+(a) FedProx
+(b) SCAFFOLD
+Figure 9: Test accuracy and test loss on the CIFAR10 dataset with varying training temperatures with advanced federated learning schemes.
+Overall, models trained with lower temperatures showed superior performance.
+Appendix D Experiment on hyperparameters
+In this section, we explore the effect of other federated learning hyperparameters (i.e., batch size, number of local epochs, number of participants, and degree of data disparity).
+Through this set of experiments, we observed that applying low temperatures during local training improves test accuracy and convergence speed, being independent from hyperparameters.
+We note that the experiments were done by applying the default setting except for the control factors.
+D.1 Batch size
+(a) Batch size 8
+(b) Batch size 16
+(c) Batch size 32
+(d) Batch size 64
+(e) Batch size 128
+Figure 10: Test accuracy and test loss over federated learning round with varying the batch size and the training temperature on the CIFAR10 dataset.
+Relatively small batch size shows better training quality with stability.
+D.2 Number of participants
+(a) Number of participants 5
+(b) Number of participants 10
+(c) Number of participants 15
+Figure 11: Test accuracy and test loss over federated learning round with varying the number of participants and the training temperature on the CIFAR10 dataset.
+When relatively many clients participate in the training, it shows better training quality with stability
+D.3 Number local epochs
+(a) Number of local epoch 1
+(b) Number of local epoch 5
+(c) Number of local epoch 10
+(d) Number of local epoch 15
+Figure 12: Test accuracy and test loss over federated learning round with varying the number of local epochs and the training temperature on the CIFAR10 dataset.
+Relatively many updates in the local training phase show better training quality with stability
+D.4 Degree of data disparity
+(a) $\alpha$=0.1
+(b) $\alpha$=0.5
+(c) $\alpha$=1.0
+Figure 13: Test accuracy and test loss over federated learning round with varying the degree of data disparity and the training temperature on the CIFAR10 dataset.
+Relatively moderate data disparity shows better training quality with stability
+Hyperparameters Appendix D total clients 50 communication rounds 100 optimizer SGD learning rate 0.001 local epoch 5 participants per round 5 batch size 32 $\alpha$ 0.5 loss function Cross Entropy
+Table 7: Default hyperparameters used in this section. $\alpha$ denotes the importance of the Dirichlet distribution.
+Appendix E Details of Centered Kernel Alignment (CKA)
+E.1 Background
+Centered kernel alignment (CKA) is a method to compute the similarity between two different neural networks by measuring the similarity of the network representation for the given input data.
+Let $X$ and $Y$ be the features generated from two different neural networks, we can calculate the CKA by Eq 3.
+As we are assuming a federated learning scenario, the clients share the same neural network architecture.
+Thus, the dimensions of the two features from each corresponding layer are identical, allowing for a direct comparison to be possible.
+We note that in our experiments, we exploited the linear CKA to measure the similarity of the feature spaces between two neural networks.
+$\begin{split}CKA(X,Y)&=\frac{||Y^{T}X||^{2}_{F}}{||X^{T}X||_{F}||Y^{T}Y||_{F}}\\ \end{split}$ (3)
+E.2 Limitation of Centered Kernel Alignment evaluation in federated learning
+Here, we discuss our rationale for measuring CKA between a client model participating in the federated learning round and a reference model trained using i.i.d.
+configurations.
+As mentioned in Section 4.4, the feature space similarity itself only reports the similarity of neural networks and does not assure high quality (i.e., accuracy) of the trained models.
+Figure 14 illustrates the possible cases that can observed in a federated learning scenario.
+We use these four cases to understand how locally trained models converge to the reference model.
+Here, the orange line indicates the update directions of local models from the federated learning training phase (i.e., diverging from the initially shared aggregated model presented as the black dot) and the red dotted line represents the average of these locally updated models that occur at the server.
+The star represents the ideal model in the representation space with high accuracy.
+(a) Case 1
+(b) Case 2
+(c) Case 3
+(d) Case 4
+(e) Case 5
+Figure 14: Illustration of possible cases in a federated learning scenario.
+The solid orange line and dotted red line correspond to the local update direction and aggregated model update direction respectively.
+The black circle and the red star depict the initial point and optimal point respectively.
+$\bullet$ High similarity between local models does not guarantee a better model (Case 1, 2): The first case arises when local models, each trained on their respective local data, produce highly similar updates due to a well-aligned data configuration.
+Such alignment implies that the training data across different clients are relatively similar in distribution and characteristics, allowing the aggregated global model to closely approach the ideal model achievable through centralized training.
+However, in non-i.i.d.
+environments, achieving this scenario is infrequent.
+In non-i.i.d.
+settings, the data distribution varies significantly across different local nodes.
+This variation can lead to local models learning disparate aspects of the data, resulting in updates that are less similar.
+Consequently, the aggregated global model may not perform as well as in an i.i.d.
+scenario, as it has to reconcile these diverse and potentially conflicting updates.
+On the other hand, Case 2 represents a critical scenario in federated learning where there is a high similarity among local model updates, but these updates are aligned in a suboptimal or incorrect direction.
+This situation can be considered one of the worst cases in federated learning.
+In addition, using these two cases, we demonstrate that the high similarity between local models does not ensure a superior update or quality of training.
+$\bullet$ A superior model can be obtained without similar local models (Case 3): Case 3 illustrates a case where a superior model can be achieved even when local models exhibit relatively low similarity.
+In this scenario, each local model learns effective but diverse features from heterogeneous data configurations, and these diverse updates are properly aggregated at the central server.
+Unlike scenarios where local models produce highly similar updates, in this case, the updates from different local models show significant variation.
+This diversity reflects the different aspects of the data each local model has learned.
+This situation arises under conditions of data heterogeneity, where the data distributions across different clients are diverse.
+Each local model captures unique features relevant to its specific data subset.
+Despite the low similarity among local model updates, the central server effectively aggregates these diverse updates.
+The aggregation process leverages the complementary information from each local model, resulting in a robust and well-generalized global model.
+This case supports the hypothesis that a proper global model can be obtained even when local models do not exhibit high similarity and is a globally ideal case for federated learning scenarios.
+$\bullet$ Relationship between model similarity and the magnitude of updates (Case 4, 5): It is important to note that the magnitude of updates can influence the perceived similarity of the network.
+If the update magnitude is small, the similarity of the network may be measured as high, whereas if the update magnitude is large, the similarity may be measured as relatively low.
+This example underscores the need to rethink the conventional emphasis on high similarity as an indicator of performance.
+It suggests that focusing solely on similarity metrics, while providing insights on how local training operations induce model changes, might overlook other critical factors that contribute to the effectiveness of the aggregated model.
+Therefore, a more nuanced approach to evaluating model updates and their impact on the accuracy performance is required, considering both the magnitude of updates and their alignment with the overall learning objectives.
+Based on the aforementioned discussions, we introduce measuring similarity with a reference model as an alternative evaluation method in this work.
+Although we cannot ensure that the reference model trained in an i.i.d.
+data configuration perfectly represents the optimal update objective, we consider this method as an effective and reasonable way to evaluate the federated learning process.
+This approach allows for a more nuanced assessment of the updates by comparing them to a standard benchmark, helping to identify whether the aggregated model is progressing toward a semi-optimal solution despite the inherent data heterogeneity.
+E.3 Experimental results
+Table 8 presents the exact numbers for the accuracy and CKA values discussed in Section 4.4 and Figure 5. Figure 15 illustrates the data distribution applied in the experiments.
+Temperature ($T$) Accuracy (%) Layer 1 CKA Layer 2 CKA Layer 3 CKA Layer 4 CKA 4.00 49.72 0.6539 0.5632 0.7149 0.7661 2.00 53.82 0.6792 0.6872 0.7417 0.7963 1.00 58.19 0.7419 0.6997 0.7644 0.8214 0.50 63.48 0.9322 0.7459 0.8181 0.8559 0.25 66.85 0.9569 0.8161 0.8418 0.8830 0.05 72.76 0.9390 0.9059 0.8898 0.9026 Reference Model ($T$=1.00, i.i.d.) 73.86 1.0000 1.0000 1.0000 1.0000
+Table 8: Accuracy and CKA values of 2 layer CNN measured on the CIFAR10 dataset in Figure 5.
+Figure 15: Visualization of the distribution of training data used in CKA evaluation. Best viewed in color.
+(a) $T$=0.05
+(b) $T$=0.25
+(c) $T$=0.5
+(d) $T$=1.0
+(e) $T$=2.0
+(f) $T$=4.0
+Figure 16: Client-wise feature space similarity matrix measured in the first convolution layers (Layer 1) by using CKA and inference accuracy in different temperatures (0.05, 0.25, 0.5, 1.0, 2.0, 4.0) and federated rounds (1, 5, 50, 100).
+(INIT: Initial dispatched model in the given round, AGG: Aggregated global model, REF: Reference model, ACC: Accuracy)
+(a) $T$=0.05
+(b) $T$=0.25
+(c) $T$=0.5
+(d) $T$=1.0
+(e) $T$=2.0
+(f) $T$=4.0
+Figure 17: Client-wise feature space similarity matrix measured in the second convolution layers (Layer 2) by using CKA and inference accuracy in different temperatures (0.05, 0.25, 0.5, 1.0, 2.0, 4.0) and federated rounds (1, 5, 50, 100).
+(INIT: Initial dispatched model in the given round, AGG: Aggregated global model, REF: Reference model, ACC: Accuracy)
+(a) $T$=0.05
+(b) $T$=0.25
+(c) $T$=0.5
+(d) $T$=1.0
+(e) $T$=2.0
+(f) $T$=4.0
+Figure 18: Client-wise feature space similarity matrix measured in penultimate layers (Layer 3) by using CKA and inference accuracy in different temperatures (0.05, 0.25, 0.5, 1.0, 2.0, 4.0) and federated rounds (1, 5, 50, 100).
+(INIT: Initial dispatched model in the given round, AGG: Aggregated global model, REF: Reference model, ACC: Accuracy)
+(a) $T$=0.05
+(b) $T$=0.25
+(c) $T$=0.5
+(d) $T$=1.0
+(e) $T$=2.0
+(f) $T$=4.0
+Figure 19: Client-wise feature space similarity matrix measured in classification layers (Layer 4) by using CKA and inference accuracy in different temperatures (0.05, 0.25, 0.5, 1.0, 2.0, 4.0) and federated rounds (1, 5, 50, 100).
+(INIT: Initial dispatched model in the given round, AGG: Aggregated global model, REF: Reference model, ACC: Accuracy)

### v3

@@
 Abstract
-Federated learning is inherently hampered by data heterogeneity: non-i.i.d.
-training data over local clients.
-We propose a novel model training approach for federated learning, FLex&\&&Chill, which exploits the Logit Chilling method.
-Through extensive evaluations, we demonstrate that, in the presence of non-i.i.d.
-data characteristics inherent in federated learning systems, this approach can expedite model convergence and improve inference accuracy.
-Quantitatively, from our experiments, we observe up to 6$\times$ improvement in the global federated learning model convergence time, and up to 3.37% improvement in inference accuracy.
+[…]
+keywords:
+Federated Learning , Temperature Scaling
+journal: Advanced Engineering Informatics
+[…]
+[…]
 1 INTRODUCTION
-Federated learning represents a paradigm enabling the training of effective models within distributed environments without explicitly exposing local data McMahan et al.
-(2017).
-This approach provides opportunities for communication and computation efficiency during training by sharing model parameters/gradients learned independently from local learning processes while suppressing raw data transmissions.
-Furthermore, since training operations take place in a decentralized manner, federated learning opens the possibility of parallelizing the model training process via the assistance of local client devices.
-It presents an attractive solution for handling and exploiting the vast amounts of data generated by many mobile and Internet of Things (IoT) device entities consisting of the overall system Park et al.
-(2023a).
-In contrast to conventional centralized training methods, federated learning faces challenges associated with delayed model convergence and limited performance.
-A fundamental challenge of federated learning arises from the disparity of training data used at each client device.
-Federated stochastic optimization encounters difficulties in determining optimal parameters, leading to increased communication overhead between local devices and the central server, as well as the need for additional training on (typically resource-limited) local devices Zhao et al.
-(2018).
-To address these challenges, a number of prior research have focused on optimizing local training operations on federated learning clients or on refining the server-side aggregation process.
-For instance, FedProx Li et al.
-(2020) controls the number of local device iterations, aiming to train models resilient to challenges posed by non-independent and non-i.i.d.
+Federated learning represents a paradigm enabling the training of effective models within distributed environments without explicitly exposing local data [mcmahan2017communication].
+This approach provides opportunities for communication and computation efficiency by sharing model parameters and gradients learned independently from local learning processes while suppressing raw data transmissions and potential data leakage [lee2023exploiting, lee2025mind].
+Furthermore, since training operations take place in a decentralized manner, federated learning opens the possibility of parallelizing the model training process via the assistance of local client resources.
+As a result, federated learning presents an attractive solution for handling and exploiting the vast amounts of data generated by many mobile and Internet of Things (IoT) device entities comprising the overall system [park2023attfl, shin2024effective].
+However, in contrast to conventional centralized training methods, federated learning faces challenges associated with delayed model convergence and limited performance.
+More specifically, a fundamental challenge of federated learning arises from the disparity of training data used at each client device.
+Given this inevitable training data disparity rising from its distributed nature, federated stochastic optimization encounters difficulties in determining optimal parameters, which can lead to increased communication overhead between local devices and the central server, as well as needing additional training on (typically resource-limited) local devices [zhao2018federated].
+To address this challenge, several prior work have proposed schemes to optimize local training operations for federated learning clients or have targeted to refine the server-side model aggregation operations.
+For instance, FedProx [li2020federated] appends a proximal term, aiming to train models resilient to challenges posed by non-independent and non-i.i.d.
 data environments.
-SCAFFOLD Karimireddy et al.
-(2020) achieves expedited convergence and improved model accuracy McMahan et al.
-(2017) by introducing a correction term in the model aggregation phase to balance the influence of each client.
+SCAFFOLD [karimireddy2020scaffold] achieves expedited convergence and improved model accuracy by introducing a correction term in the local model training phase to balance the influence of each client.
 These operations alleviate challenges posed by the non-i.i.d.
-environment, a common dataset configuration for federated learning clients.
-This paper proposes a novel model training approach for federated learning, FLex&\&&Chill (Federated Learning EXploiting logit Chilling), that targets to boost the convergence and model performance via temperature scaling, namely logit chilling, during the local model training operations.
-Note that a typical neural network generates a probability ($p_{i}$) for each class from its neural network logits ($z_{i}$) by applying a softmax operation (Eq.
-1) with the temperature $T$ conventionally set to 1.
-However, it is not uncommon in other application scenarios to exploit even higher temperature values (i.e., $T\in(1,\infty)$), for example, in knowledge distillation Hinton et al.
-(2015); Touvron et al.
-(2021) or neural network calibration Guo et al.
-(2017).
-$\xi(z,i,T)=p_{i}=\frac{exp(z_{i}/T)}{\sum_{j}exp(z_{j}/T)}$ (1)
-When employing a temperature $T~{}>$ 1, the probability distribution across classes undergoes smoothing due to the narrowed interval between logits Guo et al.
-(2017), making such an approach suitable for applications such as knowledge distillation.
-Notably, despite altering the probabilities, the order of labels’ probability remains unaffected with respect to changes in $T$.
-While these works successfully explore the impact of varying temperatures for neural network operations, they mainly focus on applying high temperatures after the model training process, mostly for inference operations.
-Since increasing $T$ reduces the training confidence, such schemes are effective in taming overconfident DNNs.
-On the other hand, in federated learning, where disparate training data is common, increasing model confidence of each individually trained client model (using small $T$) prior to parameter aggregation can be a more effective approach.
-Our investigation involves measuring both the gradient flow and data position shift within the representation space to provide guidelines in this unexplored territory.
-Specifically, this work targets to show that utilizing fractional temperature values (i.e., $T\in(0,1)$) during the training process enables effective gradient propagation towards the input layer when applied to federated learning data distribution scenarios.
-Consequently, we observed that lower temperatures facilitate more efficient data updates within the representation space when dealing with non-i.i.d.
-datasets, a common characteristic of federated learning systems.
-From our evaluations with three datasets and three baseline federated learning models, we show that these effects accelerate the federated learning model convergence time by up to $\mathbf{6.00\times}$ and improve the inference accuracy by up to $\mathbf{3.37\%}$, despite potential concerns on model instability when training with with low temperatures.
-We summarize the contributions of this work in three-fold as the following:
-  • To the best of our knowledge, this is the first work that explores the impact of integrating temperature scaling during the training process in a federated learning scenario.
-  • We offer empirical evidence that demonstrates the performance enhancements of exploiting logit chilling in federated learning scenarios, by examining the gradient flow and data position shifts within the representation space.
-  • We introduce a novel federated learning model training approach, FLex&\&&Chill, showcasing the application of low temperature usage during the training process.
-2 Background and Related Work
+data environment, a common data distribution for federated learning clients.
+We note that a potentially orthogonal approach to these prior efforts is to exploit the fact that a typical neural network generates a probability ($p_{i}$) for each target class from its neural network logits ($z_{i}$) by applying a softmax operation (Eq.
+1) with a temperature $T$ value conventionally set to 1.
+However, it is not uncommon to exploit even higher temperature values (i.e., $T\in(1,\infty)$) in use cases such as knowledge distillation [hinton2015distilling, touvron2021training] or neural network calibration [guo2017calibration].
+$p_{i}=\frac{exp(z_{i}/T)}{\sum_{j}exp(z_{j}/T)}$ (1)
+When employing a temperature $T>1$ for scaling the logits, the probability distribution across classes undergoes smoothing due to the narrowed inter-logit intervals [guo2017calibration], making higher temperature adjustments suitable for applications such as knowledge distillation.
+This is possible since, despite altering the model’s internal probabilities, the order of labels’ probabilities remains unaffected with respect to changes in $T$.
+While the use of diverse temperatures for neural network operations has been shown to be effective, works presented until now mainly focus on applying high temperatures after the model is trained, making it mostly applicable for the inference operations only.
+For instance, since increasing $T$ reduces the model confidence, such schemes are effective in curbing the overconfidence of DNNs [guo2017calibration].
+On the other hand, the effect of exploiting temperature scaling during the training process has gathered relatively limited attention in the current literature [agarwala2020temperature].
+Despite the distinctive characteristics of federated learning scenarios (e.g., data disparity and training parallelism), the benefits of employing temperature scaling during the federated training operations are left understudied.
+This paper proposes a novel model training approach for federated learning, FLex&\&Chill (Federated Learning EXploiting logit Chilling), that targets to boost the model convergence and performance via temperature scaling, namely logit chilling, during the local model training operations.
+This paper analyzes the efficacy and effectiveness of employing temperature scaling during the local training process in a federated learning scenario via theoretical convergence analysis and empirical evidence to provide guidelines in this unexplored territory.
+Specifically, we target to show that utilizing fractional temperature values (i.e., $T\in(0,1)$) during model training enables effective gradient propagation towards the input layer when applied to federated learning with non-i.i.d.
+datasets.
+From our evaluations with three datasets and three baseline federated learning models, we show that these effects accelerate the federated learning model convergence time by up to $\mathbf{6.00\times}$ and improve the inference accuracy by up to $\mathbf{3.37\%}$, despite potential concerns on model instability when training with low temperatures.
+Specifically, we summarize the contributions of this work in three folds as follows:
+  • To the best of our knowledge, this is the first work that assesses the impact of integrating temperature scaling during the training process in a federated learning scenario.
+  • We offer theoretical explanations and empirical evidence that demonstrate the performance enhancements of exploiting logit chilling in federated learning scenarios.
+  • We introduce a novel federated learning model training approach, FLex&\&Chill, showcasing the application of low-temperature usage during the training process.
+The remainder of this work is structured as follows.
+The following section will present related work in this domain to position our work among previous work.
+Section 3 presents the design of FLex&\&Chill along with deep analysis of its operations.
+We will present experimental results using three datasets in Section 4 and suggest interesting discussion points in Section 5.
+Finally, we conclude the paper in Section 6.
+We note that an open source implementation of the proposed scheme can be accessed at https://github.com/eis-lab/temperature-scaling.
+2 RELATED WORK
 2.1 Temperature Scaling
 In machine learning, a model’s output probability distribution serves diverse purposes.
-Primarily, classification models offer probabilities for each class, showcasing the model’s confidence in its predictions.
-This certainty measure allows consideration of the likelihood of each class prior to making a final classification decision.
-Moreover, this distribution offered by the model helps in understanding the uncertainty of a model on its decisions Park et al.
-(2023b); Pearce et al.
-(2021).
-For neural networks, computing probabilities from logits often involves employing softmax operations, as in Eq.
+Primarily, classification models offer a probability for each class, showcasing the model’s confidence in its predictions.
+This model-generated certainty measure allows consideration of the likelihood of each class prior to making a final classification decision.
+Moreover, this distribution offered by the model helps in understanding the uncertainty of a model on its decisions [park2023self, pearce2021understanding].
+Figure 1: Effect of varying temperature $T$ on the output distribution of the softmax function, illustrating how lower $T$ sharpens class probabilities and higher $T$ produces smoother, more uniform distributions.
+Best viewed in color.
+For neural networks, computing label probabilities from logits often involves softmax operations, which is a special case ($T=1$) of a Boltzmann distribution as in Eq.
 1.
-Here, the temperature parameter ($T>0$) controls the probability distribution’s concentration towards one class, and manages how the categorical probability is dispersed over possible classes Jang et al.
-(2016).
-Notably, this categorical probability aligns with the concept of entropy in information theory, offering a perspective where $T$ governs the level of disorder or uncertainty within the distribution.
-Specifically, with $T\in(1,\infty)$, the probability distribution tends to be smoother as $T$ increases.
-This arises from the diminished differences between logit values, enhancing the distribution’s entropy.
-In contrast, $T\in(0,1)$ sharpens the probability distribution, since a lower $T$ amplifies inter-logit differences, effectively polarizing the distribution and concentrating probabilities closer to the maximum logit value.
-Consequently, model predictions are made more confidently, by allocating higher probabilities to fewer classes, thereby reducing uncertainty in the distribution.
+Here, the temperature parameter ($T>0$) controls the concentration of probability towards the class with the largest logit and manages how the categorical probability is dispersed over possible classes [jang2016categorical].
+Figure 1 shows how the per-class probability distribution changes for varying temperature (i.e.
+$T$) values.
+Specifically, with $T\in(1,\infty)$, the probability distribution tends to be smoother as $T$ increases (green line in Fig.
+1).
+This arises from the diminished ratio between exponentials of equally scaled values.
+In contrast, $T\in(0,1)$ shifts the probability distribution into a point-mass, as a lower $T$ amplifies the ratio of the exponential of scaled logits, effectively polarizing the distribution and concentrating probabilities to the maximum logit value (blue line in Fig.
+1).
+Consequently, model predictions are made more confidently by allocating higher probabilities to fewer classes.
 Therefore, temperature scaling operates as a means to adjust the probability distribution within a neural network-based system.
-Previous work in Knowledge Distillation Hinton et al.
-(2015); Touvron et al.
-(2021) exploit the temperature value by applying a high temperature $T\in(1,\infty)$ to achieve a smoothed probability distribution.
-This manipulation aims to magnify small values within the distribution, often representing hidden information encapsulated within the model.
-Additionally, in generative models Shih et al.
-(2023); Wang et al.
-(2020), temperature scaling regulates the randomness inherent in model outputs.
-This technique allows for the control of variability levels in generated samples, enabling adjustments to generated output diversity.
-Temperature scaling is also used in model calibration Guo et al.
-(2017).
-Applying temperature adjustments to the model’s output probabilities, typically through a scaling factor $T$, aids in aligning these probabilities more accurately with the true likelihoods or uncertainties present in the data.
-Likewise, previous work has focused on applying high-temperature values for various purposes, mostly for the purpose of conservatively better understanding the full dataset without bias.
+Temperature scaling was applied in previous works for knowledge distillation [hinton2015distilling, touvron2021training], generative models [shih2023long, wang2020contextual], gradient analysis [lee2024detrigger] and model calibration [guo2017calibration].
+Applying temperature adjustments to the model’s output probabilities, typically through a scaling factor $T$, aids in aligning the probabilities more accurately with the true likelihoods or uncertainties present in the data.
+However, most prior works have focused primarily on applying temperature scaling as a post-training technique for probability calibration, leaving its effects during training largely unexplored.
+While pioneering research [agarwala2020temperature] investigated temperature scaling during training and showed its positive impact in model generalization, their analysis was limited to a centralized learning scenario.
+Furthermore, works such as FedRS propose changing local objective functions to apply temperature scaling in federated model training operations [li2021fedrs].
+However, the work is rather heuristic and does not present a thorough theoretical explanation of why altering the softmax function can improve convergence and mitigate data heterogeneity.
+In contrast, our work explicitly highlights (i) the effect of temperature scaling applied during the training phase rather than post-training, and (ii) its impact specifically in federated learning scenarios characterized by heterogeneous, non-i.i.d.
+data distributions.
 2.2 Federated Learning
-In recent years, federated learning has emerged as a promising paradigm in machine learning, particularly in scenarios where data privacy, security, and the decentralization of model training are core concerns.
+Recently, federated learning has emerged as a promising paradigm in machine learning, particularly in scenarios where data privacy, security, and the decentralization of model training are core concerns.
 Unlike traditional centralized training approaches, federated learning enables model training across distributed devices or servers while keeping the data localized; thus, addressing privacy challenges associated with centralized data aggregation.
 This decentralized learning paradigm involves iterative model training where local updates occur on individual devices, and only aggregated model updates are shared with the central server.
-Despite its benefits, federated learning poses several challenges.
+Specifically, federated learning aims to solve the following optimization problem:
+$\min_{w}F(w)=\sum_{k=1}^{K}p_{k}F_{k}(w),$
+where $K$ denotes the total number of clients participating in training, $w$ represents the global model parameters to be optimized, $p_{k}\geq 0$ is the weight corresponding to the $k$-th client (typically proportional to the size of the local dataset), and $F_{k}(w)$ denotes the local objective function defined as follows:
+$F_{k}(w)=\frac{1}{|D_{k}|}\sum_{(x,y)\in D_{k}}\ell(w;x,y),$
+where $D_{k}$ is the local dataset on client $k$, and $\ell(w;x,y)$ represents the loss function evaluating the prediction of the model $w$ for data sample $(x,y)$.
+The key challenge of federated learning lies in optimizing the global objective $F(w)$ under the constraint that the clients’ data remains distributed and private, each client trains locally without directly sharing its raw data.
+Despite its benefits and active research, federated learning poses several challenges.
 A body of work has tackled the issue of communication overhead stemming from the exchange of frequent model parameter updates.
-A number of previously proposed works have proposed schemes to effectively abstract the model parameters in various forms to relieve the communication and networking overhead Park et al.
-(2023a).
-Furthermore, federated learning systems are prone to facing issues arising from data heterogeneity caused by the diversity or differences in the characteristics, distributions, or formats of data across federated learning clients.
-Previous works have attempted to address this challenge by addressing relevant issues such as variations in feature/input variations Yu et al.
-(2021) and spatio-temporal variations Liu et al.
-(2020).
+A number of previously proposed works have proposed schemes to effectively abstract the model parameters in various forms to relieve the communication and networking overhead [park2023attfl].
+Furthermore, federated learning systems are prone to facing issues arising from data heterogeneity caused by the diversity or differences in the characteristics, distributions, or formats of data spread across distributed federated learning clients.
+Previous works have attempted to address this challenge by addressing relevant issues such as variations in feature/input variations [yu2021fed2] and spatio-temporal variations [liu2020client].
 3 DESIGN OF FLEX&CHILL
-We initiate our discussions by exploring the impact of logit scaling, also known as temperature scaling, providing empirical evidence on the rationale behind our proposed system.
-Furthermore, in this section, we introduce our system, referred to as FLex&\&&Chill, a novel local training approach designed for efficiently training a federated learning framework.
-3.1 Impact of Training with Lower Temperatures
-As aforementioned, conventionally, temperature scaling has primarily been focused on calibrating the probability distribution of a model’s output after the model training phase, with a limited exploration into its application during the training process itself.
-In this section, we present empirical results on three initial experiments specifically crafted to showcase the implications of employing temperature scaling throughout the training phase operations in different dimensions.
-Through these investigations, our aim is to establish a foundation for comprehending the potential implications and advantages associated with the utilization of temperature scaling techniques in neural network training.
-We conduct our experiments with the CIFAR10 dataset Krizhevsky et al.
-(2009) and use a 2-layered CNN architecture McMahan et al.
-(2017) as the baseline neural network architecture.
-For experimental details, see APPENDIX B
-3.1.1 Gradient flow patterns
-Figure 1: Distribution of gradient norm at input layer for correctly/incorrectly inferred samples with varying training temperatures.
+We begin discussing our framework design by examining the effect of temperature scaling (also referred to as logit scaling), providing both theoretical and empirical justification for our approach.
+On the theoretical side, we analyze how temperature scaling influences training dynamics in federated learning, with a particular focus on gradient amplification and its implications for convergence.
+Empirically, we investigate how applying temperature scaling alters the data distribution and decision boundaries within the representation space.
+Building on these insights, we introduce FLex&\&Chill, a novel local training strategy tailored to enhance the efficiency and performance of federated learning systems.
+3.1 Theoretical Analysis
+As aforementioned, conventionally, temperature scaling has primarily been focused on calibrating the probability distribution of a model’s output after the model training phase, with limited exploration into its application during the training process itself.
+In this section, we provide a theoretical analysis of the effect of temperature scaling during training with respect to the convergence in a federated learning scenario.
+Based on this investigation, we aim to establish a foundation for comprehending the potential implications and advantages associated with the utilization of temperature scaling techniques in neural network training.
+$\frac{\partial\mathcal{L}}{\partial{z_{i}}}=\frac{1}{T}(p_{i}-y_{i})=\frac{1}{T}\left(\frac{e^{z_{i}/T}}{\sum_{j}e^{z_{j}/T}}-y_{i}\right)$ (2)
+Eq.
+2 presents the gradient of cross-entropy for the softmax activation ($\mathcal{L}$) with varying temperature $T$ and hard target, one-hot encoded labels.
+For brevity, we present the final derivation result and present detailed proof in the Appendix.
+Since the gradient is multiplied by $\frac{1}{T}>1$, for low temperatures $T<1$, we can expect a boost in the magnitude of the gradient.
+Furthermore, from the perspective of the loss term $(p_{i}-y_{i})$, when the model yields an incorrect decision, the norm of the loss term significantly increases as $p_{i}\xrightarrow{}0$.
+Building upon the theoretical analysis on the convergence of federated learning by Li et al.
+[Li2020On], we can derive the inequality Eq.
+3, where $F(w^{t})$ is the loss of the neural network with the weight ($w^{t}$) at round $t$, Lipschitz constant $L$, the convexity constant $\mu$, and the upper bound of variance of the stochastic gradient at local devices $\sigma^{2}$.
+The inequality suggests that the difference between the optimal weight $w^{*}$ and the actual weight ($w_{t}$) can be minimized by iteratively updating the weight.
+$\begin{split}&\mathbb{E}[F(w^{(t+1)})-F(w^{*})]\leq\\ &(1-\frac{\eta\mu}{T})\mathbb{E}[F(w^{(t)})-F(w^{*})]+\frac{L\eta^{2}\sigma^{2}}{2T^{2}}.\end{split}$ (3)
+Here, we can identify the role of softmax temperature ($T$) regarding the convergence of federated settings.
+In the inequality, the temperature $T$ controls the speed of convergence.
+For instance, a high temperature ($T>1$) slows the convergence as it reduces the term $\frac{\eta\mu}{T}$.
+Furthermore, $T$ also affects the stability of convergence as it controls the variance term ($\frac{L\eta^{2}\sigma^{2}}{2T^{2}}$).
+$T>1$ also reduces the amplitude of the variance term, which consequently stabilizes the convergence (Proof in Appendix).
+Based on these findings, we can consider the temperature as a key hyperparameter in federated learning.
+We note that the above discussion can be extended to more complex adaptive optimization methods.
+For brevity, these details are provided in the Appendix.
+Besides the mathematical analysis, to provide an intuitive rationale for applying temperature scaling during local training, the low temperatures effectively penalize incorrect predictions strongly while rewarding correct ones, enabling aggressive learning focused on challenging samples.
+This intuition aligns with the experimental results presented in the following section.
+However, it is important to note that excessively aggressive updates may cause instability.
+3.2 Temperature v.s. Learning Rate
+Since $\eta$ and $T$ have a reciprocal relationship in Eq.
+3, we can question the relationship between the learning rate ($\eta$) and the temperature ($T$).
+Specifically, Will decreasing or increasing the temperature be equivalent to increasing or decreasing the learning rate?
+We emphasize that while temperature and learning rate exhibit similar effects on convergence, they are not identical.
+$w_{t+1}$ $=w_{t}-\eta\nabla f(w_{t})$ (4)
+$=w_{t}-\frac{\eta}{T}\left(\frac{e^{z_{i}/T}}{\sum_{j}e^{z_{j}/T}}-y_{i}\right)$ (5)
+Eq.
+4 presents the stochastic update of a neural network parameter $w_{t}$ at step $t$ with learning rate $\eta$.
+In Eq.
+5, the update rule can be written by introducing the softmax temperature parameter $T$.
+Dividing the learning rate $\eta$ by the temperature $T$ suggests that increasing $T$ would have a similar effect to reducing $\eta$, and vice versa.
+$T$ has, however, an additional effect in calculating the Boltzmann distribution inside the parentheses.
+$\eta$ only controls the magnitude of the parameter update, affecting the overall convergence dynamics.
+On the other hand, $T$ has two-fold effects: changing convergence dynamics and modulating the output distribution’s sensitivity.
+$T$ impacts the relative importance of individual gradients and their direction rather than only uniformly scaling the update size.
+Therefore, although $\eta$ and $T$ appear to play interchangeable roles for some cases, the underlying mechanisms through which they influence the training process are fundamentally different.
+3.3 Empirical Analysis
+In this section, we present empirical results via three initial experiments specifically crafted to support the analytical results and showcase the implications of employing temperature scaling throughout the training phase operations in different dimensions.
+Through these investigations, we aim to establish a foundation for comprehending the potential implications and advantages associated with the utilization of temperature scaling techniques in neural network training.
+We conduct our experiments with the CIFAR10 dataset [krizhevsky2009learning] and use a 2-layered CNN architecture [mcmahan2017communication] as the baseline neural network architecture.
+Figure 2: Distribution of gradient norm at input layer for correctly (top) / incorrectly (bottom) inferred samples with varying training temperatures.
+Input gradient norm.
+Our first experiment focuses on assessing the norm of computed gradients within the input layer for various temperature values ($T$), and we present a sample distribution plot of the norm of input layer gradients for 50K training samples in Figure 2.
+Note that here we separate the plots for the correctly inferred samples (in the forward pass) and the incorrect ones to observe their differences.
+The propagation of gradients to the input layer serves as an indicative metric for the model’s learning efficacy.
+Successful gradient propagation to the input layer implies that the model effectively utilizes and comprehends information from diverse features within the input data during the training phase [evci2022gradient].
+As shown in Figure 2, the mode of distribution shifts to larger values with decreasing temperatures for both correct and incorrect predictions.
+As mentioned, the reason behind this mode shift stems from the gradient function characteristics of softmax.
+As shown in Eq.
+2, since the gradient is multiplied by $\frac{1}{T}>1$, for low temperatures $T<1$, we can expect a gradient boost.
+Furthermore, from the perspective of the loss term $(p_{i}-y_{i})$, when the model yields an incorrect decision, the loss term significantly increases as $p_{i}\xrightarrow{}0$.
+The double-folded effect accounts for a substantial mode shift with thick and long tails in the gradient distribution for lower temperatures.
 Figure 3: Distributions depicting differences between distances to the decision boundary before and after model updates for varying training temperatures.
 Notice that lower temperatures show a noticeable shift in estimations’ positions on the representation space, suggesting their aggressiveness in modifying the model even with a small number of training samples.
-Our first experiment focuses on assessing the norm of computed gradients within the input layer for various temperature values ($T$), and we present a sample distribution plot of how the gradients flow for 50K training samples in Figure 2.
-Note that here we separate the plots for the correctly inferred samples (in the forward pass) and the incorrect ones to observe their differences.
-The propagation of gradients to the input layer serves as an indicative metric for the model’s learning efficacy.
-Successful gradient propagation to the input layer implies that the model effectively utilizes and comprehends information from diverse features within the input data during the training phase Evci et al.
-(2022).
-As shown in Figure 2, the mode of distribution shifts to larger values with decreasing temperatures for both correct and incorrect predictions.
-One reason behind this mode shift stems from the gradient function characteristics of softmax.
-Eq.
-2 presents the gradient of cross-entropy for the softmax activation ($\mathcal{L}$) with varying temperature $T$ and hard target, one-hot encoded labels.
-For brevity, we present the final derivation result and leave details in the Appendix (see Appendix A).
-Since the gradient is multiplied by $\frac{1}{T}>1$, for low temperatures $T<1$, we can expect a gradient boost.
-$\frac{\partial\mathcal{L}}{\partial{z_{i}}}=\frac{1}{T}(p_{i}-y_{i})=\frac{1}{T}(\frac{e^{z_{i}/T}}{\sum_{j}e^{z_{j}/T}}-y_{i})$ (2)
-Furthermore, from the perspective of the loss term $(p_{i}-y_{i})$, when the model yields an incorrect decision, the loss term significantly increases as $p_{i}\xrightarrow{}0$.
-The double-folded effect accounts for a substantial mode shift with thick and long tails in the gradient distribution for lower temperatures.
-3.1.2 Data position shift in the representation space
+Data position shift in the representation space.
 To better understand the impact of temperature scaling on the training process, we conducted a study into the shift of a data’s position in the representation space (i.e., the distance to the decision boundary).
-Drawing inspiration from the concept of ‘travel to decision boundary’ as proposed by Kim et al.
-Kim et al.
-(2022), we quantified the distance of a data point’s inference result to the decision boundary by determining the magnitude of epsilon ($\epsilon$) value multiplied by the adversarial perturbation Goodfellow et al.
-(2014) needed to alter the model’s decision.
+Drawing inspiration from the concept of travel to decision boundary as proposed by Kim et al.
+[kim2022curved], we quantified the distance of a data point’s inference result to the decision boundary by determining the magnitude of epsilon ($\epsilon$) value multiplied by the adversarial perturbation [goodfellow2014explaining] needed to alter the model’s decision.
 For this experiment, we configure an initial neural network and perform inference for 1,000 samples to identify the results’ positions on the representation space.
@@
 Figure 3 plots our results in the form of distributions depicting differences between distances to the decision boundary before and after model updates.
-When higher temperatures ($T={2,4}$) are applied, we can notice that the changes in inference results’ positions are not significant (i.e., plots skewed towards 0).
+When higher temperatures (T=2,4T={2,4}) are applied, we can notice that the changes in inference results’ positions are not significant (i.e., plots skewed towards 0).
 Whereas, with a small $T$, we can notice more pronounced shifts in data positions within the representation space.
 This suggests that using a small $T$ at the softmax operations can actively steer the model towards (potentially) substantial optimization to better align with the training data.
-3.1.3 Visualization in the representation space
+Visualization in the representation space
+Figure 4: Example of data points in the 2D representation space with their respective classification boundaries for different federated learning clients with varying training temperatures.
+Best viewed in color.
 To examine the effect of applying a low temperature during the model training phase, we conducted a simple example that classifies 2-dimensional data via logistic regression.
-We split the data into three different subsets and assign them as three devices with non-i.i.d data.
-The Softmax layer was applied at different temperatures ($T$=1.0, 0.5) during the local training phase and average model weights for aggregation after 10 local updates.
+We split the data into three different subsets and assigned them as three devices with non-i.i.d.
+data.
+The Softmax layer was applied at different temperatures (TT=1.0, 0.5) during the local training phase and average model weights for aggregation after 10 local updates.
 Figure 4 illustrates the classification boundaries and samples distributed in the representation space for the three clients and the aggregated model.
 The results suggest that applying $T$ = 0.5 effectively reflects the local dataset characteristics compared to $T$ = 1.0.
-Specifically, as an example, for clients 0, 1 and the aggregate model, using $T$ = 1.0 fails to make a clear distinction between the red and light green samples; whereas, when applying $T$ = 0.5, this distinction becomes pronounced.
-Figure 3: Example of data points in the 2D representation space with their respective classification boundaries for different federated learning clients with varying training temperatures.
-3.2 FLex&\&&Chill and Logit Chilling
-Based on such empirical observations, we hypothesize that training neural networks with lower temperatures can benefit the model convergence and performance for systems exploiting non-i.i.d.
-data, which is common for federated learning scenarios.
-In FLex&\&&Chill we term this as logit chilling, as we ‘chill down’ the temperature at the softmax operations, thereby altering the logit computation derived from the neural network during the training process.
-The focus of a FLex&\&&Chill-based federated learning system is to exploit logit chilling to train a client-side model with low-temperature values.
-As we observed through the studies above and as our evaluations will show, this has an effect of effectively exploiting crucial information from the training samples to actively alter the model towards the new knowledge.
-The lower temperatures used by Logit Chilling directly influence the local training process of each federated learning client.
-Notably, the seamless integration of Logit Chilling and FLex&\&&Chill with existing federated learning frameworks is a distinct advantage, allowing for a straightforward implementation within already deployed frameworks.
-Furthermore, we note that FLex&\&&Chill can be used orthogonality with other efforts that accelerate federated learning training operations such as FedProx and SCAFFOLD.
-For a detailed explanation and experimental result about the orthogonality of FLex&\&&Chill, please refer to Appendix  C.
-4 Experiment
-We evaluate the performance of FLex&\&&Chill using three datasets: (i) FEMNIST from the LEAF database Caldas et al.
-(2018a), including the Extended MNIST handwriting data Cohen et al.
-(2017) , (ii) the CIFAR10 dataset, and (iii) the CIFAR100 dataset Krizhevsky et al.
-(2009).
-As metrics, we use the average accuracy and training loss observed at the clients, selected to observe both the performance gain and the effectiveness of FLex&\&&Chill in its training operations.
+Specifically, as an example, for clients 0, 1, and the aggregate model, using $T$ = 1.0 fails to make a clear distinction between the red and light green samples; whereas, when applying $T$ = 0.5, this distinction becomes pronounced.
+3.4 FLex&\&Chill and Logit Chilling
+Based on the theoretical analysis and our empirical observations, we find that training neural networks with lower temperatures directly amplifies the gradient magnitude and accelerates convergence in the presence of non-i.i.d.
+data, which is a typical regime for federated learning.
+In particular, our analysis shows that the softmax cross-entropy gradient scales with the factor $1/T$, and the resulting convergence bound includes the term $\eta\mu/T$, indicating faster convergence when $T<1$.
+Complementing this, our empirical studies in the representation space (e.g., Figures 3 and 4) show that lower temperatures move samples farther from the decision boundary, especially for the misclassified points, leading to more aggressive and influential local updates at each client.
+We formulate these insights as an explicit optimization strategy, termed FLex&\&Chill.
+Specifically, FLex&\&Chill exploits logit chilling by “chilling down” the temperature at the softmax layer during local training on each client, thereby sharpening the output distribution and amplifying gradients throughout the network.
+The focus of a FLex&\&Chill-based federated learning system is to leverage these low-temperature updates so that each client can more strongly encode new knowledge into its local model, and the subsequent aggregation of these models at the server accelerates the global convergence toward a high-quality solution.
+As our evaluations in the following sections demonstrate that FLex&\&Chill consistently achieves faster convergence and improved accuracy compared to conventional training with $T=1$ under various non-i.i.d.
+federated learning scenarios.
+We formulate the operations of FLex&\&Chill using Algorithm 1.
+Here, typical federated learning operations take place at both the server and clients except for the fact that the server determines a target $T$ at each round and local model training operations are performed with $T$.
+Since FLex&\&Chill only manipulates the local training process which inherently provides orthogonality to our proposed method, it is extremely easy to integrate with existing federated learning frameworks (e.g., FedProx [li2020federated], SCAFFOLD [karimireddy2020scaffold], FedBN [li2021fedbn]).
+Algorithm 1 FLex&\&Chill
+Data ($D_{1},D_{2},...,D_{N}$) where $D_{i}$ is the user $i$’s local data.
+Model ($w^{1}_{t},w^{2}_{t},...,w^{N}_{t}$) where $w_{t}^{i}$ is the user $i$’s model in $t$-th round.
+Parameter: Temperature $T$
+  Server executes:
+  Initialize $w_{0}$
+  for each round t=1,2,,Rt=1,2,...,R do
+   $S_{t}\xleftarrow{}$ (random set of $m$ clients)
+   for $k\in S_{t}$ do
+    $w_{t}^{k}\xleftarrow{}$ ClientUpdate($w_{t},T$)
+   end for
+   $w_{t+1}\xleftarrow{}$ Aggregation($S_{t}$)
+  end for
+  ClientUpdate($w_{t}$, $T$):
+  $w_{t}^{k}\xleftarrow{}w_{t}$
+  for data, label $\in D_{k}$ do
+   $\mathbf{z}=w_{t}^{k}(data)/T$
+   $l(\mathbf{z})=loss(\mathbf{z};label)$
+   $w_{t}^{k}\xleftarrow{}LocalUpdate(w_{t}^{k},l(\mathbf{z}))$
+  end for
+  Return $w_{t}^{k}$ to server
+The low temperatures used by Logit Chilling directly influence the local training process of each federated learning client.
+Notably, the seamless integration of Logit Chilling and FLex&\&Chill with existing federated learning frameworks is a distinct advantage, allowing for a straightforward implementation within already deployed frameworks.
+Furthermore, as mentioned, we note that FLex&\&Chill can be used orthogonally with other efforts that accelerate federated learning training operations, such as FedProx, SCAFFOLD, and FedBN.
+We provide an open-sourced implementation of FLex&\&Chill at https://github.com/eis-lab/temperature-scaling.
+4 EVALUATION
+We evaluate the performance of FLex&\&Chill using three datasets: (i) FEMNIST from the LEAF database [caldas2018leaf], including the Extended MNIST handwriting data [cohen2017emnist], (ii) the CIFAR10 dataset, and (iii) the CIFAR100 dataset [krizhevsky2009learning].
+As metrics, we use the average accuracy and training loss observed at the clients, selected to observe both the performance gain and the effectiveness of FLex&\&Chill in its training operations.
 Furthermore, we also present the number of federated learning rounds needed to achieve a pre-defined target accuracy.
-(a) FEMNIST-DNN
-(b) CIFAR10-CNN
-(c) CIFAR100-ResNet18
-Figure 4: Average federated learning client-side model accuracy for different temperatures used in training.
-Overall, lower temperatures show higher average model accuracy for federated learning systems.
 4.1 Experiment Setup
-Dataset. We detail the datasets used in our work as follows.
-  • FEMNIST dataset: The FEMNIST dataset comprises of 10 handwritten digits (0-9) and 52 characters (26 lowercase and 26 uppercase) images contributed by 712 users, totaling 157,132 samples.
+Dataset.
+We detail the datasets used in our work as follows.
+(i) The FEMNIST dataset comprises of 10 handwritten digits (0-9) and 52 characters (26 lowercase and 26 uppercase) images contributed by 712 users, totaling 157,132 samples.
 We selected a subset of 36 users, each with diverse amounts of local data to account for the non-i.i.d.
 environment.
-We followed the implementations provided by the original work Caldas et al.
-(2018b).
-  • CIFAR10 and CIFAR100 dataset: We use baseline CIFAR10 and CIFAR100 datasets Krizhevsky et al. (2009) and manually partition data over 50 federated learning users, with each holding 1,000 samples.
+We followed the implementations provided by the original work [caldas2018leaf].
+(ii) We use baseline CIFAR10 and CIFAR100 datasets [krizhevsky2009learning] and manually partition data over 50 federated learning users, with each holding 1,000 samples.
+(a) FEMNIST
+(b) CIFAR10
+(c) CIFAR100
+Figure 5: Visualization of the distribution of training data used in FEMNIST, CIFAR10-CNN and CIFAR100-ResNet experiments. Best viewed in color.
+(a) Severe Heterogeneity ($\alpha$=0.1)
+(b) Moderate Heterogeneity ($\alpha$=0.5)
+(c) Weak Heterogeneity ($\alpha$=1.0)
+Figure 6: Visualization of training data distribution partitioned by the Dirichlet-based method. Best viewed in color.
+Figure 5 visualizes the data distribution used for our evaluations (c.f., Sec 4).
+In this section, we employ two complementary dataset‐splitting strategies.
+First, in Sections 4.24.3, and 4.4, we use pathological (shard-based) partitioning, a common approach for simulating real-world data heterogeneity.
+Here, each class’s data is divided into fixed-size shards, and each client is assigned a subset of these shards without overlap.
+By varying both the number of shards per client and the shard size, we induce controlled quantity skew (different total data amounts) and label skew (uneven class distributions) across clients.
+In image-based tasks such as CIFAR10 and CIFAR100, this shard-based assignment also implicitly introduces feature-level skew, since different clients see different subsets of the underlying intra-class variations (e.g., orientation, background, and lighting), leading to mixed heterogeneity in both label composition and input characteristics.
+Second, for all other experiments in Section 4, we adopt a Dirichlet-based dataset partitioning.
+Figure 6 shows the data distribution sampled via Dirichlet-based partitioning over the clients.
+Concretely, let $K$ be the number of classes and $C$ the number of clients.
+We first draw, for each client $i$, a class‐proportion vector
+$\mathbf{p}^{(i)}\sim\mathrm{Dirichlet}(\alpha,\dots,\alpha)\,$,
+where $\alpha>0$ is the concentration parameter. Then, for each class $k$, we allocate to client $i$ exactly
+$n_{k}^{(i)}=\bigl\lfloor\,|\mathcal{D}_{k}|\;p_{k}^{(i)}\bigr\rfloor$
+samples drawn without overlap from the global pool $\mathcal{D}_{k}$.
+In this way, each client ends up with a total of $\sum_{k}n_{k}^{(i)}$ samples whose class composition follows $\mathbf{p}^{(i)}$.
+By tuning $\alpha$, we control the degree of heterogeneity: (i) A small $\alpha$ (e.g.
+$\alpha=0.1$) yields highly skewed, non‐i.i.d.
+splits (clients see only a few classes); (ii) A moderate $\alpha$ (e.g.
+$\alpha=0.5$) induces moderate class imbalance; (iii) A large $\alpha$ (e.g.
+$\alpha\geq 1$) approaches the i.i.d.
+regime.
+This Dirichlet‐based method, therefore, provides a continuous spectrum of label and quantity skew, allowing us to systematically evaluate our algorithms under varying non‐i.i.d.
+conditions.
+Similar to the shard-based setting, when applied to image datasets, Dirichlet partitioning also induces mixed skew with clients differing not only in class proportions but also in the subset of visual factors they observe; resulting in both label and feature heterogeneity.
+We note that these two widely adopted non-i.i.d.
+benchmarks (shard-based and Dirichlet-based partitioning) already create mixed heterogeneity patterns that combine label skew, quantity skew, and, in the case of image datasets, feature-level skew arising from intra-class variability.
 Model.
@@
 We term this combination as “FEMNIST-DNN” in the rest of our evaluations.
-In the case of the CIFAR10 and CIFAR100 datasets, experiments were conducted using a two-layer CNN model McMahan et al.
-(2017) and a ResNet18 He et al.
-(2016), respectively.
-These two cases are denoted as “CIFAR10-CNN” and “CIFAR100-ResNet18.”
+In the case of the CIFAR10 and CIFAR100 datasets, experiments were conducted using a two-layer CNN model [mcmahan2017communication] and a ResNet18 [he2016deep], respectively.
+These two cases are denoted as “CIFAR10-CNN” and “CIFAR100-ResNet18.” We note that the impact of temperature scaling during the local training process is a model/dataset agnostic phenomenon.
+Layer Details Repetition FC 1 Linear(784, 512) ReLU() $\times 1$ FC 2 Linear(512, 256) ReLU() $\times 1$ FC 3 Linear(256, 128) ReLU() $\times 1$ FC 4 Linear(256, 62) $\times 1$
+Table 1: Model architectures details of DNN used to train the FEMNIST dataset
+Layer Details Repetition Conv 1 Conv2d(3, 10, k=(5, 5), s=(1, 1)) ReLU() MaxPool2d(k=(2, 2)) $\times 1$ Conv 2 Conv2d(10, 20, k=(5, 5),s=(1, 1)) ReLU() MaxPool2d(k=(2, 2)) $\times 1$ FC 1 Linear(500, 256) ReLU() $\times 1$ FC 2 Linear(256, 10) $\times 1$
+Table 2: Model architectures details of 2-layered CNN used to train the CIFAR10 dataset
+Table 1 and 2 show the detailed model architecture used for the experiments, respectively.
+For the ResNet18 model used to train the CIFAR100 dataset, we exploited PyTorch implementation [NEURIPS2019_9015] by only replacing the final classification head from Linear(512, 1000) to Linear(512, 100).
+It is important to note that low temperatures may cause unstable convergence in large models (e.g., ResNet, ViT).
+We recommend incorporating normalization layers to mitigate this effect.
+We further note that investigating methods to minimize instability caused by low temperatures with respect to different model architectures or components is an interesting direction for future work.
+In real‐world deployments, clients often employ heterogeneous model architectures tailored to their specific applications and resource constraints.
+In this work, however, we restrict our evaluation to a homogeneous setting in which all clients share the same model architecture, thereby abstracting away variations in computational and communication capabilities.
+Extending our approach to support architectural heterogeneity and diverse client resources is an important avenue for future research, as it would more closely reflect practical federated learning scenarios.
 Experiment Details.
-Each case underwent local training for 300 federated training rounds, utilizing FedAvg McMahan et al.
-(2017) as the baseline federated learning framework.
+Each case underwent local training for 300 federated training rounds, utilizing FedAvg [mcmahan2017communication] as the baseline federated learning framework.
 FedAvg is a widely used framework for federated learning.
-In each round, the server selects 10 participating clients, employing Stochastic Gradient Descent (SGD) as the optimizer with a learning rate of 0.001 Ruder (2016).
+In each round, the server selects 10 participating clients, employing Stochastic Gradient Descent (SGD) as the optimizer with a learning rate of 0.001 [ruder2016overview].
 Each client trained its model for 10 local epochs, employing a batch size of 16, and utilized cross-entropy loss as the loss function.
 Unless specified otherwise, the experiment was conducted using the same hyperparameters as mentioned in this section.
-4.2 Average Client-side Accuracy
-We report the average client-side model accuracy obtained for models trained with different temperatures in Figure 4.
+Table 3 shows the default hyperparameters used for the experiments.
+We note that we applied a learning rate decay of $10^{-5}$ to enhance training stability.
+Although the effect was minimal and did not alter the trend of the main results, we recommend its application.
+Hyperparameters FEMNIST CIFAR10 CIFAR100 CKA Eval total clients 36 50 10 communication rounds 300 100 optimizer SGD learning rate 0.001 0.005 local epoch 10 10 participants per round 10 10 batch size 16 32 shard size | $\alpha$ N/A 200 $\alpha$=0.5 loss function Cross Entropy
+Table 3: Hyperparameters used for Section 4. $\alpha$ denotes the importance of the Dirichlet distribution.
+Finally, for the experiments, we exploited a server with four GeForce RTX 2080 Ti GPUs, Intel(R) Xeon(R) Silver 4210 CPU @ 2.20GHz, and 64GB RAM.
+4.2 Average Global Model Accuracy and Loss
+(a) FEMNIST-DNN
+(b) CIFAR10-CNN
+(c) CIFAR100-ResNet18
+Figure 7: Average global model accuracy for different temperatures used in training.
+Overall, lower temperatures show higher average model accuracy for federated learning systems.
+The shaded areas indicate the standard deviation across different runs.
+Note that the FEMNIST–DNN configuration exhibited relatively low variance compared to the other settings, largely due to its simpler architecture and lower overall complexity.
+(a) FEMNIST-DNN
+(b) CIFAR10-CNN
+(c) CIFAR100-ResNet18
+Figure 8: Average global model loss for different temperatures used in training.
+Overall, lower temperatures show a lower average model loss for federated learning systems.
+The shaded areas indicate the standard deviation across different runs.
+We report the average global model accuracy obtained for models trained with different temperatures in Figure 7.
 For all three datasets (and the three respective models used for each test case) we see an increasing accuracy trend observed for lower temperatures.
-Note that for the FEMNIST-DNN case with $T$=0.05, we noticed that the accuracy converges after 120 rounds; thus, perform an early exit for this case.
-One interesting aspect we can notice is that the accuracy for the $T$=0.05 case, which is the lowest temperature we tested for, does not show the best performance for the CIFAR100-ResNet18 experiment.
-While the $T$=0.05 case shows the best performance of FEMNIST-DNN and CIFAR10-CNN, for CIFAR100-ResNet18, we see an accuracy worse than the $T$=0.5 case.
+Note that for the FEMNIST-DNN case with TT=0.05, we noticed that the accuracy converges after 120 rounds; thus, perform an early exit for this case.
+One interesting aspect we can notice is that the accuracy for the TT=0.05 case, which is the lowest temperature we tested for, does not show the best performance for the CIFAR100-ResNet18 experiment.
+While the TT=0.05 case shows the best performance of FEMNIST-DNN and CIFAR10-CNN, for CIFAR100-ResNet18, we see an accuracy worse than the TT=0.5 case.
 This suggests that training with the lowest possible temperature value is not always ideal, rather this result highlights the importance of identifying a proper training temperature as we will further discuss in Section 5.
-A closer zoom into the plots suggests that the best-performing configuration can change over the training round and can vary for different datasets, but overall, we can notice that exploiting a temperature value lower than “1” in the training process can help improve the classification accuracy.
-The average accuracy for all federated learning clients after 300 training rounds presented in Table 1 confirm our observations, indicating that with $T$=0.05, the CIFAR10 case shows a 3.37% improvement in accuracy compared to the $T$=1 case.
+A closer zoom into the plots suggests that the best-performing configuration can change over the training round and can vary for different datasets, but overall, we can notice that exploiting a temperature value lower than 1 in the training process can help improve the classification accuracy.
+The average accuracy for all federated learning clients after 300 training rounds presented in Table 4 confirm our observations, indicating that with TT=0.05, the CIFAR10 case shows a 3.37% improvement in accuracy compared to the TT=1 case.
 Even for other cases, the use of a low-temperature value shows promising results in terms of accuracy.
@@
 Rather, our results suggest the need to carefully explore the fractional temperature space in the federated learning training process.
-We note that the performance for model performance and convergence speed tends to be similar regardless of federated learning hyperparameters (see Appendix D).
-4.3 Federated Learning Convergence Analysis
+We note that the performance for model performance and convergence speed tends to be similar regardless of federated learning hyperparameters (i.e., batch size, local epoch, number of participants).
+Next, we examine the validation loss of the global model.
+In Figure 8 we present the average loss observed at each federated learning training round.
+Again, we note that we perform early exit for FEMNIST-DNN at 120 rounds when TT=0.05 due to quick model convergence.
+We can notice here that the use of lower temperatures in the training phase generally helps to quickly lower the validation loss during the local training phase for federated learning.
+At the same time, when observing the CIFAR100-ResNet18 plots, we can see that the loss for TT=0.05 does not exhibit stable convergence, serving as a reason behind the lower accuracy reported for CIFAR100-ResNet18.
+From Table 4 where we report the average validation loss after 300 training rounds (120 for the FEMNIST-DNN @ TT=0.05), we can make similar observations.
+Especially, for FEMNIST-DNN, we see a six-fold improvement in loss compared to TT=1.
+For the CIFAR100-ResNet18 case, however, we notice that the lowest loss convergence is at TT=0.25 rather than 0.05.
+As noted in Section 3, this observation emphasizes the need to properly consider and explore fractional temperature values for client-side model training to balance between the boosting convergence and the instability during local training or the risk of overfitting.
+4.3 Federated Learning Convergence
 FEMNIST-DNN CIFAR10-CNN CIFAR100-ResNet18 Temperature ($T$) Accuracy (%) Loss Num.
@@
 of Round $T=4.0$ 73.60 (-5.06) 0.893 299 (0.44×) 37.58 (-3.76) 1.724 299 (0.84×) 28.05 (-0.20) 2.984 288 (0.69×) $T=2.0$ 77.20 (-1.46) 0.749 205 (0.64×) 40.35 (-0.99) 1.638 270 (0.93×) 27.89 (-0.36) 2.918 288 (0.69×) $T=1.0$ 78.66 (0.00) 0.684 132 (1.00×) 41.34 (0.00) 1.594 251 (1.00×) 28.25 (0.00) 2.918 200 (1×) $T=0.5$ 79.78 (+1.13) 0.654 88 (1.50×) 41.52 (+0.18) 1.551 161 (1.56×) 29.77 (+1.52) 2.875 178 (1.12×) $T=0.25$ 79.50 (+0.85) 0.650 57 (2.32×) 43.41 (+2.07) 1.532 132 (1.90×) 30.63 (+2.38) 2.997 152 (1.32×) $T=0.05$ 80.86 (+2.20) 0.643 22 (6.00×) 44.71 (+3.37) 1.528 106 (2.37×) 27.05 (-2.72) 4.228 239 (0.74×)
-Table 1: Average classification accuracy, model training loss, and number of federated learning training rounds needed to achieve the maximum accuracy of the $T$=4 cases for each dataset-model combination.
-Finally, in the last columns for each test case in Table 1 we present the number of federated learning training rounds needed to achieve a preset target accuracy for different temperature values used for model training.
-This number provides an understanding of how efficient the local training processes are when different temperature values are applied.
-Specifically, in this experiment, we configure the target accuracy to be the maximum accuracy achieved within 300 training rounds with $T$=4.
-For example, for the FEMNIST-DNN and CIFAR10-CNN cases, the maximum accuracy for $T$=4 was achieved at the 299th round, while CIFAR100-ResNet18 achieved such accuracy after 288 rounds.
-From the results in Table 1 we can notice that with $T$=0.05, the FEMNIST-DNN case achieves the target accuracy after only 22 training rounds.
-We see similar trends of low temperatures positively affecting the training efficiency throughout our results.
-This is from the general aggressiveness of training behaviors when applying lower temperature training.
-Given a single training sample, while high temperature-based training conservatively absorbs the available information, lower temperatures are more progressive in making model changes.
-Given the small datasets used for local training and the need to effectively deliver its knowledge to the server for aggregation, we see FLex&\&&Chill as an effective way to train client-side models in the federated learning process.
-4.4 Feature Space Similarity Analysis
-Figure 5: Test accuracy of the global model and feature space similarity in each layer between reference model trained with i.i.d.
-dataset and the global model trained with non-i.i.d.
-dataset with different temperatures applied during the local training.
-(a) $T$=0.05
-(b) $T$=1.0
-Figure 6: Client-wise feature space similarity matrix measured in penultimate layers by using CKA and inference accuracy in different temperatures (0.05, 1.0) and federated rounds (1, 5, 50, 100).
-(INIT: Initial dispatched model in the given round, AGG: Aggregated global model, REF: Reference model, ACC: Accuracy)
-To better understand the mechanism and rationale behind the performance of FLex&\&&Chill, we analyze the similarity of the feature space between global (aggregated) models trained with different temperatures against a reference model trained in an i.i.d.
-data configuration.
-Specifically, we employ the Centered Kernel Alignment (CKA) Kornblith et al.
-(2019) for computing the similarity of feature spaces, a widely used approach in evaluating model divergence in federated learning Collins et al.
-(2022); Luo et al.
-(2021).
-Note that a limitation of CKA measures is that they only report network similarity without considering training quality.
-Thus, we use a well-trained reference model from an i.i.d.
-environment to observe CKA variations of the penultimate and classification layers between the global models and the reference model.
-We note that the reference model is trained with $T$=1, while comparison models (e.g., global models from federated learning) are trained in non-i.i.d.
-environments with different $T$.
-In the experiments, we used a 2-layered CNN and the CIFAR-10 dataset as in Section 4.1, with the dataset distributed among 10 clients using a Dirichlet distribution Hsu et al.
-(2019).
-All clients participated in the federated training process.
-For detailed discussions on the experimental setup and CKA, please refer to Appendix E.
-As the accuracy and CKA plots for different layers in Figure 5 show, introducing the non-i.i.d.
-environment itself drops the model accuracy (73.86% $\xrightarrow{}$ 58.19%) at $T$=1, with the CKA of the penultimate/classification layers between the two models remaining at 0.7644/0.8214 after 300 rounds.
-Nevertheless, with $T$=0.05, despite the dataset disparity, the global model achieved a comparable accuracy of 72.76% with relatively higher CKA values (0.8898/0.9026) in both the penultimate and classification layers.
-We point out that improved gradient flows via applying low temperatures help models to learn similar local features (i.e., layers 1, 2), affecting the following layers (i.e., layers 3, 4) in the early stage of the learning process.
-As a result, applying low temperatures during local training effectively improves the overall accuracy performance in a non-i.i.d.
-data distribution setting and helps models learn similar feature spaces compared to the ideal i.i.d.
-data configuration.
-We now target to better understand how local model updates progress with different $T$.
-For brevity of the paper we only report the CKA at the penultimate layer, and plot all other cases in Appendix E.
-Specifically, Figure 6 plots the model accuracy (ACC) and measured CKA at the penultimate layer of a client’s local model with (i) the initially distributed model (INIT), (ii) the aggregated model (AGG), (iii) the ideal reference model trained with i.i.d data distribution (REF), and (iv) other client models while $T$=0.05 and 1.0.
-In Round 1, clients’ local models trained with $T$=1 show high feature space similarity against each other while failing to learn the features of the ideal reference model.
-With $T$=0.05, the CKA between local models are relatively low while showing high similarity with the reference model.
-After five rounds (and beyond), local models trained with both $T$=1.0 and $T$=0.05 showed high similarity amongst themselves and the initial/aggregated models, but the $T$=0.05 case continuously shows higher similarity with REF and reports higher accuracy.
-Hence, we argue that adapting low temperatures allows local models to learn effective and diverse features to quickly converge the federated learning process.
-4.5 Impact of Dataset Dispersion and Federated Learning on FLex&\&&Chill
-We now examine the impact of FLex&\&&Chill on different model training configurations.
+Table 4: Average classification accuracy, model test loss, and number of federated learning training rounds needed to achieve the maximum accuracy of the TT=4 cases for each dataset-model combination.
+Finally, in the final columns for each test case in Table 4, we present the number of federated learning training rounds needed to achieve a preset target accuracy for different temperature values used for model training.
+This quantity serves as our notion of “convergence speed” and provides an understanding of how efficiently the local training processes operate under different temperatures.
+Specifically, we define the target accuracy as follows: for each dataset–model configuration, we first run the training with $T=4$ for 300 rounds, and record the highest test accuracy that the global model (after server aggregation) achieves within those 300 rounds.
+The highest accuracy obtained by the global model trained with $T=4$ serves as the target accuracy for all temperature settings.
+For example, in the FEMNIST-DNN and CIFAR10-CNN experiments, the global model trained with $T=4$ reaches its maximum accuracy at round 299, whereas for CIFAR100-ResNet18, the corresponding maximum occurs at round 288.
+[…]
+[…]
+4.4 Model stability of low-temperature training during federated learning aggregation
+Figure 9: Accuracy drop observed before and after server-side model aggregation. The aggregation process of client models trained with lower temperatures better retains the individual knowledge.
+Figure 9 plots the changes in accuracy observed for both participating and non-participating clients at each federated learning round.
+Specifically, once the model is aggregated, each client computes its accuracy with both the pre-update and the post-update model and computes their difference.
+Note that pre-update models for clients participating in the round would have updated their local models (very recently) with their own data.
+As discussed in Section 3.3, low temperatures expedite local training even with a small amount of training data.
+Nevertheless, the effect of low-temperature training on the model generalization and aggregation stability of federated learning has not yet been thoroughly studied.
+To inspect the stability of federated learning aggregation for models trained with low temperatures, we measured the performance change of each client-side model before and after the server-side aggregation round.
+Thus, the pre-model performance will be high.
+However, since each model is trained aggressively over heterogeneous data, the aggregation process could puddle up this knowledge and the aggregated model may lose information.
+Ideally, with their local information properly aggregated, the performance drop will be minimal.
+Non-participating clients on the other hand, would hope to have an improved updated model of the federated learning process.
+As Figure 9 shows the models trained with low temperatures retain the knowledge better than the medium and hot models for participating clients (solid plots groups at the bottom).
+In the early rounds, the models with high temperatures show less performance drop, but the fast learning speed of the low-temperature model amortizes the performance drop, which results in faster learning of the aggregated model.
+Furthermore, we can also notice that the aggregation process does not interfere with the performance of non-participating clients (grouped on top), implying that federated learning with low-temperature training preserves the generalization power.
+4.5 Impact of Dataset Dispersion and Federated Learning on FLex&\&Chill
+We now examine the impact of FLex&\&Chill on different model training configurations.
 Specifically, using the CIFAR10-CNN test case, we examine two additional configurations: (i) centralized learning with i.i.d.
@@
 By comparing these results with the non-i.i.d.
-federated learning results, we target to understand the impact of FLex&\&&Chill on federated learning and dataset disparity in greater detail.
-As Table 2 shows, applying FLex&\&&Chill, thus a lower $T$ does not induce a positive impact on the performance of centralized learning scenarios.
-In fact, for centralized learning, controlling $T$ for the training process does not show a noticeable impact.
+federated learning results, we target to understand the impact of FLex&\&Chill on federated learning and dataset disparity in greater detail.
+As Table 5 shows, applying FLex&\&Chill, thus a lower $T$ does not induce a positive impact on the performance of centralized learning scenarios.
+In fact, for centralized learning, controlling $T$ for the training process does not show a positive impact.
 We see the same phenomena for the federated learning system with i.i.d.
@@
 (%) Loss 4.0 58.05 1.152 48.22 1.473 37.58 1.724 2.0 63.27 1.041 51.90 1.403 40.35 1.638 1.0 65.18 1.044 54.66 1.325 41.34 1.594 0.5 64.19 1.030 53.19 1.320 41.52 1.551 0.25 63.88 1.092 54.64 1.265 43.41 1.532 0.05 60.98 1.227 47.60 1.701 44.71 1.528
-Table 2: Average inference accuracy and model training loss for CIFAR10-CNN with different configurations.
-The impact of FLex&\&&Chill is prominent for federated learning scenarios with non-i.i.d.
-datasets, whereas its impact is less than elsewise.
+Table 5: Average inference accuracy and model test loss for CIFAR10-CNN with different configurations.
+The impact of FLex&\&Chill is prominent for federated learning scenarios with non-i.i.d.
+datasets, whereas its impact is less than else-wise.
+Figure 10: Average entropy before aggregation (Before Agg) and after aggregation (After Agg) for different data disparity configurations.
+In federated learning, where non-i.i.d.
+environments are common, the loss of information during model aggregation can be a significant factor contributing to performance degradation.
+Figure 10 shows the entropy of the model before and after server-side aggregation in two different settings: non-i.i.d.
+and i.i.d.
+data distribution.
+In the federated learning setup, the model’s entropy increases after aggregation, indicating that the probability distribution produced by the model has become smoother, with the smoothing effect more pronounced in non-i.i.d.
+environments.
+By applying a lower $T$, we carefully argue that the smoothed probability distribution can be sharpened, which appears to enhance both convergence and final performance.
 We do point out that holding i.i.d.
 datasets in federated learning will show overall higher performance compared to non-i.i.d.
-configurations (regardless from $T$), but is not a practical case to assume when deploying federated learning systems.
+configurations (regardless of $T$), but is not practical to assume when deploying federated learning systems.
 Rather, as we discussed, the impact of exploiting low training temperatures is beneficial when dealing with non-i.i.d.
 datasets.
-4.6 Model stability of low-temperature training during federated learning aggregation
-Figure 7: Accuracy drop observed before and after server-side model aggregation. The aggregation process of client models trained with lower temperatures better retains the individual knowledge.
-Figure 7 plots the changes in accuracy observed for both participating and non-participating clients at each federated learning round.
-Specifically, once the model is aggregated, each client computes its accuracy with both the pre-update and the post-update model and computes their difference.
-Note that pre-update models for clients participating in the round would have updated their local models (very recently) with their own data.
-As discussed in Section 3, low temperatures expedite local training even with a small amount of training data.
-Nevertheless, the effect of low-temperature training on the model generalization and aggregation stability of federated learning has not yet been thoroughly studied.
-To inspect the stability of federated learning aggregation for models trained with low temperatures, we measured the performance change of each client-side model before and after the server-side aggregation round.
-Thus, the pre-model performance will be high.
-However, since each model is trained aggressively over heterogeneous data, the aggregation process could puddle up this knowledge and the aggregated model may lose information.
-Ideally, with their local information properly aggregated, the performance drop will be minimal.
-Non-participating clients on the other hand, would hope to have an improved updated model of the federated learning process.
-As Figure 7 shows the models trained with low temperatures retain the knowledge better than the medium and hot models for participating clients (solid plots groups at the bottom).
-In the early rounds, the models with high temperatures show less performance drop, but the fast learning speed of the low-temperature model amortizes the performance drop, which results in faster learning of the aggregated model.
-Furthermore, we can also notice that the aggregation process does not interfere with the performance of non-participating clients (grouped on top), implying that federated learning with low-temperature training preserves the generalization power.
+(a) Severe heterogeneity. ($\alpha$=0.1)
+(b) Moderate heterogeneity. ($\alpha$=0.5)
+(c) Weak heterogeneity. ($\alpha$=1.0)
+Figure 11: Test accuracy and test loss over federated learning rounds with varying the degree of data disparity and the training temperature on the CIFAR10 dataset.
+Relatively moderate data disparity shows better training quality with stability.
+The shaded areas indicate the standard deviation across different runs.
+To assess the robustness of FLex&\&Chill under varying degrees of data heterogeneity, we partition the dataset using a Dirichlet distribution [hsu2019measuring].
+A lower Dirichlet parameter $\alpha$ corresponds to a higher degree of heterogeneity, while a higher $\alpha$ indicates a more homogeneous setting.
+We consider $\alpha=\{0.1,0.5,1.0\}$ to represent severe, moderate, and weak heterogeneity levels, respectively.
+Figure 11 presents the global model accuracy over 300 training rounds.
+As depicted in the figure, a lower temperature consistently leads to superior model performance across all settings.
+4.6 Feature Space Similarity
+Figure 12: Test accuracy of the global model and feature space similarity in each layer between reference model trained with i.i.d.
+dataset and the global model trained with non-i.i.d.
+dataset with different temperatures applied during the local training.
+(a) TT=0.05
+(b) TT=1.0
+Figure 13: Client-wise feature space similarity matrix measured in penultimate layers by using CKA and inference accuracy in different temperatures (0.05, 1.0) and federated rounds (1, 5, 50).
+(INIT: Initial dispatched model in the given round, AGG: Aggregated global model, REF: Reference model, ACC: Accuracy)
+To better understand the mechanism and rationale behind the performance of FLex&\&Chill, we analyze the similarity of the feature space between global (aggregated) models trained with different temperatures against a reference model trained in an i.i.d.
+data configuration.
+Specifically, we employ the Centered Kernel Alignment (CKA) [kornblith2019similarity] for computing the similarity of feature spaces, a widely used approach in evaluating model divergence in federated learning [collins2022fedavg, luo2021no, shin2024effective].
+CKA is a similarity measure between two different neural networks by measuring the similarity of the network representation for the given input data.
+Let $X$ and $Y$ be the features generated from two different neural networks, we can calculate the CKA by Eq 6.
+As we are assuming a federated learning scenario, the clients share the same neural network architecture.
+Thus, the dimensions of the two features from each corresponding layer are identical, allowing for a direct comparison to be possible (detailed discussion provided in Appendix).
+We note that in our experiments, we exploited the linear CKA to measure the similarity of the feature spaces between two neural networks.
+[…]
+Note that a limitation of CKA measures is that they only report network similarity without considering training quality.
+Thus, we use a well-trained reference model from an i.i.d.
+environment to observe CKA variations of the penultimate and classification layers between the global models and the reference model.
+We note that the reference model is trained with TT=1, while comparison models (e.g., global models from federated learning) are trained in non-i.i.d.
+environments with different $T$.
+In the experiments, we used a 2-layered CNN and the CIFAR-10 dataset as in Section 4.1, with the dataset distributed among 10 clients using a Dirichlet distribution with $\alpha=0.5$, and all clients participated in the federated training process.
+Temperature ($T$) Accuracy (%) Layer 1 CKA Layer 2 CKA Layer 3 CKA Layer 4 CKA 4.00 49.72 0.6539 0.5632 0.7149 0.7661 2.00 53.82 0.6792 0.6872 0.7417 0.7963 1.00 58.19 0.7419 0.6997 0.7644 0.8214 0.50 63.48 0.9322 0.7459 0.8181 0.8559 0.25 66.85 0.9569 0.8161 0.8418 0.8830 0.05 72.76 0.9390 0.9059 0.8898 0.9026 Reference Model (TT=1.00, i.i.d.) 73.86 1.0000 1.0000 1.0000 1.0000
+Table 6: Final accuracy and CKA values of 2 layer CNN measured on the CIFAR10 dataset in Figure 12.
+[…]
+We now aim to better understand how local model updates progress with different $T$.
+For brevity of the paper, we only report the CKA at the penultimate layer.
+Specifically, Figure 13 plots the model accuracy (ACC) and measured CKA at the penultimate layer of a client’s local model with (i) the initially distributed model (INIT), (ii) the aggregated model (AGG), (iii) the ideal reference model trained with i.i.d.
+data distribution (REF), and (iv) other client models while TT=0.05 and 1.0.
+In Round 1, clients’ local models trained with TT=1 show high feature space similarity against each other while failing to learn the features of the ideal reference model.
+With TT=0.05, the CKA between local models is relatively low while showing high similarity with the reference model.
+After five rounds (and beyond), local models trained with both TT=1.0 and TT=0.05 showed high similarity amongst themselves and the initial/aggregated models, but the TT=0.05 case continuously shows higher similarity with REF and reports higher accuracy.
+Hence, we argue that adapting low temperatures allows local models to learn effective and diverse features to converge the federated learning process quickly.
+4.7 FLex&\&Chill with Advanced FL Schemes
+(a) FedProx
+(b) SCAFFOLD
+(c) FedBN
+Figure 14: Test accuracy and test loss on the CIFAR10 dataset with varying training temperatures with advanced federated learning schemes.
+Overall, models trained with lower temperatures showed superior performance.
+SCAFFOLD FedProx FedBN Temperature ($T$) Accuracy (%) Loss Num.
+of Round Accuracy (%) Loss Num.
+of Round Accuracy (%) Loss Num.
+of Round $T=4.0$ 32.32 (-11.60) 1.935 300 (0.39×) 29.95 (-13.41) 1.989 300 (0.76×) 43.77 (-6.60) 1.562 292 (0.61×) $T=2.0$ 38.40 (-5.52) 1.729 222 (0.55×) 37.42 (-5.94) 1.739 300 (0.76×) 47.32 (-3.05) 1.478 212 (0.84×) $T=1.0$ 43.92 (+0.00) 1.562 121 (1.00×) 43.36 (+0.00) 1.555 227 (1.00×) 50.37 (+0.00) 1.397 178 (1.00×) $T=0.5$ 47.90 (+3.98) 1.464 95 (1.27×) 46.99 (+3.63) 1.457 202 (1.12×) 53.81 (+3.44) 1.312 143 (1.24×) $T=0.25$ 52.02 (+8.10) 1.324 45 (2.69×) 50.78 (+7.42) 1.356 143 (1.59×) 55.95 (+5.58) 1.244 88 (2.02×) $T=0.05$ 60.10 (+16.18) 1.127 13 (9.31×) 61.15 (+17.79) 1.106 49 (4.63×) 56.31 (+5.94) 1.235 82 (2.17×)
+Table 7: Average classification accuracy, model test loss, and number of federated learning training rounds needed to achieve the maximum accuracy of the TT=4 cases for FedProx, SCAFFOLD, and FedBN.
+Finally, as discussed in Section 3.4, FLex&\&Chill is designed as an orthogonal enhancement applicable to a wide range of federated learning algorithms.
+To further demonstrate this property, we additionally trained the 2-layer CNN with FedBN [li2021fedbn] alongside FedProx [li2020federated] and SCAFFOLD [karimireddy2020scaffold], which represent more advanced optimization schemes compared to naive FedAvg.
+Figure 14 and Table 7 report the global model accuracy over 300 federated learning rounds across different temperature settings.
+As detailed earlier in Table 2, the model used in our main experiments do not include batch normalization; thus, we use a variant of the 2-layer CNN with batch normalization when evaluating FedBN.
+This architectural difference leads to slight numerical deviations compared to other experiments.
+Nevertheless, the overall trend remains consistent with lower temperatures (0.05, 0.25, 0.5) outperforming the standard or higher temperatures (1.0, 2.0, 4.0) across all three algorithms (i.e., FedProx, SCAFFOLD, and FedBN).
+This result further highlights the orthogonality and broad applicability of FLex&\&Chill across diverse federated optimization strategies.
+4.8 Experiments on Transformer Architecture
+Figure 15: Test accuracy and test loss over federated learning rounds with varying the training temperature on the Vision Transformer and CIFAR10 dataset configuration.
+The shaded areas indicate the standard deviation across different runs.
+Layer Details Repetition Patch Embedding Conv2d(3, 256, k=(4, 4), s=(4, 4)) Flatten() Add CLS token + Positional Embedding $\times 1$ Transformer Encoder Layer Multi‐head Attention (d_model=256, nhead=4) MLP (dim_feedforward=1024, activation=GELU, dropout=0.1) $\times 3$ LayerNorm LayerNorm(normalized_shape=256) $\times 1$ Classification Head (FC) Linear(256, 10) $\times 1$
+Table 8: Model architecture details of the ViT used for CIFAR-10 classification
+We assess our method’s applicability to transformer architectures by running FedAvg on a Vision Transformer (ViT) [dosovitskiy2020image] trained on CIFAR-10.
+Five temperature values ($T\in\{0.25,0.5,1.0,2.0,4.0\}$) are compared to evaluate their effects on convergence speed and final accuracy.
+Table 8 details the ViT configuration used, and Figure 15 shows test accuracy and loss over communication rounds.
+As with our convolutional experiments, lower temperatures accelerate convergence and improve final accuracy in the transformer setting.
+This behavior arises because ViTs, as with most classification models, use a softmax output layer and cross-entropy loss [lee2024tazza], so temperature scaling similarly sharpens probability distributions and amplifies meaningful gradients.
+4.9 Experiments on Time-Series Dataset
+Figure 16: Test accuracy and test loss over federated learning rounds with varying the training temperature on the 1D-CNN and UCI-HAR dataset configuration.
+The shaded areas indicate the standard deviation across different runs.
+Layer Details Repetition Conv Block 1 Conv1d(1, 16, k=7, s=1, p=3) BatchNorm1d(16) ReLU() MaxPool1d(k=2) $\times 1$ Conv Block 2 Conv1d(16, 32, k=5, s=1, p=2) BatchNorm1d(32) ReLU() MaxPool1d(k=2) $\times 1$ Conv Block 3 Conv1d(32, 64, k=5, s=1, p=2) BatchNorm1d(64) ReLU() MaxPool1d(k=2) $\times 1$ Conv Block 4 Conv1d(64, 128, k=3, s=1, p=1) BatchNorm1d(128) ReLU() AdaptiveAvgPool1d(output=1) $\times 1$ Classification Head (FC) Linear(128, 6) $\times 1$
+Table 9: Model architecture details of the 1D CNN HAR classification.
+We assess our method’s applicability to IoT- and mobile-oriented federated scenarios [seo2025fedunet, xiao2025federated, xing2022efficient] by running FedAvg on a 1-dimensional CNN trained on an Inertial Measurement Unit (IMU) sensor dataset (UCI-HAR) [ucihar].
+This time-series setting is closely related to recent federated learning studies on human activity recognition and multitask time-series classification [lee2025gmt], which employ contrastive learning or feature-based distillation to enhance representation quality under heterogeneous data [xiao2025federated, xing2022efficient].
+These approaches primarily redesign the training objective (e.g., by adding contrastive or distillation losses), while keeping the underlying softmax temperature fixed.
+In contrast, FLex&\&Chill is orthogonal to such methods as it modifies the local training dynamics by applying low-temperature scaling at the softmax layer, and can, in principle, be combined with these feature-centric objectives.
+Five temperature values ($T\in\{0.25,0.5,1.0,2.0,4.0\}$) are compared to evaluate their effects on convergence speed and final accuracy.
+Table 9 details the 1D-CNN configuration used, and Figure 16 shows test accuracy and loss over communication rounds.
+As with our image-based experiments, lower temperatures accelerate convergence and improve final accuracy in the time-series sensor data classification task.
 5 DISCUSSION
+Based on our experiences in designing FLex&\&Chill, we outline a set of interesting discussion points that point towards meaningful future research directions.
 $\bullet$ Identifying the optimal temperature.
@@
 However, since the softmax layer training temperature can be considered a hyperparameter, optimal temperatures can vary depending on the applications and model architectures.
-Thus, like other hyperparameters used in model training (e.g., batch size, local epoch count), identifying an optimal temperature value is challenging to analytically solve, being beyond the scope of this paper.
+Thus, like other hyperparameters used in model training (e.g., batch size, local epoch), identifying an optimal temperature value is challenging to analytically solve, being beyond the scope of this paper.
 Nevertheless, based on our experiences, we can offer heuristics for selecting temperatures regarding the model capacity and task complexity.
@@
 While our paper focuses on understanding the advantages of maintaining a constant low temperature for the training process, this may not be the ideal approach when targeting maximum performance.
-In reality, varying temperatures can be applied in the training process with respect to their significance Shin et al.
-(2022); Zeng et al.
-(2021).
+In reality, varying temperatures can be applied in the training process with respect to their significance [shin2022fedbalancer, zeng2021mercury].
 That would mean that the training process can exploit lower temperatures for the samples that are “harder” to learn, while higher temperatures are used for those easier to understand.
 Alternatively, a more course-grain approach of dynamically adjusting the temperature on a per-federated learning round-basis can be another meaningful approach to adapt towards changing local data characteristics.
-$\bullet$ Supporting heterogeneous federated learning.
-While many federated learning frameworks assume homogeneous devices, recent work points out that this is not always the case Park and Ko (2023).
-In a federated learning network consisting of devices with heterogeneous computing/network capabilities, the issue of “stragglers”, which take longer to complete local training, becomes important to address.
-Leveraging insights gleaned from our empirical observations, applying lower temperatures for potential stranglers can be an interesting way of uniforming the local training duration.
-This would require a prior analysis on the conditions expected at each device and a global consensus on the temperature configuration.
-We find this as an interesting topic that requires further investigation.
+$\bullet$ Supporting computation heterogeneity.
+While many federated learning frameworks assume homogeneous devices, recent work points out that this is not always the case [PARK23fedhm].
+In a federated learning consisting of devices with heterogeneous computing and network capabilities, the issue of “stragglers,” which take longer to complete local training, becomes important to address [shin2024effective].
+Leveraging insights from our empirical observations, applying lower temperatures for potential stragglers can be an interesting solution; it would require prior analysis of each device’s expected conditions and a global temperature configuration consensus.
+We find this an interesting direction of future research.
+Future work could profile each client’s compute and network resources to assign and negotiate personalized temperatures, lower on less constrained devices to speed convergence, and relatively higher on powerful clients to reduce overhead.
+$\bullet$ Developmental Tendencies and Challenges.
+We identify several developmental tendencies and challenges that can guide future research.
+Beyond using a fixed temperature, dynamic or client-adaptive temperature scheduling represents a natural next step, particularly as our results indicate that low-temperature training preserves client-specific knowledge more effectively (as also suggested by the smaller post-aggregation accuracy drop in Figure 9).
+This makes FLex&\&Chill a promising foundation for integration with personalization frameworks and heterogeneous model deployments.
+At the same time, important challenges remain, including determining the optimal temperature for different model capacities and task complexities, ensuring stability for large networks under very low temperatures, and understanding performance limits under extreme non-i.i.d.
+scenarios such as fully disjoint label distributions.
+Addressing these open issues will further clarify the scope and robustness of temperature-based training in federated learning.
 6 CONCLUSION
-This research introduces a novel training approach for federated learning systems, FLex&\&&Chill, aimed at enhancing convergence and improving model performance by exploiting a temperature scaling method termed logit chilling in federated learning local model training operations.
-Specifically, our work revolves around the implications of employing fractional temperature values ($T\in(0,1)$) during the model training process.
-Our findings, derived from an extensive set of evaluations using three datasets and three widely applied neural networks for federated learning, demonstrate that exploiting the FLex&\&&Chill approach in the client-side model training process holds the potential to expedite model convergence by inducing more pronounced changes despite a limited number of local samples.
-Consequently, we show that by doing so, we can also achieve an overall higher inference accuracy.
-While the impact of lower temperature varies with respect to the neural network architecture and dataset, our evaluations reveal up to 6.00$\times$ improvement in federated learning model convergence speed and 3.37% improvement in inference accuracy.
-We see this work as a first step into devising a more fine-tuned (and optimal) training mechanism for federated learning systems.
-Appendix A Gradient of softmax with varying temperature
+[…]
+Acknowledgements
+[…]
+Appendix
+$\bullet$ Gradient of Softmax with Varying Temperature
 Proof.
-Let cross-entropy loss for softmax as $\mathcal{L}=-\sum_{j=1}^{C}y_{j}\log{(p_{j})}$, where $C$ is the number of classes and $y_{j}$ and $p_{j}$ corresponds to the hard label (i.e., onehot-encoded label) and probability for given class.
+Let cross-entropy loss for softmax as $\mathcal{L}=-\sum_{j=1}^{C}y_{j}\log{(p_{j})}$ where $C$ is the number of classes and $y_{j}$ and $p_{j}$ corresponds to the hard label (i.e., onehot-encoded label) and probability for given class.
 Here, the probability $p_{j}$ is computed by exploiting the softmax with varying temperatures which can be formalized as follows.
@@
 Here, $T$ is the temperature, and $z_{j}$ is the logit used as the input of the softmax function.
-$\begin{split}\frac{\partial\mathcal{L}}{\partial z_{i}}&=-\sum_{j=1}^{C}\frac{\partial y_{j}\log{(p_{j})}}{\partial z_{i}}\\ &=\sum_{j=1}^{C}y_{j}\frac{\partial\log{(p_{j})}}{\partial z_{i}}\\ &=-\sum_{j=1}^{C}y_{j}\frac{\partial\log{(p_{j})}}{\partial z_{i}}\\ &=-\sum_{j=1}^{C}y_{j}\frac{1}{p_{j}}\frac{\partial p_{j}}{\partial z_{i}}\\ &=-\frac{y_{i}}{p_{i}}\frac{\partial p_{i}}{\partial z_{i}}-\sum_{j\neq i}^{C}\frac{y_{j}}{p_{j}}\frac{\partial p_{j}}{\partial z_{i}}\\ &=-\frac{y_{i}}{p_{i}}\frac{1}{T}p_{i}(1-p_{i})-\sum_{j\neq i}^{C}\frac{y_{j}}{p_{j}}(-p_{j}p_{i})\\ &=-\frac{y_{i}}{T}+\frac{1}{T}y_{i}p_{i}+\frac{1}{T}\sum_{j\neq i}^{C}y_{j}p_{i}\\ &=-\frac{y_{i}}{T}+\frac{1}{T}\sum_{j=1}^{C}y_{j}p_{i}\\ &=-\frac{y_{i}}{T}+\frac{1}{T}p_{i}\sum_{j=1}^{C}y_{j}\\ &=\frac{1}{T}(p_{i}-y_{i})\\ \therefore\frac{\partial\mathcal{L}}{\partial z_{i}}&=\frac{1}{T}(p_{i}-y_{i})\\ &=\frac{1}{T}(\frac{e^{z_{i}/T}}{\sum_{j}e^{z_{j}/T}}-y_{i})\end{split}$
-Appendix B Experimental Details
-B.1 Model Architecture
-Layer Details Repetition FC 1 Linear(784, 512) ReLU() ×$\times 1$ FC 2 Linear(512, 256) ReLU() ×$\times 1$ FC 3 Linear(256, 128) ReLU() ×$\times 1$ FC 4 Linear(256, 62) ×$\times 1$
-Table 3: Model architectures details of DNN used to train the FEMNIST dataset
-Layer Details Repetition Conv 1 Conv2d(3, 10, k=(5, 5), s=(1, 1)) ReLU() MaxPool2d(k=(2, 2)) ×$\times 1$ Conv 2 Conv2d(10, 20, k=(5, 5),s=(1, 1)) ReLU() MaxPool2d(k=(2, 2)) ×$\times 1$ FC 1 Linear(500, 256) ReLU() ×$\times 1$ FC 2 Linear(256, 10) ×$\times 1$
-Table 4: Model architectures details of 2-layered CNN used to train the CIFAR10 dataset
-For the ResNet18 model used to train the CIFAR100 dataset, we exploited PyTorch implementation Paszke et al.
-[2019] by only replacing the final classification head from Linear(512, 1000) to Linear(512, 100).
-B.2 Hyperparameters
-Hyperparameters FEMNIST-DNN CIFAR10-CNN CIFAR100-ResNet18 CKA Evaluation total clients 36 50 10 communication rounds 300 100 optimizer SGD learning rate 0.001 0.005 local epoch 10 10 participants per round 10 10 batch size 16 32 shard size | $\alpha$ N/A 200 $\alpha$=0.5 loss function Cross Entropy
-Table 5: Hyperparameters used for Section 4. $\alpha$ denotes the importance of the Dirichlet distribution.
-B.3 Data distribution
-(a) FEMNIST
-(b) CIFAR10
-(c) CIFAR100
-Figure 8: Visualization of the distribution of training data used in FEMNIST, CIFAR10-CNN and CIFAR100-ResNet experiments. Best viewed in color.
-B.4 Hardware
-For the experiments, we exploited server with four GeForce RTX 2080 Ti GPUs, Intel(R) Xeon(R) Silver 4210 CPU @ 2.20GHz, and 64GB RAM.
-Appendix C Orthogonality of FLex&\&&Chill
-C.1 FLex&\&&Chill As An Algorithm
-We formulate the operations of FLex&\&&Chill using Algorithm 1.
-Here, typical federated learning operations take place at both the server and clients except for the fact that the server determines a target $T$ at each round and local model training operations are performed with $T$.
-Since FLex&\&&Chill only manipulates the local training process which inherently provides orthogonality to our proposed method, it is extremely easy to integrate with existing federated learning frameworks.
-Algorithm 1 FLex&\&&Chill
-Data ($D_{1},D_{2},...,D_{N}$) where $D_{i}$ is the user $i$’s local data.
-Model ($w^{1}_{t},w^{2}_{t},...,w^{N}_{t}$) where $w_{t}^{i}$ is the user $i$’s model in $t$-th federated round.
-Parameter: Temperature $T\in(0,1)$
-1:Server executes:
-2:Initialize $w_{0}$
-3:for each round $t=1,2,...,R$ do
-4:    $S_{t}\xleftarrow{}$ (random set of $m$ clients)
-5:    for $k\in S_{t}$ do
-6:         $w_{t}^{k}\xleftarrow{}$ ClientUpdate($w_{t},T$)
-7:    end for
-8:    for $k\in S_{t}$ do
-9:         $w_{t+1}\xleftarrow{}$ Aggregation($w_{t}^{k}$)
-10:    end for
-11:end for
-12:ClientUpdate($w_{t}$, $T$):
-13:$w_{t}^{k}\xleftarrow{}w_{t}$
-14:for data, label $\in D_{k}$ do
-15:    $\mathbf{z}=w_{t}^{k}(data)/T$
-16:    $l(\mathbf{z})=loss(\mathbf{z};label)$
-17:    $w_{t}^{k}\xleftarrow{}LocalUpdate(w_{t}^{k},l(\mathbf{z}))$
-18:end for
-19:Return $w_{t}^{k}$ to server
-C.2 FLex&\&&Chill with Advanced FL Schemes
-As mentioned in Section 3.2, FLex&\&&Chill is an orthogonal approach to enhance the federated learning process.
-To demonstrate this, we trained the 2-layer CNN with FedProx Li et al.
-[2020] and SCAFFOLD Karimireddy et al.
-[2020] representing more recently proposed FL schemes compared to FedAvg used in this work.
-Hyperparameters FedProx SCAFFOLD total clients 50 communication rounds 300 optimizer SGD learning rate 0.001 local epoch 5 participants per round 5 batch size 32 $\alpha$ 0.5 loss function Cross Entropy $\mu$ 0.01 N/A
-Table 6: Default hyperparameters used in this section. $\alpha$ denotes the importance of the Dirichlet distribution.
-(a) FedProx
-(b) SCAFFOLD
-Figure 9: Test accuracy and test loss on the CIFAR10 dataset with varying training temperatures with advanced federated learning schemes.
-Overall, models trained with lower temperatures showed superior performance.
-Appendix D Experiment on hyperparameters
-In this section, we explore the effect of other federated learning hyperparameters (i.e., batch size, number of local epochs, number of participants, and degree of data disparity).
-Through this set of experiments, we observed that applying low temperatures during local training improves test accuracy and convergence speed, being independent from hyperparameters.
-We note that the experiments were done by applying the default setting except for the control factors.
-D.1 Batch size
-(a) Batch size 8
-(b) Batch size 16
-(c) Batch size 32
-(d) Batch size 64
-(e) Batch size 128
-Figure 10: Test accuracy and test loss over federated learning round with varying the batch size and the training temperature on the CIFAR10 dataset.
-Relatively small batch size shows better training quality with stability.
-D.2 Number of participants
-(a) Number of participants 5
-(b) Number of participants 10
-(c) Number of participants 15
-Figure 11: Test accuracy and test loss over federated learning round with varying the number of participants and the training temperature on the CIFAR10 dataset.
-When relatively many clients participate in the training, it shows better training quality with stability
-D.3 Number local epochs
-(a) Number of local epoch 1
-(b) Number of local epoch 5
-(c) Number of local epoch 10
-(d) Number of local epoch 15
-Figure 12: Test accuracy and test loss over federated learning round with varying the number of local epochs and the training temperature on the CIFAR10 dataset.
-Relatively many updates in the local training phase show better training quality with stability
-D.4 Degree of data disparity
-(a) $\alpha$=0.1
-(b) $\alpha$=0.5
-(c) $\alpha$=1.0
-Figure 13: Test accuracy and test loss over federated learning round with varying the degree of data disparity and the training temperature on the CIFAR10 dataset.
-Relatively moderate data disparity shows better training quality with stability
-Hyperparameters Appendix D total clients 50 communication rounds 100 optimizer SGD learning rate 0.001 local epoch 5 participants per round 5 batch size 32 $\alpha$ 0.5 loss function Cross Entropy
-Table 7: Default hyperparameters used in this section. $\alpha$ denotes the importance of the Dirichlet distribution.
-Appendix E Details of Centered Kernel Alignment (CKA)
-E.1 Background
-Centered kernel alignment (CKA) is a method to compute the similarity between two different neural networks by measuring the similarity of the network representation for the given input data.
-Let $X$ and $Y$ be the features generated from two different neural networks, we can calculate the CKA by Eq 3.
-As we are assuming a federated learning scenario, the clients share the same neural network architecture.
-Thus, the dimensions of the two features from each corresponding layer are identical, allowing for a direct comparison to be possible.
-We note that in our experiments, we exploited the linear CKA to measure the similarity of the feature spaces between two neural networks.
-$\begin{split}CKA(X,Y)&=\frac{||Y^{T}X||^{2}_{F}}{||X^{T}X||_{F}||Y^{T}Y||_{F}}\\ \end{split}$ (3)
-E.2 Limitation of Centered Kernel Alignment evaluation in federated learning
-Here, we discuss our rationale for measuring CKA between a client model participating in the federated learning round and a reference model trained using i.i.d.
-configurations.
-As mentioned in Section 4.4, the feature space similarity itself only reports the similarity of neural networks and does not assure high quality (i.e., accuracy) of the trained models.
-Figure 14 illustrates the possible cases that can observed in a federated learning scenario.
-We use these four cases to understand how locally trained models converge to the reference model.
-Here, the orange line indicates the update directions of local models from the federated learning training phase (i.e., diverging from the initially shared aggregated model presented as the black dot) and the red dotted line represents the average of these locally updated models that occur at the server.
-The star represents the ideal model in the representation space with high accuracy.
+$\frac{\partial\mathcal{L}}{\partial z_{i}}$ $=-\sum_{j=1}^{C}\frac{\partial y_{j}\log{(p_{j})}}{\partial z_{i}}$
+$=\sum_{j=1}^{C}y_{j}\frac{\partial\log{(p_{j})}}{\partial z_{i}}$
+$=-\sum_{j=1}^{C}y_{j}\frac{\partial\log{(p_{j})}}{\partial z_{i}}$
+$=-\sum_{j=1}^{C}y_{j}\frac{1}{p_{j}}\frac{\partial p_{j}}{\partial z_{i}}$
+$=-\frac{y_{i}}{p_{i}}\frac{\partial p_{i}}{\partial z_{i}}-\sum_{j\neq i}^{C}\frac{y_{j}}{p_{j}}\frac{\partial p_{j}}{\partial z_{i}}$
+$=-\frac{y_{i}}{p_{i}}\frac{1}{T}p_{i}(1-p_{i})-\sum_{j\neq i}^{C}\frac{y_{j}}{p_{j}}(-p_{j}p_{i})$
+$=-\frac{y_{i}}{T}+\frac{1}{T}y_{i}p_{i}+\frac{1}{T}\sum_{j\neq i}^{C}y_{j}p_{i}$
+$=-\frac{y_{i}}{T}+\frac{1}{T}\sum_{j=1}^{C}y_{j}p_{i}$
+$=-\frac{y_{i}}{T}+\frac{1}{T}p_{i}\sum_{j=1}^{C}y_{j}$
+$=\frac{1}{T}(p_{i}-y_{i})$
+$\therefore\frac{\partial\mathcal{L}}{\partial z_{i}}$ $=\frac{1}{T}(p_{i}-y_{i})$
+$=\frac{1}{T}(\frac{e^{z_{i}/T}}{\sum_{j}e^{z_{j}/T}}-y_{i})$
+$\bullet$ Convergence Analysis
+Stochastic Optimization: SGD
+We first start the convergence analysis with basic SGD optimizer [dean2012large], incorporating both temperature scaling $T$ and learning rates $\eta$.
+Assumption 1. $L$-smoothness: local loss function $F_{k}(w)$ is L-smooth.
+[…]
+Assumption 2. $\mu$-strong convexity: loss function $F(w)$ is $\mu$-strong convex.
+[…]
+Assumption 3.
+Bounded Variance: We assume that the gradient of the local loss function ($\nabla F_{k}(w)$) has finite bounded variance which means that the difference between the local gradient and the global gradient is bounded.
+1. Local update
+$w_{k}^{(t+1)}=w_{k}^{(t)}-\eta\nabla F_{k}(w_{k})^{(t)}$
+2. Global model update
+$w^{(t+1)}=\sum_{k=1}^{K}p_{k}w_{k}^{(t+1)}$
+3. Difference between local and global update
+$\mathbb{E}[F(w^{(t+1)})-F(w^{*})]\leq$
+$(1-\eta\mu)\mathbb{E}[F(w^{(t)})-F(w^{*})]+\frac{L\eta^{2}\sigma^{2}}{2}$
+[…]
+4. Incorporating temperature scaling
+Recall that we can rewrite the local gradient as follows (see Section 3-3.1):
+$\nabla F_{k}(w)=\frac{1}{T}\nabla F_{k}(w;T=1)$
+[…]
+$(1-\eta\mu/T)\mathbb{E}[F(w^{(t)})-F(w^{*})]+\frac{L\eta^{2}\sigma^{2}}{2T^{2}}\blacksquare.$
+Quantitative Trade‐off between Temperature and Learning Rate.
+Starting from
+$\mathbb{E}[F(w^{(t+1)})-F(w^{*})]\;\leq\;$
+$\underbrace{\bigl(1-\tfrac{\eta\mu}{T}\bigr)}_{\rho}\,\mathbb{E}[F(w^{(t)})-F(w^{*})]+\frac{L\,\eta^{2}\,\sigma^{2}}{2\,T^{2}},$
+we identify:
+  • Contraction factor: $\rho=1-\eta\mu/T$. To ensure $\rho<1$, one needs
+    $0<\eta<\frac{T}{\mu}.$
+    A larger $T$ (for fixed $\eta$) yields $\rho$ closer to 1, i.e., slower exponential decay.
+  • Steady‐state error term: $\frac{L\,\eta^{2}\,\sigma^{2}}{2\,T^{2}}\,$, which scales like $(\eta/T)^{2}$.
+Hence after $t$ rounds,
+$\mathbb{E}[F(w^{(t)})-F(w^{*})]\approx\rho^{t}\,[F(w^{(0)})-F(w^{*})]+O\!\bigl((\eta/T)^{2}\bigr).$
+Noting $\rho^{t}\approx\exp(-t\,\eta\mu/T)$, the effective learning rate is $\alpha=\eta/T$.
+Optimal Effective Stepsize and Temperature Scheduling Define $\alpha=\eta/T$. Then:
+$\rho=1-\alpha\,\mu,\quad\text{steady‐state error}=O(\alpha^{2}).$
+Balancing convergence speed and variance suggests a diminishing $\alpha$, e.g.
+$\alpha_{t}=\frac{1}{\mu\,t},\quad\Longrightarrow\quad\eta_{t}=\frac{T}{\mu\,t}.$
+Such a schedule ensures both rapid early descent and vanishing steady‐state error.
+Summary.
+  • Both contraction speed and variance error depend on the ratio $\eta/T$.
+  • Higher temperature $T$ for fixed $\eta$ slows convergence and reduces variance term.
+  • A time‐decaying stepsize $\eta_{t}\propto T/t$ attains the best trade‐off.
+Adaptive Optimization: Adam
+We now extend the above convergence analysis to the widely used Adam optimizer [kingma2014adam], incorporating both temperature scaling $T$ and adaptive per‐coordinate learning rates.
+Assumption 4. Bounded Gradients & Second Moments. For all $k,t$ and coordinates $i$,
+$|\nabla_{i}F_{k}(w^{(t)})|\leq G,\quad\mathbb{E}\bigl[\nabla_{i}F_{k}(w^{(t)})^{2}\bigr]\leq G^{2}.$
+Let $\beta_{1},\beta_{2}\in[0,1)$ be the Adam decay rates, and $\epsilon>0$ a small constant.
+1. Local Adam update on client $k$ at round $t$:
+$m_{k}^{(t)}$ $=\beta_{1}\,m_{k}^{(t-1)}\;+\;(1-\beta_{1})\,\nabla F_{k}(w_{k}^{(t)};T),$
+$v_{k}^{(t)}$ $=\beta_{2}\,v_{k}^{(t-1)}\;+\;(1-\beta_{2})\,\bigl[\nabla F_{k}(w_{k}^{(t)};T)\bigr]^{2},$
+$\hat{m}_{k}^{(t)}$ $=\frac{m_{k}^{(t)}}{1-\beta_{1}^{t}},\quad\hat{v}_{k}^{(t)}=\frac{v_{k}^{(t)}}{1-\beta_{2}^{t}},$
+$w_{k}^{(t+1)}$ $=w_{k}^{(t)}\;-\;\eta\;\frac{\hat{m}_{k}^{(t)}}{\sqrt{\hat{v}_{k}^{(t)}}+\epsilon}.$
+2. Global aggregation:
+$w^{(t+1)}\;=\;\sum_{k=1}^{K}p_{k}\,w_{k}^{(t+1)}.$
+Under Assumptions 1–4 and for a suitably chosen constant stepsize $\eta$, one can show [reddi2019convergence] that
+$\frac{1}{T}\,\nabla F_{k}(w)\Longrightarrow\frac{1}{T}\,\mathbb{E}\bigl[\|\nabla F(w^{(t)})\|_{1}\bigr]\;\leq\;O\!\Bigl(\tfrac{1}{\sqrt{T\,t}}\Bigr)+O\!\Bigl(\tfrac{1}{T\,t}\Bigr).$
+More precisely, after $T_{\max}$ total gradient steps (across all clients and rounds), the averaged gradient norm satisfies
+$\frac{1}{T_{\max}}\sum_{t=1}^{T_{\max}}\mathbb{E}\bigl[\|\nabla F(w^{(t)})\|_{1}\bigr]\;\leq\;\frac{D_{1}}{\sqrt{T\,T_{\max}}}+\frac{D_{2}}{T\,T_{\max}},$
+for constants $D_{1},D_{2}$ depending on $G,\beta_{1},\beta_{2},\epsilon,L$ and the initialization.
+Trade‐offs under Adam & Temperature Scaling
+  • Temperature scaling reduces each client’s effective gradient by $1/T$, slowing convergence by a factor $1/\sqrt{T}$ in the dominant $O(1/\sqrt{T\,t})$ term.
+  • Adam’s normalization by $\sqrt{\hat{v}_{k}^{(t)}}$ mitigates the noise variance term, yielding a faster $\tilde{O}(1/\sqrt{t})$ decay even when gradients are highly heterogeneous across clients.
+  • Define $\tilde{\alpha}=\eta/T$.
+Then the leading bound scales as $O\bigl(1/(\sqrt{\tilde{\alpha}\,t})\bigr)$, suggesting one should choose $\eta\propto T$ and possibly decay $\eta_{t}\propto T/\sqrt{t}$ to balance convergence and steady‐state variance.
+Summary.
+  • Incorporating Adam yields an $O(1/\sqrt{T\,t})$ convergence rate under temperature‐scaled gradients.
+  • Temperature $T$ still slows down the convergence rate via a $1/\sqrt{T}$ factor.
+  • A practical schedule is $\eta_{t}=\Theta\bigl(T/\sqrt{t}\bigr)$ to achieve the best trade‐off between bias and variance under adaptive optimization.
+Adaptive Aggregation: FedAdam
+FedAdam [reddi2020adaptive] is a variant of FedAvg that incorporates the adaptive moment estimates of Adam at the server side.
+In FedAdam, each client still performs local SGD (or optionally local Adam) updates, but the global aggregation step applies an Adam‐style update to the server’s model state.
+Concretely, let
+$g^{(t)}\;=\;\sum_{k=1}^{K}p_{k}\,\bigl(w_{k}^{(t+1)}-w^{(t)}\bigr)$
+be the aggregated model increment. FedAdam maintains server‐side moment estimates
+$m^{(t)}$ $=\beta_{1}\,m^{(t-1)}+(1-\beta_{1})\,g^{(t)},$
+$v^{(t)}$ $=\beta_{2}\,v^{(t-1)}+(1-\beta_{2})\,\bigl(g^{(t)}\bigr)^{2},$
+$\hat{m}^{(t)}$ $=\frac{m^{(t)}}{1-\beta_{1}^{t}},\quad\hat{v}^{(t)}=\frac{v^{(t)}}{1-\beta_{2}^{t}},$
+and updates the global model by
+$w^{(t+1)}=w^{(t)}-\eta_{s}\;\frac{\hat{m}^{(t)}}{\sqrt{\hat{v}^{(t)}}+\epsilon},$
+where $\eta_{s}$ is the server‐side stepsize and $(\beta_{1},\beta_{2},\epsilon)$ are the usual Adam hyperparameters.
+Trade‐off between FedAdam and Local Adam ++ Aggregation
+  • Server‐side adaptivity vs.
+client‐side noise: FedAdam smooths the noisy aggregated gradients $g^{(t)}$ using $m^{(t)},v^{(t)}$, reducing variance at the server.
+In contrast, applying Adam locally on each client before averaging retains some heterogeneity in second‐moment estimates across clients.
+  • Effective learning rate scaling: With temperature scaling $T$, each client’s model increment scales as $g^{(t)}\propto\eta/T$.
+The server‐side Adam then applies another $\eta_{s}/T$ factor to the update, so the combined effective stepsize is $\alpha_{\mathrm{FedAdam}}\approx(\eta\,\eta_{s})/T^{2}$.
+  • Convergence bound: Under similar bounded‐gradient assumptions, one can show (extending the analysis of [reddi2019convergence]) that the server‐averaged gradient norm satisfies
+    $\frac{1}{T_{\max}}\sum_{t=1}^{T_{\max}}$ $\mathbb{E}\|\nabla F(w^{(t)})\|_{1}\;\leq\;$
+    $O\!\Bigl(\tfrac{1}{\sqrt{T^{2}\,T_{\max}}}\Bigr)+O\!\Bigl(\tfrac{1}{T^{2}\,T_{\max}}\Bigr),$
+    highlighting the $1/T^{2}$ slowdown due to two stages of scaling (client and server).
+  • Practical scheduling: To balance convergence speed and variance, one may choose
+    $\eta_{t}=\Theta\!\bigl(\tfrac{T}{\sqrt{t}}\bigr),\quad\eta_{s,t}=\Theta\!\bigl(\tfrac{T}{\sqrt{t}}\bigr),$
+    yielding an overall decay $O\bigl(1/\sqrt{t}\bigr)$ despite the double scaling.
+Summary.
+FedAdam augments federated learning with server‐side adaptivity, further damping heterogeneity and noise but introducing an additional temperature scaling factor.
+Its convergence rate inherits a $1/\sqrt{T^{2}\,t}$ dependence, suggesting careful tuning of both client and server learning rates in proportion to $T$.
+$\bullet$ Details on Centered Kernel Alignment (CKA)
+CKA as a feature-space alignment metric in our study
+In this work, we use Centered Kernel Alignment (CKA) as our primary metric to quantify feature-space alignment between neural networks.
+CKA is widely used to compare internal representations across models because it is architecture-agnostic and can be computed layer-wise.
+In our experiments (Section 4.6), we mainly measure CKA between a global model trained under non-i.i.d.
+federated settings and a reference model trained on an i.i.d.
+configuration.
+This allows us to assess how closely the federated model recovers the feature space of an ideal centralized learner.
+Empirically, we observe that lower training temperatures lead to faster and stronger feature-space alignment with the i.i.d.
+reference model, especially in deeper layers.
+As shown in Figure 12 and Table 6, models trained with small temperatures such as $T=0.25$ and $T=0.05$ achieve higher CKA values in the penultimate and classification layers and reach these levels in fewer rounds than $T=1.0$ or higher.
+This behavior closely follows the accuracy convergence trends: configurations that exhibit faster CKA growth also converge more rapidly and reach higher final accuracy.
+In other words, CKA in our study is not used in isolation but is interpreted jointly with test accuracy and convergence speed, and together they provide evidence that low-temperature training accelerates both representation alignment and optimization.
+We also study feature-space similarity at the client level (Figure 13), where we compare each client model with the initial global model, the aggregated model, and the i.i.d.
+reference model.
+For low-temperature configurations, clients quickly learn representations that are both diverse across clients and better aligned with the reference model, and these representations are effectively aggregated at the server.
+This pattern is consistent with the observation that FLex&\&Chill promotes faster global convergence while preserving useful client-specific information.
 (a) Case 1
@@
 Figure 17: Illustration of possible cases in a federated learning scenario.
-The solid orange line and dotted red line correspond to the local update direction and aggregated model update direction respectively.
-The black circle and the red star depict the initial point and optimal point respectively.
-$\bullet$ High similarity between local models does not guarantee a better model (Case 1, 2): The first case arises when local models, each trained on their respective local data, produce highly similar updates due to a well-aligned data configuration.
-Such alignment implies that the training data across different clients are relatively similar in distribution and characteristics, allowing the aggregated global model to closely approach the ideal model achievable through centralized training.
-However, in non-i.i.d.
-environments, achieving this scenario is infrequent.
-In non-i.i.d.
-settings, the data distribution varies significantly across different local nodes.
-This variation can lead to local models learning disparate aspects of the data, resulting in updates that are less similar.
-Consequently, the aggregated global model may not perform as well as in an i.i.d.
-scenario, as it has to reconcile these diverse and potentially conflicting updates.
-On the other hand, Case 2 represents a critical scenario in federated learning where there is a high similarity among local model updates, but these updates are aligned in a suboptimal or incorrect direction.
-This situation can be considered one of the worst cases in federated learning.
-In addition, using these two cases, we demonstrate that the high similarity between local models does not ensure a superior update or quality of training.
-$\bullet$ A superior model can be obtained without similar local models (Case 3): Case 3 illustrates a case where a superior model can be achieved even when local models exhibit relatively low similarity.
-In this scenario, each local model learns effective but diverse features from heterogeneous data configurations, and these diverse updates are properly aggregated at the central server.
-Unlike scenarios where local models produce highly similar updates, in this case, the updates from different local models show significant variation.
-This diversity reflects the different aspects of the data each local model has learned.
-This situation arises under conditions of data heterogeneity, where the data distributions across different clients are diverse.
-Each local model captures unique features relevant to its specific data subset.
-Despite the low similarity among local model updates, the central server effectively aggregates these diverse updates.
-The aggregation process leverages the complementary information from each local model, resulting in a robust and well-generalized global model.
-This case supports the hypothesis that a proper global model can be obtained even when local models do not exhibit high similarity and is a globally ideal case for federated learning scenarios.
-$\bullet$ Relationship between model similarity and the magnitude of updates (Case 4, 5): It is important to note that the magnitude of updates can influence the perceived similarity of the network.
-If the update magnitude is small, the similarity of the network may be measured as high, whereas if the update magnitude is large, the similarity may be measured as relatively low.
-This example underscores the need to rethink the conventional emphasis on high similarity as an indicator of performance.
-It suggests that focusing solely on similarity metrics, while providing insights on how local training operations induce model changes, might overlook other critical factors that contribute to the effectiveness of the aggregated model.
-Therefore, a more nuanced approach to evaluating model updates and their impact on the accuracy performance is required, considering both the magnitude of updates and their alignment with the overall learning objectives.
-Based on the aforementioned discussions, we introduce measuring similarity with a reference model as an alternative evaluation method in this work.
-Although we cannot ensure that the reference model trained in an i.i.d.
-data configuration perfectly represents the optimal update objective, we consider this method as an effective and reasonable way to evaluate the federated learning process.
-This approach allows for a more nuanced assessment of the updates by comparing them to a standard benchmark, helping to identify whether the aggregated model is progressing toward a semi-optimal solution despite the inherent data heterogeneity.
-E.3 Experimental results
-Table 8 presents the exact numbers for the accuracy and CKA values discussed in Section 4.4 and Figure 5. Figure 15 illustrates the data distribution applied in the experiments.
-Temperature ($T$) Accuracy (%) Layer 1 CKA Layer 2 CKA Layer 3 CKA Layer 4 CKA 4.00 49.72 0.6539 0.5632 0.7149 0.7661 2.00 53.82 0.6792 0.6872 0.7417 0.7963 1.00 58.19 0.7419 0.6997 0.7644 0.8214 0.50 63.48 0.9322 0.7459 0.8181 0.8559 0.25 66.85 0.9569 0.8161 0.8418 0.8830 0.05 72.76 0.9390 0.9059 0.8898 0.9026 Reference Model ($T$=1.00, i.i.d.) 73.86 1.0000 1.0000 1.0000 1.0000
-Table 8: Accuracy and CKA values of 2 layer CNN measured on the CIFAR10 dataset in Figure 5.
-Figure 15: Visualization of the distribution of training data used in CKA evaluation. Best viewed in color.
-(a) $T$=0.05
-(b) $T$=0.25
-(c) $T$=0.5
-(d) $T$=1.0
-(e) $T$=2.0
-(f) $T$=4.0
-Figure 16: Client-wise feature space similarity matrix measured in the first convolution layers (Layer 1) by using CKA and inference accuracy in different temperatures (0.05, 0.25, 0.5, 1.0, 2.0, 4.0) and federated rounds (1, 5, 50, 100).
-(INIT: Initial dispatched model in the given round, AGG: Aggregated global model, REF: Reference model, ACC: Accuracy)
-(a) $T$=0.05
-(b) $T$=0.25
-(c) $T$=0.5
-(d) $T$=1.0
-(e) $T$=2.0
-(f) $T$=4.0
-Figure 17: Client-wise feature space similarity matrix measured in the second convolution layers (Layer 2) by using CKA and inference accuracy in different temperatures (0.05, 0.25, 0.5, 1.0, 2.0, 4.0) and federated rounds (1, 5, 50, 100).
-(INIT: Initial dispatched model in the given round, AGG: Aggregated global model, REF: Reference model, ACC: Accuracy)
-(a) $T$=0.05
-(b) $T$=0.25
-(c) $T$=0.5
-(d) $T$=1.0
-(e) $T$=2.0
-(f) $T$=4.0
-Figure 18: Client-wise feature space similarity matrix measured in penultimate layers (Layer 3) by using CKA and inference accuracy in different temperatures (0.05, 0.25, 0.5, 1.0, 2.0, 4.0) and federated rounds (1, 5, 50, 100).
-(INIT: Initial dispatched model in the given round, AGG: Aggregated global model, REF: Reference model, ACC: Accuracy)
-(a) $T$=0.05
-(b) $T$=0.25
-(c) $T$=0.5
-(d) $T$=1.0
-(e) $T$=2.0
-(f) $T$=4.0
-Figure 19: Client-wise feature space similarity matrix measured in classification layers (Layer 4) by using CKA and inference accuracy in different temperatures (0.05, 0.25, 0.5, 1.0, 2.0, 4.0) and federated rounds (1, 5, 50, 100).
-(INIT: Initial dispatched model in the given round, AGG: Aggregated global model, REF: Reference model, ACC: Accuracy)
+The solid orange line and dotted red line correspond to the local update direction and aggregated model update direction, respectively.
+The black circle and the red star depict the initial point and an ideal high-accuracy model in the representation space.
+Limitation and interpretation of CKA in federated learning
+Despite its usefulness, CKA alone does not guarantee high model quality.
+This subsection explains how we interpret CKA in the context of federated learning and why we additionally compare against a reference model.
+Figure 17 illustrates several possible cases in an abstract representation space.
+The orange line denotes local update directions for client models starting from a shared global model (black circle), and the red dotted line represents the effective aggregated update at the server.
+The red star indicates an ideal model with high accuracy.
+$\bullet$ High similarity between local models does not guarantee a better model (Case 1, 2): In Case 1, local models produce highly similar updates because their data distributions are well aligned.
+In such a near-i.i.d.
+scenario, the aggregated global model can approach the ideal centralized solution.
+However, this case is relatively rare in realistic non-i.i.d.
+settings.
+In Case 2, local model updates are also highly similar but are aligned in a suboptimal direction.
+Even though the feature spaces of the local models are similar to each other, the global update may move away from the desirable region, resulting in poor accuracy.
+These cases highlight that high similarity among local models or their updates does not, by itself, ensure good performance.
+$\bullet$ A superior model can be obtained without similar local models (Case 3): Case 3 demonstrates that a strong global model can emerge even when local models exhibit relatively low pairwise similarity.
+Under heterogeneous data distributions, each client may learn distinct but useful features.
+If these diverse updates are properly aggregated, the resulting global model can achieve high accuracy and robust generalization.
+This scenario aligns with our empirical findings where low-temperature configurations encourage more aggressive and diverse local updates that still aggregate into a better-aligned and better-performing global model.
+In this sense, FLex&\&Chill tends to behave closer to Case 3 than Case 2.
+$\bullet$ Relationship between model similarity and the magnitude of updates (Case 4, 5): Cases 4 and 5 emphasize that the magnitude of local updates also affects measured similarity.
+When update magnitudes are very small, models can appear highly similar even if they have not moved meaningfully toward the optimal solution.
+Conversely, large but beneficial updates may temporarily reduce similarity while still improving accuracy.
+This observation suggests that focusing solely on absolute similarity values, without considering update magnitude and direction, can be misleading.
+Based on these considerations, we do not treat CKA as a standalone proxy for performance.
+Instead, we measure CKA with respect to a reference model trained in an i.i.d.
+configuration and analyze its evolution together with accuracy.
+This reference-based comparison provides a more informative baseline for assessing how closely federated models recover an ideal feature space.
+In our experiments, federated models trained with lower temperatures consistently achieve higher CKA values relative to this reference and do so more quickly, which correlates well with faster convergence and higher final accuracy.
+Regarding extreme non-i.i.d.
+scenarios (for example, very small Dirichlet concentration parameters or nearly disjoint class partitions), our results indicate that absolute CKA values can indeed be lower due to stronger representation drift, which is consistent with prior observations in federated learning.
+However, within the same non-i.i.d.
+setting,FLex&\&Chill still yields higher CKA values and better accuracy than training with $T=1.0$, indicating that low-temperature training partially compensates for drift by amplifying informative gradients and promoting more effective feature-space alignment.
+We therefore view CKA, when interpreted relative to a reference model and in conjunction with accuracy, as a meaningful but not sufficient indicator for understanding how Logit Chilling shapes representations in heterogeneous federated environments.
+$\bullet$ Experiment on Hyperparameters
+In this section, we explore the effect of other federated learning hyperparameters (i.e., number of participants (Figure 18), batch size (Figure 19), and number of local epochs (Figure 20)).
+Through this set of experiments, we observed that applying low temperatures during local training improves test accuracy and convergence speed, being independent of hyperparameters.
+We note that the experiments were done by applying the default setting except for the control factors.
+One can view participant count, batch size, and local epochs through the lens of gradient variance.
+Sampling more clients per round increases the chance that the aggregated update better approximates the true gradient, reducing variance.
+Likewise, using smaller batch sizes with more local epochs yields more frequent updates that capture finer-grained information from each client’s data, which also dampens variance.
+Because low temperatures can sometimes introduce training instability, tuning these variance-reducing parameters, more clients, smaller batches, or additional local epochs may help stabilize convergence at low temperatures.
+(a) Number of participants 5
+(b) Number of participants 10
+(c) Number of participants 15
+Figure 18: Test accuracy and test loss over federated learning rounds with varying the number of participants and the training temperature on the CIFAR10 dataset.
+The shaded areas indicate the standard deviation across different runs.
+When relatively many clients participate in the training, it shows better training quality with stability.
+(a) Batch size 8
+(b) Batch size 16
+(c) Batch size 32
+(d) Batch size 64
+(e) Batch size 128
+Figure 19: Test accuracy and test loss over federated learning rounds with varying the batch size and the training temperature on the CIFAR10 dataset.
+The shaded areas indicate the standard deviation across different runs.
+A relatively small batch size shows better training quality with stability.
+(a) Number of local epoch 1
+(b) Number of local epoch 5
+(c) Number of local epoch 10
+(d) Number of local epoch 15
+Figure 20: Test accuracy and test loss over federated learning rounds with varying the number of local epochs and the training temperature on the CIFAR10 dataset.
+The shaded areas indicate the standard deviation across different runs.
+Relatively many updates in the local training phase show better training quality with stability.
+$\bullet$ Extended Stability Analysis Under Long-Horizon Training
+To further assess the stability of low-temperature training beyond the main experimental horizon, we conducted an extended study with 1,000 federated training rounds.
+This experiment aims to evaluate whether the aggressive updates induced by Logit Chilling remain stable under prolonged optimization.
+Figure 21 presents the convergence trajectories for multiple temperature settings.
+Across all configurations, models trained with lower temperatures continue to exhibit stable convergence behavior.
+Notably, temperatures below one consistently outperform the standard/high temperature setting ($T\in\{1.0,2.0,4.0\}$) throughout the extended training horizon, reaffirming the robustness and long-term benefits of applying temperature scaling during local training.
+These results demonstrate that the accelerated convergence observed in the main experiments does not introduce long-term instability, even when the training duration is substantially increased.
+Figure 21: Average global test accuracy and test loss for 1,000 federated training rounds across different temperature settings.
+Lower temperatures maintain stable convergence and continue to outperform the standard/high temperature ($T\in\{1.0,2.0,4.0\}$).
+$\bullet$ Analysis on Model Calibration
+Figure 22: Reliability diagrams illustrating the calibration performance of models trained with different temperatures.
+Each diagram compares predicted confidence against empirical accuracy, showing that temperature scaling during training does not introduce notable miscalibration.
+Figure 23: Prediction confidence distributions of models trained with varying temperatures.
+The histograms show how the confidence levels are distributed across samples, providing insight into potential over- or under-confidence behaviors under different temperature settings.
+Figure 24: Expected Calibration Error (ECE) for models trained with different temperatures.
+Error bars indicate standard deviation across runs.
+The results demonstrate minimal differences in calibration quality across temperature settings.
+Since temperature scaling was originally introduced for model calibration, one might be concerned that training with low temperatures could lead to over- or under-confidence.
+Model calibration measures how well a model’s predicted probabilities correspond to the true likelihood of correctness.
+A well-calibrated model outputs high confidence only when its predictions are likely to be correct, and expresses appropriate uncertainty when the outcome is less certain.
+Thus, calibration evaluates not only what the model predicts but also how confident it is in those predictions.
+For example, a model that frequently produces high-confidence predictions but achieves only moderate accuracy is considered over-confident, whereas a model that predicts with low confidence despite high accuracy is under-confident.
+A standard quantitative metric for evaluating calibration is the Expected Calibration Error (ECE), defined as
+$\mathrm{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}\left|\mathrm{acc}(B_{m})-\mathrm{conf}(B_{m})\right|,$
+where $B_{m}$ denotes the set of samples whose predicted confidences fall into the $m$-th bin, $\mathrm{acc}(B_{m})$ is the empirical accuracy for that bin, $\mathrm{conf}(B_{m})$ is the average predicted confidence, and $n$ is the total number of samples.
+Lower ECE values indicate better calibration.
+To assess calibration quality in our setting, we examine the prediction probability distributions, the model’s average confidence, and its accuracy.
+Concretely, we present the calibration results (Figure 22), the prediction confidence distributions (Figure 23), and the corresponding ECE values (Figure 24) for models trained with different temperatures (0.05, 0.25, 0.5, 1.0, 2.0, 4.0).
+We note that we use the same temperature for testing as for training.
+Figure 23 shows that models trained with higher temperatures produce fewer samples in the high-confidence region, which is consistent with their lower performance at the same training round and indicates that they are less certain about the given inputs.
+In contrast, models trained with lower temperatures assign higher confidence to a larger portion of samples, and Figure 22 shows that these high-confidence regions remain well calibrated: the empirical accuracy in the highest confidence bins does not deteriorate, and the reliability curves stay close to the diagonal.
+Finally, as shown in Figure 24, across our experimental settings, we do not observe meaningful differences in ECE among models trained with different temperatures.
+Taken together, the ECE values, confidence distributions, and reliability diagrams indicate that applying temperature scaling during training does not introduce notable calibration degradation, nor systematic over-confidence or under-confidence, in our setting.
+$\bullet$ Comparison to Knowledge Distillation
+In this section, we clarify how our use of temperature scaling in FLex&\&Chill differs fundamentally from conventional knowledge distillation techniques.
+Although both rely on a temperature parameter applied to the softmax function, their goals, where the temperature is applied, and the corresponding training regimes are distinct.
+In classical knowledge distillation, a high-capacity teacher model is first trained in a centralized manner.
+A temperature value larger than 1 is then applied to the teacher’s logits to produce softened probability distributions.
+These softened probabilities are used as soft labels to train a student model, typically by minimizing a distillation loss such as the Kullback–Leibler divergence between the teacher and student outputs, often combined with a standard cross-entropy loss on hard labels.
+The key role of temperature in this setting is to smooth the teacher’s output distribution so that relative relationships among non-argmax classes become more informative to the student.
+Importantly, in conventional knowledge distillation, the temperature is introduced primarily to generate softened targets and to define the distillation loss: both the teacher and the student outputs are typically divided by the same temperature $T>1$ inside the softmax when computing the distillation term, while the hard-label loss and inference-time predictions of the student still use temperature $T=1$.
+In other words, the temperature is mostly confined to the soft-label matching objective, rather than being used as a global training knob that systematically controls all aspects of the student’s optimization dynamics.
+In contrast, our work does not involve a teacher–student architecture or soft-label transfer.
+FLex&\&Chill applies temperature scaling directly to the softmax of the model that is being trained on each client, and this is done throughout the federated training process rather than as a post-training calibration or label-smoothing step.
+Moreover, we operate in the low-temperature regime where the temperature lies between 0 and 1, which sharpens the output distribution and, as analyzed in our convergence theory, amplifies gradient magnitudes and accelerates optimization in the presence of non-i.i.d.
+data.
+In other words, temperature in our setting is used as an optimization control knob that shapes gradient flow and feature-space alignment across heterogeneous clients, not as a mechanism for compressing or transferring knowledge from a pre-trained teacher network.
+These differences imply that FLex&\&Chill is complementary to, rather than a variant of, conventional knowledge distillation.